### v1

# Extending Mean-Field Variational Inference via Entropic Regularization: Theory and Computation

Abstract
Variational inference (VI) has emerged as a popular method for approximate inference for high-dimensional Bayesian models. In this paper, we propose a novel VI method that extends the naive mean field via entropic regularization, referred to as $\Xi$-variational inference ($\Xi$-VI). $\Xi$-VI has a close connection to the entropic optimal transport problem and benefits from the computationally efficient Sinkhorn algorithm. We show that $\Xi$-variational posteriors effectively recover the true posterior dependency, where the dependence is downweighted by the regularization parameter. We analyze the role of dimensionality of the parameter space on the accuracy of $\Xi$-variational approximation and how it affects computational considerations, providing a rough characterization of the statistical-computational trade-off in $\Xi$-VI. We also investigate the frequentist properties of $\Xi$-VI and establish results on consistency, asymptotic normality, high-dimensional asymptotics, and algorithmic stability. We provide sufficient criteria for achieving polynomial-time approximate inference using the method. Finally, we demonstrate the practical advantage of $\Xi$-VI over mean-field variational inference on simulated and real data.
KEYWORDS: Variational inference, Optimal transport, Mean-field approximation, Statistical-computational tradeoff, High-dimensional Bayesian inference.
1 Introduction
Variational inference (VI) is a widely used method for approximate probabilistic inference. VI approximates a difficult-to-compute distribution by positing a family of simpler distributions and minimizing the KL divergence between the family and the target. In Bayesian modeling, which is a common application of VI, the target is a posterior distribution of latent variables given observations $\textrm{p}(\theta\,|\,\mathbf{x})$ and the variational family is of distributions of the latent variables $\textrm{q}(\theta)\in\mathbb{Q}(\Theta)$. VI approximates the posterior with
$\textrm{q}^{*}(\theta)=\arg\min_{\textrm{q}\in\mathbb{Q}}D_{\textsc{kl}}\left(\textrm{q}(\theta)\parallel\textrm{p}(\theta\,|\,\mathbf{x})\right).$ (1.1)
To set up a variational inference, we need to select the family of distributions over which to optimize. In many applications, practitioners use the mean-field or fully factorized family. This is the family of product distributions, i.e., where each variable is independent and endowed with its own distributional factor. For example, if there are $D$ latent variables $\theta=\{\theta_{1},\ldots,\theta_{D}\}$ then the corresponding mean-field family is
$\textrm{q}(\theta)=\prod_{j=1}^{D}\textrm{q}_{j}(\theta_{j}).$ (1.2)
Thanks to this simple family, the variational optimization is computationally efficient (to a local optimum). But this efficiency comes at a cost, and mean-field VI suffers in accuracy [13].
In this paper, we develop a new way of doing variational inference. The idea is to optimize over all distributions of the latent variables, i.e., $\textrm{q}\in\mathbb{P}(\Theta)$, but to regularize the variational objective function to encourage simpler distributions that are “more like the mean-field.” At one end of the regularization path we effectively optimize over the mean-field family, providing traditional mean-field VI (MFVI). At the other end we optimize over all distributions, providing exact inference (but at prohibitive cost). Between these extremes, our method smoothly trades off efficiency and accuracy.
In detail, consider a probabilistic model $\textrm{p}(\theta,\mathbf{x})=\textrm{p}(\theta)\textrm{p}(\mathbf{x}\,|\,\theta)$ and the goal to approximate the posterior $\textrm{p}(\theta\,|\,\mathbf{x})$. Denote the prior $\pi(\theta):=\textrm{p}(\theta)$ and the log likelihood $\ell(\mathbf{x}\,;\,\theta):=\log\textrm{p}(\mathbf{x}\,|\,\theta)$. We propose to approximate the posterior by optimizing an entropy-regularized variational objective over the entire space of distributions $\mathbb{P}(\Theta)$,
$\textrm{q}_{\lambda}^{*}(\theta)=\arg\max_{\textrm{q}\in\mathbb{P}(\Theta)}\mathbb{E}_{\textrm{q}}\left[\ell(\mathbf{x}\,;\,\theta)\right]-D_{\textsc{kl}}\left(\textrm{q}(\theta)\parallel\pi(\theta)\right)-\lambda D_{\textsc{kl}}\left(\textrm{q}(\theta)\parallel\prod_{j=1}^{D}\textrm{q}_{j}(\theta_{j})\right).$ (1.3)
In this objective, the first two terms comprise the usual evidence lower bound (ELBO) [34, 13]. When optimized relative to the full set of distributions of $\theta$, maximizing the ELBO provides the variational representation of the exact posterior. The third term, however, is a penalty term. It encourages the optimal q to resemble a factorized distribution. By varying the regularizer $\lambda$, we interpolate between the exact posterior and its factorized approximation.
We will study the theory and application of Eq. 1.3, which we call $\Xi$-VI (pronounced “ksee VI”). First we show that we can solve this optimization by iterating between (1) calculating approximate posterior marginals for each variable and (2) solving a problem of entropic optimal transport (EOT) with a multi-marginal Sinkhorn algorithm [22, 41]. We then approximate this solution in two steps, where we first use traditional VI to approximate the marginals, e.g., with black-box variational inference [57] or expectation propagation [45], and then solve the EOT problem.
We prove that $\Xi$-VI gives Frequentist guarantees including posterior consistency and the Bernstein-von Mises theorem. Further, we show how to theoretically reason about the regularization parameter to optimize the statistical-computation trade-off that $\Xi$-VI implies. Specifically, we characterize high-dimensional regimes where $\Xi$-VI is either mean-field or Bayes-optimal. Empirically, we apply $\Xi$-VI to multivariate Gaussians, linear regression with a Laplace prior, and hierarchical modeling of the 8-schools dataset. We find that in addition to outperforming mean-field VI and other VI methods, $\Xi$-VI undergoes a phase transition as we vary the regularization parameter, with evidence of $\lambda=D$ being a critical threshold.
The rest of the paper is structured as follows. Section 2 introduces $\Xi$-VI. Section 2.4 addresses computational complexities in implementing it. Section 3 establishes theoretical guarantees for the $\Xi$-variational posterior, including posterior consistency, Bernstein-von Mises theorem, high-dimensional bounds, finite-sample convergence, and algorithmic stability. Section 4 provides an empirical study. Section 5 concludes the paper.
Related Work.
This work draws upon several lines of research on the statistical and computational guarantees of VI.
In recent years, many researchers have studied the statistical theory around VI. Some of the results include asymptotic normality [32, 31, 12, 72], contraction rates [79, 78], nonasymptotic bounds [2, 1, 76], and properties in high-dimensional models [7, 60, 59, 47, 36, 48, 56]. Our works add to the existing theory by establishing Frequentist guarantees for a new class of non-mean-field variational methods.
Other research has explored computational guarantees for VI. These results include convergence rates of coordinate ascent [49, 54, 78, 74, 11], convergence of black-box variational inference [37], and the statistical-computational tradeoff in Gaussian VI [9]. A related thread studies VI using gradient flow techniques [77, 40, 23, 33]. Our work continues the quest of [9] to explore statistical-computational tradeoff as we move beyond the mean-field VI, while expanding the interface between OT and VI by discovering novel ways to deploy entropic OT tools in VI.
Mean-field VI is efficient, but also suffers from limitations. Some of these limitations have been highlighted in multivariate Gaussians [13], state space models [70], piecewise constant models [79], and spike covariance models [28]. Several methods have been proposed to address these failure modes, including structural VI [73, 58], Copula VI [64, 65], linear response correction [30, 61], TAP correction [51, 26, 16, 17]. However, these methods generally lack statistical and computational theory, or are problem-specific. A related method is variational boosting [43, 42], which corrects for multi-modality in the true posterior. Our method, $\Xi$-VI, adds to the growing non-mean-field toolbox as an automatic and theoretically principled variational method that produces dependent posterior approximation with no extra modeling assumptions.
2 $\Xi$-variational inference
Again, we consider a general probabilistic model
$\textrm{p}(\theta,\mathbf{x})=\pi(\theta)\exp\{\ell(\mathbf{x}\,;\,\theta)\},$ (2.1)
where $\pi(\theta)$ is the prior of the unknown parameter and $\ell(\mathbf{x}\,;\,\theta)$ is the log likelihood of the data under $\theta$. Our goal is to find a distribution $q\in\mathbb{P}(\Theta)$ to approximate the posterior $\textrm{p}(\theta\,|\,\mathbf{x})$.
2.1 The $\Xi$ variational objective
Consider a distribution $\textrm{q}\in\mathbb{P}(\Theta)$ with marginal distributions $\textrm{q}_{i}\in\mathbb{P}(\Theta_{i})$. We define the expressivity functional as the KL divergence between q and the product of its marginals:
$\Xi(\textrm{q})=D_{\textsc{kl}}\left(\textrm{q}(\theta)\parallel\prod_{i=1}^{D}\textrm{q}_{i}(\theta_{i})\right).$ (2.2)
Expressivity quantifies the strength of dependence of q as a joint distribution of $D$ random variables. In the language of information theory, it is the multivariate mutual information of a random variable $\theta\sim\textrm{q}(\theta)$. In the language of variational inference, it quantifies the deviation of $\textrm{q}(\theta)$ from its KL-projection on the mean-field family.
With expressivity in hand, we define $\Xi$-variational inference ($\Xi$-VI) as the penalized optimization problem in Eq. 1.3. In this problem, $\lambda\geq 0$ is a user-defined regularization parameter. The optimal $\textrm{q}_{\lambda}^{*}(\theta)$ is called the $\Xi$-variational posterior. When it is not unique, $\textrm{q}^{*}_{\lambda}$ is one of the distributions that solves Eq. 1.3.
Remark.
When $\lambda=0$, $\textrm{q}^{*}_{\lambda}$ is the exact posterior. When $\lambda=\infty$, $\textrm{q}^{*}_{\infty}$ is a mean-field variational posterior.
Remark.
Eq. 1.3 is equivalent to maximizing a regularized ELBO:
$\textrm{q}^{*}_{\lambda}=\mathop{\mathrm{argmax}}_{\textrm{q}\in\mathbb{P}(\Theta)}\text{ELBO}\left(\textrm{q}\right)-\lambda\Xi(\textrm{q}),\quad\lambda\geq 0.$
In Section 2.2, we derive a procedure for solving $\Xi$-VI. Although $\Xi$-VI is in general a non-convex problem, our analysis shows it has a tractable convex substructure in close connection to the entropic optimal transport problem.
2.2 Solution via Entropic Optimal Transport
We now derive an algorithm for solving $\Xi$-VI. We will use the tools from entropic optimal transport (EOT) [67, 50]. From now on, we assume the prior $\pi$ to be a product distribution.
Let $m_{i}(\theta_{i})$ denote a marginal distribution of $\theta_{i}$ and let $\mathcal{M}(\Theta)$ denote the space of product distributions of $\theta$,
$\mathcal{M}(\Theta)=\left\{m(\theta):m(\theta)=\prod_{i=1}^{D}m_{i}(\theta_{i})\right\}.$ (2.3)
Given a set of $D$ marginals let $\mathcal{C}\left(m_{1},\ldots,m_{D}\right)$ denote the set of $D$-dimensional joint distributions where $m_{j}(\theta_{j})$ is the $j^{th}$ marginal,
$\mathcal{C}\left(m_{1},\ldots,m_{D})\right)=\{q(\theta_{1},\ldots,\theta_{D}):q_{j}(\theta_{j})=m_{j}(\theta_{j})\,,\,j=1,\ldots,D\}.$ (2.4)
The set $\mathcal{C}(m_{1},\cdots,m_{D})$ is called the set of couplings over the distributions $\{m_{1},\cdots,m_{D}\}$.
As shorthand, we write $\mathcal{C}(m)$ as the set of couplings over the marginal distributions of $m(\theta)$. Note the set $\mathcal{C}(m)$ is convex and closed in the Wasserstein distance [50], and we assume that there exists $\textrm{q}\in\mathcal{C}(m)$ with finite (Boltzmann) entropy.
With these definitions in place, we can write $\Xi$-VI as a double minimization problem,
$\min\limits_{m\in\mathcal{M}(\Theta)}\min\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}[-\ell(\bm{x};\theta)]+\lambda D_{\text{KL}}(\textrm{q}\parallel m)+D_{\text{KL}}(q\parallel\pi).$ (2.5)
The equation follows from expressing the minimization set $\mathbb{P}(\Theta)$ as $\{\textrm{q}\in\mathcal{C}(m),m\in\mathcal{M}(\Theta)\}$, while the objective stays the same.
The outer variational problem minimizes the objective function with respect to the space of marginal distributions. Given a set of marginals, the inner variational problem minimizes the objective function over its set of couplings.
2.2.1 Inner Variational Problem
We can view the inner variational problem as an entropic optimal transport (EOT) problem [50]. Here we fix $m$ and optimize $\textrm{q}^{*}$, simple derivation shows
$\textrm{q}^{*}(\theta)=\mathop{\mathrm{argmin}}\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}[-\ell(\bm{x};\theta)]+(\lambda+1)D_{\text{KL}}(\textrm{q}\parallel m).$ (2.6)
From now on, we assume that the log-likelihood function $\ell(\bm{x};\cdot)$ is uniformly bounded over $\Theta$, i.e. $\ell(\bm{x};\cdot)\in L^{\infty}(\Theta)$. Our next result states that the solution to the EOT problem Eq. 2.6 has a unique representation.
Theorem 1 (Structure Theorem for Multi-Marginal EOT).
Assume $\inf_{\textrm{q}\in\mathcal{C}(m)}-\mathbb{E}_{\textrm{q}}[\ell(\bm{x};\theta)]+(\lambda+1)\Xi(\textrm{q})<\infty$ and $|\ell(\bm{x};\theta)|<\infty,[m]-$almost surely.
Then there exists a unique minimizer $\textrm{q}^{*}$ to Eq. 2.6 that is absolutely continuous with respect to $m$ (denoted $\textrm{q}^{*}\ll m$) and
  1. 1)
    There exist measurable functions $\phi_{i}:\Theta_{i}\to\mathbb{R}$, $i\in[D]$, such that
    $\textrm{q}^{*}(\theta)=\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)m(\theta),$ (2.7)
    $[m]$-almost surely. The set of functions $\phi:=(\phi_{1},\cdots,\phi_{D})$ is called the set of EOT potentials. Each $\phi_{i}$ is $[m_{i}]$-almost surely measurable and unique up to an additive constant. Moreover, if $\mathbb{E}_{m_{i}}[\phi_{i}]\geq 0$ then $\phi_{i}\in L^{\infty}(m_{i})$ for $i\in[D]$.
  2. 2)
    Conversely, suppose $\textrm{q}\in\mathcal{C}(m)$ admits a density of the form Eq. 2.7, $[m]$-almost surely, for a set of functions $\phi_{i}:\Theta_{i}\to\mathbb{R},i\in[D]$. Then q minimizes the EOT problem Eq. 2.6 and $\phi_{i}$, $i\in[D]$, are the set of EOT potentials.
This result first appears heuristically in [15]. For $D=2$, the uniform boundedness assumption can be relaxed to $\ell(\bm{x};\cdot)$ being integrable (Theorem 4.2, [50]). See Proofs of Section 2 for the proof.
Remark.
The solution structure (Eq. 2.7) consists of three components: a log-likelihood term scaled by $1/(\lambda+1)$, a set of potential functions $\phi_{i}$’s and marginal distributions $m_{i}$’s. The regularization parameter $\lambda$ controls the temperature of the energy term $\ell(\bm{x};\theta)$. Intuitively, $\lambda$ divides a sample size of $n$ between the true posterior and the mean-field solution by a factor of $1/(\lambda+1)$ and $\lambda/(\lambda+1)$, respectively. In doing so, $\lambda$ quantifies the tradeoff between the likelihood and a product distribution. Higher $\lambda$ penalizes the variational posterior to being close to naive mean field and lower $\lambda$ allows the solution to better approximate the true posterior. When $\lambda=0$, the likelihood term is untempered and the variational solution is the true posterior. When $\lambda=\infty$, the product term dominates and the solution matches the mean-field variational posterior.
Remark.
$\Xi$-variational posterior has a connection with posterior tempering [44, 76], and it is useful to view $\textrm{q}^{*}_{\lambda}$ as a nonlinear tilt of the $1/(\lambda+1)$-tempered posterior. Write
$f_{i}(\theta_{i}):=\phi_{i}(\theta_{i})+\log m(\theta_{i})-\log\pi(\theta_{i}).$
Then we can represent $\textrm{q}^{*}_{\lambda}$ as a nonlinear tilt of the tempered posterior ${\textrm{q}^{*}_{0}}^{\lambda}$,
$\textrm{q}^{*}_{\lambda}(\theta)\propto\exp\left(\sum_{i=1}^{D}f_{i}(\theta_{i})\right){\textrm{q}^{*}_{0}}^{\lambda},\quad\text{where}\quad{\textrm{q}^{*}_{0}}^{\lambda}(\theta)\propto\exp\left(\frac{1}{\lambda+1}\ell(\bm{x};\theta)\right)\pi(\theta).$ (2.8)
This formula Eq. 2.8 extends the result that the mean-field variational posterior of the quadratic interaction model amounts to a linear tilting of the prior [69], as $f_{i}(\theta_{i})$ can be viewed as a tilting to the $1/(\lambda+1)$-tempered posterior.
Remark.
The path (or curve) formed by $\{\textrm{q}^{*}_{\lambda},\lambda\in\bar{\mathbb{R}}_{+}\}$ can be viewed as a smooth interpolation between the mean-field variational posterior and the true posterior. The interpolation is best illustrated by the multivariate Gaussian example (Section 4.1), where the regularization downweights the off-diagonal entries of the precision matrix by a factor of $1/(\lambda+1)$. The smoothness comes from the fact that $\lambda<\infty$, $\textrm{q}^{*}_{\lambda}$ matches the dependent structure of the true posterior. Varying $\lambda$ reduces the posterior dependency but induces no sharp structural change, except at $\lambda=\infty$.
We now derive a procedure for computing the $\Xi$-variational posterior. Note the EOT problem Eq. 2.6 is convex. We write down its Lagrangian dual problem:
$\phi^{*}=\arg\max\limits_{\phi\in\prod_{i=1}^{D}L^{1}(m_{i})}\sum_{i=1}^{D}\mathbb{E}_{m_{i}}[\phi_{i}(\theta_{i})]-\mathbb{E}_{m}\left[\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)\right].$ (2.9)
The solutions are identifiable up to an additive constant. For illustration, we fix the solution by imposing $D-1$ identifiability constraints:
$\mathbb{E}_{m_{1}}\phi_{1}(\theta_{1})=\cdots=\mathbb{E}_{m_{D-1}}\phi_{D-1}(\theta_{D})=0.$ (2.10)
Let $L_{p}^{\prime}(m_{i})$ denote the set of zero-mean potentials in the $L_{p}$ space with $m_{i}$ as the base measure, i.e. $L_{p}^{\prime}(m_{i})=\{\phi_{i}\in L_{p}(m_{i}):\mathbb{E}_{m_{i}}\phi_{i}=0\}$, and define the space $\bm{E}(m):=\prod_{i=1}^{D-1}L_{1}^{\prime}(m_{i})\times L^{1}(m_{D})$. We find the potentials by maximizing the dual EOT problem,
$\phi^{*}=\arg\max\limits_{\phi\in\bm{E}(m)}\sum_{i=1}^{D}\mathbb{E}_{m_{i}}[\phi_{i}(\theta_{i})]-\mathbb{E}_{m}\left[\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)\right].$ (2.11)
We can solve Eq. 2.11 with a block coordinate ascent algorithm called the Sinkhorn algorithm [22]. We describe the updates here, and discuss its implementation in detail in Section 2.4.
Define
$\Lambda^{t+1}(\theta_{i},\theta_{-i}):=\sum_{j=1}^{i-1}\phi_{j}^{t+1}(\theta_{j})+\sum_{j=i+1}^{D}\phi_{j}^{t}(\theta_{j})+\frac{1}{\lambda+1}\ell(\bm{x};\theta_{i},\theta_{-i}).$
The algorithm is as follows. Initialize $m^{0}\in\mathcal{M}(\Theta),\phi^{0}\in E(m_{0})$. Then for $i\in[D]$, it iteratively updates each $\phi_{i}$,
$\phi_{i}^{t+1}=\mathop{\mathrm{argmax}}_{\phi_{i}\in L^{1}_{0}(m_{i}^{t})}\mathbb{E}_{m_{i}^{t}}\phi_{i}(\theta_{i})+\sum_{j=1}^{i-1}\mathbb{E}_{m_{j}^{t}}\phi_{j}^{t+1}(\theta_{j})+\sum_{j=i+1}^{D}\mathbb{E}_{m_{j}^{t}}\phi_{j}^{t}(\theta_{j})-\mathbb{E}_{m^{t}}\left[\exp(\Lambda^{t+1}(\theta_{i},\theta_{-i}))\right].$
This update yields an explicit formula
$\phi_{i}^{t+1}(\theta_{i})=-\log\mathbb{E}_{m_{-i}^{t}}\exp(\Lambda^{t+1}(\theta_{i},\theta_{-i}))+\eta_{i}^{t},\quad\forall\theta_{i}\in\Theta_{i},$ (2.12)
where
$\eta_{i}^{t}=\begin{cases}\mathbb{E}_{m_{i}^{t}}\log\mathbb{E}_{m_{-i}^{t}}\exp(\Lambda^{t+1}(\theta_{i},\theta_{-i})),\quad\mbox{for }i\leq D-1,\\ 0\quad\mbox{ otherwise. }\end{cases}$
The updated EOT potentials satisfy the identifiability constraints under Eq. 2.10. The solution $\textrm{q}^{*}_{\lambda}$ calculated with these EOT potentials is a valid probability distribution, and its normalizing constant is included in the calculation for $\phi_{D}(\theta_{D})$.
2.2.2 Outer Variational Problem
We now turn to the outer variational problem of Eq. 2.5. After updating the potentials with Eq. 2.12, we optimize over the marginals $m_{i}$’s,
$\min\limits_{m\in\mathcal{M}(\Theta)}\mathbb{E}_{m}\left[\underbrace{(\lambda+1)\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})\right)\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)}_{\text{surrogate loss}}\right]+D_{\text{KL}}(m\parallel\pi).$ (2.13)
Eq. 2.13 is equivalent to a mean-field VI problem with a surrogate log-likelihood. To solve Eq. 2.13, we use a method based on coordinate ascent variational inference (CAVI) [13].
First denote $\Theta_{-i}:=\prod_{j\neq i}\Theta_{j}$, $\theta_{-i}:=(\theta_{[D]\backslash\{i\}})$, and $m_{-i}^{t}:=\prod_{j<i}m_{j}^{t+1}(\theta_{j})\prod_{j>i}m_{j}^{t}(\theta_{j})$. Now define $\eta^{t+1}_{i}(\theta_{i})$ as follows:
$\eta^{t+1}_{i}(\theta_{i}):=-\mathbb{E}_{m_{-i}^{t}}\left[(\lambda+1)\left(\sum_{i=1}^{D}\phi_{i}^{t+1}(\theta_{i})\right)\exp\left(\sum_{i=1}^{D}\phi_{i}^{t+1}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta_{i},\theta_{-i}\right)\right)\right].$ (2.14)
Given marginals $m^{t}=(m_{1}^{t}\cdots m_{D}^{t})$ and potentials $\phi^{t+1}=(\phi_{1}^{t+1}\cdots\phi_{D}^{t+1})\in E(m^{t})$, CAVI iteratively updates each marginal $i\in[D]$ by solving a minimization problem,
$m_{i}^{t+1}=\mathop{\mathrm{argmin}}_{m_{i}\in M(\Theta_{i})}-\mathbb{E}_{m_{i}}[\eta^{t+1}_{i}(\theta_{i})]+D_{\text{KL}}(m_{i}\parallel\pi_{i}).$
This leads to an explicit formula for the update of each marginal,
$m_{i}^{t+1}(\theta_{i})\propto\pi_{i}(\theta_{i})\exp(\eta^{t+1}_{i}(\theta_{i})),\quad\forall\theta_{i}\in\Theta_{i}.$ (2.15)
2.3 Coordinate ascent for $\Xi$-VI
Input: Log-likelihood $\ell\left(\bm{x};\theta\right)$, prior $\pi$, tolerance $\epsilon$, regularization parameter $\lambda$.
Initialize: Marginals $m_{1}^{0},\cdots,m_{D}^{0}$, EOT potentials $\phi_{1}^{0},\cdots,\phi_{D}^{0}$, $t=0$;
while The ELBO has not converged do
       for $i\in[D]$ do
             Update $\phi_{i}^{t+1}(\theta_{i})$ using Eq. 2.12
       end for
      for $i\in[D]$ do
             Update $m_{i}^{t+1}(\theta_{i})$ using Eq. 2.15 // Challenging step
       end for
      Compute $\textrm{q}^{t+1}(\theta)=\exp\left(\sum_{i=1}^{D}\phi_{i}^{t+1}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)\prod_{i=1}^{D}m_{i}^{t+1}(\theta_{i})$;
       Compute $\text{ELBO}(\textrm{q}^{t+1})=\mathbb{E}_{\textrm{q}^{t+1}}\left[\ell(\bm{x};\theta)+\log\pi(\theta)\right]-\mathbb{E}_{\textrm{q}^{t+1}}\left[\log\textrm{q}^{t+1}(\theta)\right]$;
       Increment $t=t+1$;
end while
Output: $\textrm{q}(\theta)$.
Algorithm 1 Coordinate Ascent Algorithm
We have described coordinate updates for the inner and outer variational problems of Eq. 2.5. Using these update, the full coordinate ascent algorithm is in Algorithm 1. It alternates between updating EOT potentials $\phi_{i}$ (Eq. 2.12) and marginals $m_{i}$ (Eq. 2.15). It monitors change in ELBO as the criterion of convergence, which is equivalent (up to a scalar) to the KL divergence between the variational posterior and the exact posterior. It results in an optimal solution of the form
$\textrm{q}^{*}_{\lambda}(\theta)=\exp\left(\sum_{i=1}^{D}\phi^{*}_{\lambda,i}(\theta_{i})+\frac{1}{\lambda+1}\ell(\bm{x};\theta)\right)m^{*}_{\lambda}(\theta),$ (2.16)
where $\phi^{*}_{\lambda},m^{*}_{\lambda}$ are the optimal variational EOT potentials and marginals for given $\lambda$.
Unfortunately, Algorithm 1 is difficult to implement because we cannot calculate the expectations needed in Eq. 2.12 or Eq. 2.14. When we represent $\phi_{i}$’s implicitly, there is no practically stable MFVI for implicit log-likelihood, especially when the model is high-dimensional.
As an approximation, Algorithm 2 is a one-step coordinate ascent algorithm that simplifies the iterative updates to a single round of updates. It uses Eq. 2.15 and Eq. 2.12. In the first step, it computes a set of pseudomarginals $\{\tilde{m}_{i}\}_{i\in[D]}$, and draws samples from them. In the second step, it uses the samples in a multi-marginal Sinkhorn algorithm to compute the optimal EOT coupling.
Input: Likelihood $\ell\left(\bm{x};\theta\right)$, prior $\pi$, tolerance $\epsilon$, regularization parameter $\lambda$.
Initialize: Marginals $m_{1}^{0},\cdots,m_{D}^{0}$, EOT potentials $\phi_{1}^{0},\cdots,\phi_{D}^{0}$;
Compute $\tilde{m}_{1}(\theta_{1}),\cdots,\tilde{m}_{D}(\theta_{D})$ using an approximate inference algorithm.
Compute $\tilde{\phi}_{\lambda,1},\cdots,\tilde{\phi}_{\lambda,D}$ using a multi-marginal Sinkhorn algorithm.
Compute $\tilde{\textrm{q}}_{\lambda}(\theta)=\exp\left(\sum_{i=1}^{D}\tilde{\phi}_{\lambda,i}(\theta_{i})+\frac{1}{\lambda+1}\ell(\bm{x};\theta)\right)\prod_{i=1}^{D}\tilde{m}_{i}(\theta_{i})$;
Output: $\tilde{\textrm{q}}_{\lambda}(\theta)$.
Algorithm 2 One-Step Coordinate Ascent Algorithm
Note, in the first step, we can use any algorithm for approximating the posterior marginals, e.g., variational inference [13], expectation propagation (EP) [45], or MCMC [62]. We recommend choosing an algorithm like EP that is likely to produce an overdispersed approximation. Intuitively, it produces more variation in the initial samples, which leads to better downstream approximations. We demonstrate this empirically in Section 4.
Even the one-step coordinate ascent algorithm can be computationally challenging, particularly due to the complexities of solving for the optimal EOT coupling with a growing number of marginals. Section 2.4 outlines specific conditions for Algorithm 2 to be polynomial-time solvable.
2.4 Implementation and Computational Complexity
Algorithm 2 is a two-stage algorithm to approximate the $\Xi$-variational posterior. A natural question to ask is how tractable is it in high-dimensional models? In the first stage, existing methods for the pseudomarginal computation (such as BBVI) are known to scale well in high dimensions [37]. But the second stage is an EOT computation, which is not necessarily scalable [5]. To this end, we discuss sufficient conditions for the polynomial-time solvability of the Sinkhorn algorithm.
We use a discretized representation to formulate and implement our algorithm. The set of marginals $\{m_{i}\}_{i\in[D]}$ is a nonnegative matrix $\bm{M}=(\bm{M}_{1},\cdots,\bm{M}_{D})\in\mathbb{R}^{M\times D}$, where each $\bm{M}_{i}$ contains $M$ design points; the EOT potentials $\{\phi_{i}\}_{i\in[D]}$ are a matrix $\bm{F}=(\bm{F}_{1},\cdots,\bm{F}_{D})\in\mathbb{R}^{M\times D}$; the negative loss $-\ell(\bm{x}^{n};\theta)$ is a cost tensor $\bm{C}=(C_{i_{1},...,i_{D}})\in(\mathbb{R}^{M})^{\otimes D}$; the variational posterior q is a nonnegative tensor $\bm{Q}=(Q_{i_{1},...,i_{D}})\in(\mathbb{R}^{M})^{\otimes D}$.
2.4.1 The Sinkhorn algorithm for high-dimensional models
Numerically, the EOT problem (2.6) can be posed as a linear programming problem with $DM$ constraints and $M^{D}$ minimizing variables,
$\min\limits_{\bm{Q}>0,r_{i}(\bm{Q})=r_{i}(\bm{M})}\langle\bm{C},\bm{Q}\rangle+(\lambda+1)\langle\log\bm{Q}-\log\bm{M},\bm{Q}\rangle.$
Define $H(\mathbf{Q}):=-\langle\mathbf{Q},\log(\mathbf{Q})-\mathbf{1}_{M\times\ldots\times M}\rangle$. We introduce dual variables $\bm{B}_{1},\cdots,\bm{B}_{D}$ to incorporate marginalization constraints,
$\min\limits_{\bm{Q}>0}\langle\bm{C},\bm{Q}\rangle-(\lambda+1)H(\bm{Q})-\sum_{i=1}^{D}\bm{B}_{i}^{T}(r_{i}(\bm{Q})-r_{i}(\bm{M})).$ (2.17)
Now perform a change of variables $\bm{F}_{i}:=\frac{1}{\lambda+1}\bm{B}_{i}-\log r_{i}(\bm{M}_{i})$. Since the objective in Eq. 2.17 is strongly convex, it has a unique minimizer $\bm{Q}^{*}$.
A direct calculation recovers the structure of the solution,
$\bm{Q}^{*}_{k_{1}\cdots k_{D}}=\frac{\exp\left[\sum_{i=1}^{D}\bm{F}_{ik_{i}}-\frac{1}{\lambda+1}\bm{C}_{k_{1}\cdots k_{D}}\right]\bm{M}_{k_{1}\cdots k_{D}}}{\sum_{1\leq k_{i}\leq M,\forall i\in[D]}\exp\left[\sum_{i=1}^{D}\bm{F}_{ik_{i}}-\frac{1}{\lambda+1}\bm{C}_{k_{1}\cdots k_{D}}\right]\bm{M}_{k_{1}\cdots k_{D}}},$ (2.18)
where each $M$-dimensional vector $\bm{F}_{j}$ corresponds to the $j$th EOT potential.
Plugging Eq. 2.18 into Eq. 2.17 yields the dual problem:
$\bm{F}^{*}=\arg\max\limits_{\bm{F}_{1},\cdots,\bm{F}_{D}}\log\left(\sum_{1\leq k_{i}\leq M,\forall i\in[D]}\exp\left[\sum_{i=1}^{D}\bm{F}_{ik_{i}}-\frac{1}{\lambda+1}\bm{C}_{k_{1}\cdots k_{D}}\right]\bm{M}_{k_{1}\cdots k_{D}}\right)-\sum_{i=1}^{D}\bm{F}_{i}^{T}\bm{M}_{i}.$ (2.19)
Eq. 2.19 can be solved via the Sinkhorn algorithm, which uses the updating formulas in Section 2.2. The Sinkhorn algorithm solves for $\bm{F}^{*}$ by iteratively performing log-sum-exp updates between $\bm{F}_{1},\cdots,\bm{F}_{D}$. See Algorithm 3.
Input: Cost tensor $\bm{C}$, marginals $\bm{M}$, tolerance $\epsilon$, regularization parameter $\lambda$.
Initialize: $\bm{F}_{i}=-\frac{1}{D(\lambda+1)}\mathbf{1}-\log\mathbf{M}_{i}$ for $i\in[D]$;
while $\bm{E}>\epsilon$ do
       Choose a greedy coordinate $j=\mathop{\mathrm{argmin}}_{i\in[D]}\|r_{i}(\bm{Q})-\bm{M}_{i}\|_{1}$.
      for $i\in[D]$ do
             Update $\bm{F}_{i}=\begin{cases}\bm{F}_{i}-\log\left(r_{i}(\bm{Q})\right)+\log(\bm{M}_{i})\quad\mbox{i = j},\\ \bm{F}_{i}\quad\mbox{otherwise}.\end{cases}$;
             Compute $r_{i}(\bm{Q})=\frac{\sum_{1\leq k_{j}\leq M,\forall j\neq i}\exp\left[\sum_{j=1}^{D}\bm{F}_{jk_{j}}-\frac{1}{\lambda+1}\bm{C}_{k_{1}\cdots k_{D}}\right]\bm{M}_{k_{1}\cdots k_{D}}}{\sum_{1\leq k_{j}\leq M,\forall j\in[D]}\exp\left(\sum_{i=1}^{D}\bm{F}_{ik_{i}}-\frac{1}{\lambda+1}\bm{C}_{k_{1}\cdots k_{D}}\right)\bm{M}_{k_{1}\cdots k_{D}}}$;
       end for
      Set $\bm{E}=\sum_{i=1}^{D}\|r_{i}(\bm{Q})-\bm{M}_{i}\|_{1}$;
end while
Output: An $M\times D$ matrix $\bm{F}$.
Algorithm 3 (Multmarginal) Sinkhorn Algorithm
2.4.2 The complexity of the Sinkhorn algorithm
Under the assumption that the cost tensor is uniformly bounded, the best-known Sinkhorn algorithm achieves a complexity bound of $O(D^{3}M^{D}(\lambda+1)^{-2})$ [41]. Unfortunately, the dependence scaling in $D$ cannot be improved in general [38]. The polynomial-time solvability of the Sinkhorn algorithm requires additional assumptions on the cost tensor. [4] shows that if the cost has a graphical structure with bounded treewidth or if the cost tensor is low rank plus sparse, then multimarginal EOT is solvable in $\text{poly}(M,D)$ time. Their result is stated as follows:
Proposition 1 ([5]).
Consider cost tensor $\bm{C}\in(\mathbb{R}^{M})^{\otimes D}$ that satisfies one of the following:
  1. 1.
    $\bm{C}$ has graphical structure with constant junction tree width $\omega$; or
  2. 2.
    $\bm{C}=\bm{R}+\bm{S}$ where $\bm{R}\in(\mathbb{R}^{M})^{\otimes D}$ has constant multilinear rank and $\bm{S}$ has $\text{poly}(M,D)$ sparsity.
Then for any $\lambda\geq 0$, the Algorithm 3 terminates in $\text{poly}(M,D,\bm{C}_{\max}/\epsilon,\frac{1}{\lambda+1})$ time.
Remark.
The bounded treewidth assumption guarantees polynomial-time solvability of the junction tree algorithm [69]. Models that satisfy the bounded treewidth assumption include state space models, topic models, and linear regression models with sparse design.
Remark.
The assumption of low-rank plus sparsity is less used in the Bayesian literature. Loosely speaking, the low-rank assumption requires the true posterior to be a mixture of product distributions. The error factor means that the exact posterior need only match a mixture of product distributions up to a $\text{poly}(M,D)$ sparse remainder.
For general graphs $G$ with bounded treewidth, [25] proposed implementing the Sinkhorn algorithm using the junction tree method. It has the following complexity:
Corollary 1.
Assume the cost tensor $\mathbf{C}\in(\mathbb{R}^{M})^{\otimes D}$ has constant treewidth $\omega$. Consider Algorithm 3 implemented with the junction tree method [25]. For any $\lambda\geq 0$, it converges in $O(D^{3}M^{\omega+1}(\lambda+1)^{-1}\epsilon^{-1})$ iterations.
This result is a modified version of Theorem 4 in [25], which shows the computational complexity drops with increasing $\lambda$. Also observe that achieving polynomial dependence on $D$ only requires the graph’s treewidth $\omega(G)$ to grow slower than $\log(D)$. Consequently, the polynomial-time solvability might be achievable for “locally tree-like” graphs.
Counterintuitively, for $\lambda$ scaling faster than $D$, the computational complexity decreases as $D$ increases. Yet, as Corollary 3 and Corollary 4 show in Section 3, when $\lambda$ scales faster than $D$, the variational posterior effectively reduces to the naive mean-field approximation.
Remark.
Well-known examples exist that violate the conditions in Proposition 1. For example, an Ising model over a complete $D\times D$ graph has a treewidth of $D$, and its cost tensor is neither low rank nor sparse. In fact, implementing the Sinkhorn algorithm for an Ising model on a complete graph is known to be NP-hard [4]. For these problems, we further approximate by grouping the variables and coupling marginals represented by the group. For example, if we have an Ising model with $100$ variables, we could group first $50$ and last $50$ variables and perform an EOT computation with two $50$-dimensional marginals instead of $50$ one-dimensional marginals. This still produces a strict improvement over MFVI, the procedure runs in polynomial time when the group number is fixed since the complexity of the Sinkhorn algorithm is blind to the dimension of each marginal [3].
While the Sinkhorn step determines the polynomial-time solvability of Algorithm 2, the overall solution accuracy depends crucially on the choice of pseudomarginal algorithm. For completeness, we review two algorithms in Appendix B, namely EP and BBVI.
3 Asymptotic Theory
In this section, we derive the asymptotic theory for $\Xi$-VI.
Define $\mathbb{P}_{p}(\Theta):=\{\textrm{q}\in\mathbb{P}(\Theta):\mathbb{E}_{\textrm{q}}[\|\theta\|^{p}]<\infty\}$. For $p\geq 1$, the ($p^{\text{th}}$)-Wasserstein distance is defined as $W_{p}(\textrm{q}_{0},\textrm{q}_{1}):=(\inf_{\textrm{q}\in\mathcal{C}(\textrm{q}_{0},\textrm{q}_{1})}\mathbb{E}_{\textrm{q}}[\|X-Y\|^{p}])^{1/p}$. The space $(\mathbb{P}_{2}(\Theta),W_{2})$ forms a metric space [67]. We denote $\text{BW}(\mathbb{R}^{D})$ as the subspace of $\mathbb{P}_{2}(\mathbb{R}^{D})$ consisting of Gaussian distributions, known as the Bures-Wasserstein space [10]. The metric topology of $\mathbb{P}_{2}(\Theta)$ and $\text{BW}(\mathbb{R}^{D})$ are crucial for our analysis.
We assume that observations $X_{1},\ldots,X_{n}\stackrel{{\scriptstyle iid}}{{\sim}}P_{\theta_{0}}$ for a true parameter $\theta_{0}\in\Theta$, and we assume that the posterior $\textrm{q}^{*}_{0}$ is in the space $(\mathbb{P}_{2}(\Theta),W_{2})$. We also make explicit the dependence on $n$ of the regularizer $\lambda_{n}$ and the data $\bm{x}^{(n)}$. With this setup, the $\Xi$-variational posterior is
$^{*}_{\lambda_{n}}=\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathbb{P}_{2}(\Theta)}\mathbb{E}_{\textrm{q}}\left[-\ell(\bm{x}^{(n)};\theta)\right]+D_{\text{KL}}(\textrm{q}\parallel\pi)+\lambda_{n}\Xi(\textrm{q}).$ q ) . (3.1)
Our goal is to establish the asymptotic properties of $\textrm{q}^{*}_{\lambda_{n}}$. In Section 3.1, we prove consistency and asymptotic normality (Bernstein von-Mises theorem) for $\textrm{q}^{*}_{\lambda_{n}}$ in finite-dimensional settings. In Section 3.2, we establish the asymptotic properties for $\textrm{q}^{*}_{\lambda_{n}}$ in high-dimensional models where $D$ grows with $n$, with an emphasis on explaining the phase transition behavior under different regimes of $\lambda_{n}$. In Section 3.3, we establish the convergence of $\textrm{q}^{*}_{\lambda_{n}}$ when $n$ is fixed and $\lambda_{n}$ tends to $\infty$ or 00, and algorithmic stability when we replace Algorithm 1 with Algorithm 2.
3.1 Asymptotic Normality in Finite Dimension
We establish the posterior consistency and asymptotic normality of $\Xi$-variational posteriors for finite-dimensional models. Specifically, asymptotic normality states that $\Xi$-VI converges in the limit to one of three quantities: the mean-field minimizer of a normal distribution, the normal distribution itself, or a $\Xi$-variational normal approximation. Furthermore, we identify thresholds for a phase transition in the behavior of the variational posterior. Essentially, it depends on the limiting value of $\lambda_{n}$ as the sample size $n$ increases.
For two positive sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ or $a_{n}=O(b_{n})$ or $b_{n}\gtrsim a_{n}$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for all $n$. The constant $C$ does not depend on $n$. The relation $a_{n}\asymp b_{n}$ holds if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ are true. We write $a_{n}\prec b_{n}$ or $a_{n}=o(b_{n})$ if $a_{n}\leq c_{n}b_{n}$ for all $n$, for some sequence $c_{n}$ that converges to zero, $c_{n}\to 0$. We write $a_{n}\succ b_{n}$ if $b_{n}=o(a_{n})$.
We make the following assumptions.
Assumption 1 (Prior Mass).
The prior $\pi$ has Lebesgue-density $\pi(\theta):=\exp(\nu_{0}(\theta))$, where the function $\nu_{0}:\Theta\mapsto\mathbb{R}$ is twice continuously differentiable and bounded in a neighborhood of $\theta_{0}$. For some $C>0$, we have
$\sup_{\|\theta-\theta_{0}\|_{2}\leq Cn^{-1/2}}\|\nabla\nu_{0}(\theta)\|_{2}\lesssim\sqrt{n},\quad\text{and}\quad\sup_{\|\theta-\theta_{0}\|_{2}\leq Cn^{-1/2}}\|\nabla^{2}\nu_{0}(\theta)\|_{2}\lesssim n.$
Assumption 2 (Consistent Testability Assumptions).
For every $\epsilon>0$, there exists a sequence of tests $\phi_{n}$ such that
[…]
Assumption 3 (Local Asymptotic Normality (LAN) Assumptions).
For every compact set $K\subset\mathbb{R}^{D}$, there exists random vectors $\Delta_{n,\theta_{0}}$ bounded in probability and nonsingular matrix $V_{\theta_{0}}$ such that
$\sup_{h\in K}|\ell(\bm{x}^{(n)};\theta_{0}+\delta_{n}h)-\ell(\bm{x}^{(n)};\theta_{0})-h^{T}V_{\theta_{0}}\Delta_{n,\theta_{0}}+\frac{1}{2}h^{T}V_{\theta_{0}}h|\overset{P_{\theta_{0}}}{\to}0,$
where $\delta_{n}$ is a $D\times D$ diagonal matrix. For $D=1$, we commonly have $\delta_{n}=n^{-1/2}$.
The first assumption ensures the prior is light-tailed. It is satisfied by, for example, the flat prior or the sub-Gaussian prior.
The second assumption guarantees the existence of a sequence of uniformly consistent tests for $H_{0}:\theta=\theta_{0}$ versus $H_{1}:\|\theta-\theta_{0}\|_{2}\geq\epsilon$ based on the data. This condition is satisfied when we have a consistent sequence of estimators $T_{n}$ for $\theta$ and set $\phi_{n}(\theta):=I\{T_{n}-\theta\geq\epsilon/2\}$, or when the Hellinger distance between $\{p_{\theta},\|\theta-\theta_{0}\|_{2}\geq\epsilon\}$ and $p_{\theta_{0}}$ is lower bounded by some positive constant $\delta$ [29].
The third assumption says that the likelihood can be asymptotically approximated by the likelihood of a normal location model centered at $\theta_{0}$ after rescaling. The rescaling sequence $\delta_{n}$ determines the optimal convergence rate of the posterior. For finite-dimensional, correctly-specified models, this rate is typically $n^{-1/2}$ [29].
In line with Assumption 3, we implement a change of variable,
$h=\delta_{n}^{-1}(\theta-\theta_{0}-\delta_{n}\Delta_{n,\theta_{0}}).$ (3.2)
Our main result shows that the rescaled $\Xi$-variational posterior satisfies a Bernstein-von-Mises phenomenon with phase transition.
Theorem 2 (Bernstein von-Mises Theorem).
Let $\tilde{\textrm{q}}_{\lambda_{n}}$ be the distribution of the rate-adjusted parameter $h$ defined in Eq. 3.2. The distribution $\tilde{\textrm{q}}_{\lambda_{n}}$ converges in the Wasserstein metric to a normal distribution under the following three regimes:
  1. 1.
    If $\lambda_{n}\to\infty$, then $W_{2}\left(\tilde{\textrm{q}}_{\lambda_{n}},N(0,((V_{\theta_{0}})_{\text{diag}})^{-1})\right)\overset{P_{\theta_{0}}}{\to}0$, where $(V_{\theta_{0}})_{\text{diag}}$ is the diagonal submatrix of $V_{\theta_{0}}$.
  2. 2.
    If $\lambda_{n}\to 0$, then $W_{2}\left(\tilde{\textrm{q}}_{\lambda_{n}},N(0,V_{\theta_{0}}^{-1})\right)\overset{P_{\theta_{0}}}{\to}0$.
  3. 3.
    If $\lim_{n\to\infty}\lambda_{n}=\lambda_{\infty}$ for some $\lambda_{\infty}\in(0,\infty)$, then
    $W_{2}\left(\tilde{\textrm{q}}_{\lambda_{n}},\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathbb{P}_{2}(\mathbb{R}^{D})}D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{\infty}\Xi(\textrm{q})\right)\overset{P_{\theta_{0}}}{\to}0.$
The result aligns well with intuition. When $\lambda_{n}$ approaches infinity, $\textrm{q}^{*}_{\lambda_{n}}$ converges to the mean-field approximation. When $\lambda_{n}$ approaches zero, the constraint set in the Lagrangian dual problem increases to include the true limiting posterior $N(0,V_{\theta_{0}}^{-1})$. When $\lambda_{n}$ converges to some finite value $\lambda_{\infty}$, the variational posterior stabilizes at the $\Xi$-variational posterior to the true limiting posterior. In the regime where $\lim_{n\to\infty}\lambda_{n}$ does not exists but $\lambda_{n}=O(1)$, $\textrm{q}^{*}_{\lambda_{n}}$ converges to a a “biased” estimate of the true Gaussian posterior $N(0,V_{\theta_{0}}^{-1})$ along a subsequence of $\lambda_{n}$ that converges as $n\to\infty$.
The Bernstein von-Mises Theorem implies the (weak) posterior consistency for $\textrm{q}^{*}_{\lambda_{n}}$.
Corollary 2.
The $\Xi$-variational posterior is consistent in $[P_{\theta_{0}}]$-probability, i.e. $W_{2}(\textrm{q}^{*}_{\lambda_{n}},\delta_{\theta_{0}})\overset{P_{\theta_{0}}}{\to}0$.
The convergence in Corollary 2 is stated in the Wasserstein metric, which is slightly stronger than the typical metric used in posterior consistency results. The convergence in the Wasserstein metric is equivalent to weak convergence plus the convergence of the second moments (Theorem 5.11, [63]). Thus, posterior consistency and the Bernstein–von Mises theorem (Theorem 2) can be framed in terms of the weak convergence and $L_{2}$ convergence for the corresponding measures.
3.2 Asymptotic Results with Growing Dimension
In this section, we explore the asymptotic limit of $\textrm{q}^{*}_{\lambda_{n}}$ in the settings where $D$ increases with $n$. Specifically, we characterize regimes of $\lambda_{n}$ where $\textrm{q}^{*}_{\lambda_{n}}$ approximates either a mean-field or the exact posterior. We begin with a theorem for the general model and then focus on the high-dimensional linear regression model. We assume that the parameter space is a unit cube, i.e. $\Theta=[-1,1]^{D}$, but the results are extendable to any compact set in $\mathbb{R}^{D}$. Two assumptions underpin our analysis:
Assumption 4 (Prior Assumptions).
The prior is a product distribution given as $\pi(\theta)=\exp(\nu_{0}(\theta))$, where the function $\nu_{0}:\Theta\mapsto\mathbb{R}$ is twice continuously differentiable and bounded in a neighborhood of $\theta_{0}$.
Assumption 5 (Model Assumptions).
The log-likelihood function $\ell(\bm{x}^{(n)};\cdot)$ is uniformly bounded over $\Theta$ and twice continuously differentiable. Moreover, the gradient $\nabla\ell(\bm{x}^{(n)};\cdot)$ and Hessian $\nabla^{2}\ell(\bm{x}^{(n)};\cdot)$ are uniformly bounded over $\Theta$.
Assumption 4 is a prior smoothness assumption. Assumption 5 guarantees the smoothness of the log-likelihood function by requiring uniform boundedness of the function value, its gradient, and Hessian. Both assumptions are reasonable given the compactness of $\Theta$.
General Case.
When $\ell(\bm{x}^{(n)};\cdot)$ is twice continuously differentiable, we define its $i^{\text{th}}$ partial differential $\theta\mapsto\partial_{i}\ell(\bm{x}^{(n)};\theta)$ as
$\partial_{i}\ell(\bm{x}^{(n)};\theta):=\frac{\partial}{\partial\theta_{i}}\ell(\bm{x}^{(n)};\theta)=[\nabla\ell(\bm{x}^{(n)};\theta)]_{i},$
and $(i,j)^{\text{th}}$ second partial differential $\theta\mapsto\partial_{ij}\ell(\bm{x}^{(n)};\theta)$ as
$\partial_{ij}\ell(\bm{x}^{(n)};\theta):=\frac{\partial}{\partial\theta_{i}}\left(\frac{\partial}{\partial\theta_{j}}\ell(\bm{x}^{(n)};\theta)\right)=[\nabla^{2}\ell(\bm{x}^{(n)};\theta)]_{ij}.$
We quantify the fluctuation of the log-likelihood function using the concept of oscillation, defined as $\omega_{\Theta}(f):=\sup_{\Theta}f(\theta)-\inf_{\Theta}f(\theta)$ for $f:\Theta\mapsto\mathbb{R}$. We now state the main result:
Theorem 3.
Let Assumption 4 and Assumption 5 hold. Let $\textrm{q}^{*}_{\lambda_{n}}$ be an optimizer of Eq. 3.1. Define:
$a:=\omega_{\Theta}(\ell(\bm{x}^{(n)};\theta)),\quad b_{i}:=\omega_{\Theta}(\partial_{i}\ell(\bm{x}^{(n)};\theta)),\quad c_{ij}:=\begin{cases}\omega_{\Theta}(\partial_{ij}\ell(\bm{x}^{(n)};\theta))&\text{for }i=j,\\ \sup_{\theta\in\Theta}|\partial_{ij}\ell(\bm{x}^{(n)};\theta)|&\text{for }i\neq j.\end{cases}$ (3.3)
When $\lambda_{n}\succ D^{-1/2}\max\left(\sqrt{a\sum_{i=1}^{D}c_{ii}},\sqrt{\sum_{i=1}^{D}b_{i}^{2}},\sqrt{\sum_{i=1}^{D}\sum_{j=1}^{D}c_{ij}^{2}},D^{1/2}\right)$, there exists a sequence of product distributions $m_{\lambda_{n}}^{*}$ such as, for any $1$-Lipschitz function $\psi:\mathbb{R}\mapsto\mathbb{R}$, as $n\to\infty$,
$\sup_{\bm{x}^{(n)}\in\mathbb{X}^{n}}\left|\frac{1}{D}\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{m_{\lambda_{n}}^{*}}[\psi(\theta_{i})]\right)\right|\overset{P_{\theta_{0}}}{\to}0.$ (3.4)
When $\lambda_{n}\prec D\Xi^{-1}(\textrm{q}^{*}_{0})$, for any $1$-Lipschitz function $\psi:\mathbb{R}\mapsto\mathbb{R}$, as $n\to\infty$,
$\sup_{\bm{x}^{(n)}\in\mathbb{X}^{n}}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})-\frac{1}{D}\sum_{i=1}^{D}\mathbb{E}_{\textrm{q}^{*}_{0}}[\psi(\theta_{i})]\right)^{2}\right]\overset{P_{\theta_{0}}}{\to}0,$ (3.5)
The proof uses the theory nonlinear large deviation [75] and the properties of displacement convex functionals [6]. See Proofs of Section 3.2.
Eq. 3.4 defines a mean-field regime, where a product distribution produces indistinguishable $1$-Lipschitz statistics (first-order statistics) as $\textrm{q}^{*}_{\lambda_{n}}$ asymptotically. This regime characterizes when $\Xi$-VI can be replaced by MFVI. While the threshold for $\lambda_{n}$ consists of four terms, the critical term is $\sqrt{\sum_{i,j}c_{ij}^{2}}$ and the other terms are typically well controlled. Roughly, this means we can determine the equivalence between $\Xi$-VI and MFVI by comparing $\lambda_{n}$ to $(D^{-1/2})$-rescaled Frobinius norm of the Fisher information.
Eq. 3.5 defines a Bayes optimal regime, where $\Xi$-VI asymptotically recovers $1$-Lipschitz statistic of the exact posterior. When the dimension $D$ remains constant as $n$ increases, $\Xi(\textrm{q}^{*}_{0})$ converges to zero due to posterior consistency, thus any bounded sequence of $\lambda_{n}$ provides Bayes optimality. When $D$ grows with $n$ but at a slow rate (e.g. $D\lesssim n^{-1/3}$), we may still expect a form of posterior consistency to hold the Bayesian optimal regime to contain non-trivial choices of $\lambda_{n}$.
To match the computational complexity in Section 2.4, we provide sufficient conditions for $\lambda_{n}\succ D$ to be contained in the mean-field regime.
Corollary 3.
Under the setting of Theorem 3, if $a\lesssim D,b_{i}\lesssim D,c_{ii}\lesssim D$ for $i\in[D]$ and $c_{ij}\lesssim 1$ for $i\neq j$, then for $\lambda_{n}\succ D$, there exists a product distribution $m_{\lambda_{n}}^{*}$ satisfying, for any $1$-Lipschitz function $\psi:\mathbb{R}\mapsto\mathbb{R}$, as $n\to\infty$,
$\sup_{\bm{x}^{(n)}\in\mathbb{X}^{n}}\left|\frac{1}{D}\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{m_{\lambda_{n}}^{*}}[\psi(\theta_{i})]\right)\right|\overset{P_{\theta_{0}}}{\to}0.$ (3.6)
The assumptions of Corollary 3 are met, for example, when the model satisfies Assumption 5 with uniformly bounded gradient and Hessian of the log-likelihood. Alternatively, it is also sufficient when 1) the gradient and diagonal Hessian of the log-likelihood scale slower than $D$ entry-wise and 2) the off-diagonal Hessian is uniformly bounded.
The thresholds of Theorem 3 could be too abstract to be practically useful. The definition of mean-field and Bayes optimal regimes are also somewhat restrictive because they only show equivalence in terms of $1$-Lipschitz statistics. To refine our result, we focus on the case of high-dimensional linear regression models.
High-Dimensional Linear Model.
We observe $\{(x_{i},y_{i}):1\leq i\leq n\}$, $y_{i}\in\mathbb{R}$, $x_{i}\in\mathbb{R}^{D}$. Let $\bm{y}=[y_{1},\cdots,y_{n}]\in\mathbb{R}^{n}$ and $\bm{X}=[x_{1},\ldots,x_{n}]^{T}\in\mathbb{R}^{n\times D}$. We consider a high-dimensional Bayesian linear regression model where both $n,D$ are tending to infinity:
$\bm{y}=\bm{X}\theta+\epsilon,\quad\epsilon\sim N(0,\sigma^{2}I_{n}),\quad\theta\sim\pi.$ (3.7)
We assume that prior $\pi$ is supported on $\Theta$. Moreover, we introduce some shorthand notations:
Definition 1.
For matrix $\bm{B}:=\frac{1}{\sigma^{2}}\bm{X}^{T}\bm{X}$, we denote the diagonal and off-diagonal submatrix of $\bm{B}$ as $\bm{B}_{\text{diag}}$ and $\bm{B}_{\text{off}}$, respectively. Define $\bm{B}_{\text{diag}}^{\prime}$ as diagonal matrix with $[\bm{B}_{\text{diag}}^{\prime}]_{ii}:=1/([\bm{B}]^{-1})_{ii}$.
The $\Xi$-VI for Bayesian linear model is given by
$^{*}_{\lambda_{n}}=\arg\min_{\textrm{q}\in\mathbb{P}_{2}(\Theta)}\mathbb{E}_{\textrm{q}}\left[\frac{\|\bm{y}-\bm{X}\theta\|^{2}}{2\sigma^{2}}\right]+D_{\text{KL}}(\textrm{q}\parallel\pi)+\lambda_{n}\Xi(\textrm{q}).$ q ) . (3.8)
In addition to Assumption 4 and Assumption 5, we posit a model curvature assumption.
Assumption 6 (Curvature Assumption).
The eigenvalues of $\bm{X}^{T}\bm{X}$ is lower bounded by a constant $\kappa_{2}>0$. Moreover, the prior is of the form $\pi=\exp(\nu_{0})$ where $\nu_{0}$ is $\kappa_{1}$-concave for $\kappa_{1}\geq 0$.
The assumption implies $D\leq n$, and the exact posterior $\textrm{q}^{*}_{0}$ is $(\kappa_{2}+\kappa_{1})$-log concave. It is satisfied by a wide range of design matrix $\bm{X}$ and the prior choices including Gaussian prior or Laplace prior with a positive scale parameter.
Our main result characterizes the asymptotic regimes of $\textrm{q}^{*}_{\lambda_{n}}$ for model (3.7).
Theorem 4.
Let Assumption 4Assumption 5,and Assumption 6 hold. Let $\textrm{q}^{*}_{\lambda_{n}}$ be an optimizer of Eq. 3.1. Then the following holds:
When $\lambda_{n}\succ\sqrt{\text{tr}(\bm{B}_{\text{off}}^{2})}$, there exists a sequence of product distributions $m_{\lambda_{n}}^{*}$ such as, as $n\to\infty$,
$\sup_{\bm{x}^{(n)}\in\mathbb{R}^{n}}W_{2}(\textrm{q}^{*}_{\lambda_{n}},m_{\lambda_{n}}^{*})\overset{P_{\theta_{0}}}{\to}0.$ (3.9)
When $\lambda_{n}\succ\sqrt{D^{-1}\text{tr}(\bm{B}_{\text{off}}^{2})}$, there exists a sequence of product distributions $m_{\lambda_{n}}^{*}$ such as, for any $1$-Lipschitz function $\psi:\mathbb{R}\mapsto\mathbb{R}$, as $n\to\infty$,
$\sup_{\bm{x}^{(n)}\in\mathbb{R}^{n}}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})-\frac{1}{D}\sum_{i=1}^{D}\mathbb{E}_{m_{\lambda_{n}}^{*}}[\psi(\theta_{i})]\right)^{2}\right]\overset{P_{\theta_{0}}}{\to}0.$ (3.10)
When $\lambda_{n}\prec\left[\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))\right]^{-1}$, as $n\to\infty$,
$\sup_{\bm{x}^{(n)}\in\mathbb{R}^{n}}W_{2}(\textrm{q}^{*}_{\lambda_{n}},\textrm{q}^{*}_{0})\overset{P_{\theta_{0}}}{\to}0.$ (3.11)
When $\lambda_{n}\prec D\left[\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))\right]^{-1}$, for any $1$-Lipschitz function $\psi:\mathbb{R}\mapsto\mathbb{R}$, as $n\to\infty$,
$\sup_{\bm{x}^{(n)}\in\mathbb{R}^{n}}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})-\frac{1}{D}\sum_{i=1}^{D}\mathbb{E}_{\textrm{q}^{*}_{0}}[\psi(\theta_{i})]\right)^{2}\right]\overset{P_{\theta_{0}}}{\to}0.$ (3.12)
Eq. 3.9 and Eq. 3.10 form the mean-field regimes, which refine the regime of Theorem 3. When $\lambda_{n}$ scales faster than $\sqrt{\text{tr}(\bm{B}_{\text{off}}^{2})}$, $\textrm{q}^{*}_{\lambda_{n}}$ converges in Wasserstein metric to a product distribution. This means all moment statistics can be asymptotically transported between $\Xi$-VI and MFVI. When $\lambda_{n}$ scales faster than $\sqrt{\text{tr}(\bm{B}_{\text{off}}^{2})/D}$, we can transport any $1$-Lipschitz statistic between $\Xi$-VI and MFVI. As $\lambda_{n}$ increases, $\textrm{q}^{*}_{\lambda_{n}}$ shares more distributional information with $m_{\lambda_{n}}^{*}$.
Eq. 3.11 and Eq. 3.12 define the Bayes optimal regimes. When $\lambda_{n}$ increases slower than $\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\cdot\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))^{-1}$, $\textrm{q}^{*}_{\lambda_{n}}$ converges to the exact posterior in the Wasserstein metric. By relaxing a factor of $D$, $\textrm{q}^{*}_{\lambda_{n}}$ achieves asymptotic Bayes optimality for all $1$-Lipschitz statistics. The first term, $\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}$, measures the discrepancy between the interaction matrices used in the exact and mean-field posterior inference [47]. The second term, $\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))^{-1}$, measures the size of the true posterior covariance. For sufficiently regular models, we could use the Bernstein von-Mises results $D\lesssim n^{1/3}$ [53] or $D\lesssim n^{1/2}$ [35] to approximate the posterior covariance with the inverse Fisher information matrix, namely $(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))^{-1}\approx\text{tr}(\bm{B})$. Thus, when the dimension $D$ grows slowly with $n$, the Bayes optimal regimes in Eq. 3.11 and Eq. 3.12 correspond to $\lambda_{n}\prec\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\bm{B})$ and $\lambda_{n}\prec D\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\bm{B})$, which are determined by $\bm{B}$.
It is desirable to find an overlap between the Bayes optimal and mean-field regimes, as the $\textrm{q}^{*}_{\lambda_{n}}$ in this regime could be computed efficiently with MFVI while closely approximating the exact posterior. Such a regime would depend on the target of posterior inference. For $1$-Lipschitz statistics, the overlap regime is characterized by
$\sqrt{\text{tr}(\bm{B}_{\text{off}}^{2})}\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))\prec D^{3/2}.$ (3.13)
This criterion is satisfied, for example, when $\text{tr}(\bm{B}_{\text{off}}^{2})\prec D,\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}\lesssim 1$ and $\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))\lesssim D$, which recovers the Bayes optimal condition for MFVI [47, 48]. But our criterion is more flexible: for example, it is also satisfied $\text{tr}(\bm{B}_{\text{off}}^{2})\lesssim D,\|\bm{B}-\bm{B}_{\text{diag}}^{\prime}\|_{2}\lesssim 1$ and $\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))\prec D$.
When no choice of $\lambda_{n}$ satisfies the overlap criterion, there is a gap between the mean-field and Bayes optimal regimes, indicating that the mean-field approximation fails to produce accurate posterior inference. Achieving accurate posterior inference thus requires paying an additional computational cost that scales inversely with $\lambda_{n}$, as discussed in Section 2.4.
Let the eigenvalues of $\bm{B}_{\text{off}}$ as $\eta_{D}\geq\ldots\geq\eta_{1}$. Then $\text{tr}(\bm{B}_{\text{off}}^{2})=\sum_{i=1}^{D}\eta_{i}^{2}$, and the mean-field regime Eq. 3.9 corresponds to $\lambda_{n}\succ\sqrt{\sum_{i=1}^{D}\eta_{i}^{2}}$. To match the complexity bound in Section 2.4, we provide sufficient conditions for $\lambda_{n}\succ D$ to be in the mean-field regime Eq. 3.9.
Corollary 4.
Let $\eta_{D}\geq\ldots\geq\eta_{1}$ denote the eigenvalues of $\bm{B}_{\text{off}}$. If $\sum_{i=1}^{D}\eta_{i}^{2}\lesssim D^{2}$ as $n\to\infty$, then for $\lambda_{n}\succ D$, there exists $m_{\lambda_{n}}^{*}\in\mathcal{M}(\Theta)$ such that $\sup_{\bm{x}^{(n)}\in\mathbb{R}^{n}}W_{2}(\textrm{q}^{*}_{\lambda_{n}},m_{\lambda_{n}}^{*})\overset{P_{\theta_{0}}}{\to}0$.
Data preprocessing often involves normalizing features to have unit variances. Thus, the requirement that $\sum_{i=1}^{D}\eta_{i}^{2}\lesssim D^{2}$ is often met in practice.
3.3 Finite-Sample Convergence
In this section, we establish convergence results for $\textrm{q}^{*}_{\lambda}$ when $\lambda$ tends to 00 or $\infty$, while keeping $n$ fixed [^1] We also drop the dependence of $\lambda$ on $n$ because $n$ is fixed.. Understanding this setting helps with computational considerations, as it justifies substituting the marginals of $\textrm{q}^{*}_{\lambda}$ with a set of pseudomarginals. Moreover, we establish the convergence behaviors of the $\Xi$-variational posterior for both large and small $\lambda$ values. These results are pertinent for Bayesians, since the classical Bayesian perspective treats the observed data as known [8].
We present three main results: we establish convergence of $\Xi$-variational posterior to the mean-field variational posterior as $\lambda$ tends to infinity, and convergence to the exact posterior as $\lambda$ tends to zero. Based on these results, we prove stability results for $\textrm{q}^{*}_{\lambda}$ when we replace its marginals with another set of marginals, which is used to justify Algorithm 2.
Let us define a cost function $C_{\lambda}$ over $\lambda\in\bar{\mathbb{R}}_{+}$ as $\mathcal{C}_{\lambda}:=\max\limits_{\textrm{q}\in\mathbb{P}_{2}(\Theta)}\text{ELBO}(\textrm{q})-\lambda\Xi(\textrm{q})$.
Limits as $\lambda\to\infty$ or $\lambda\to 0$.
We start with the convergence of $\textrm{q}^{*}_{\lambda}$ and $\mathcal{C}_{\lambda}$ as $\lambda$ tends to infinity.
Theorem 5.
Assume that $D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})<\infty$ for some $\textrm{q}\in\mathcal{M}(\Theta)$. For each $\lambda\in\bar{\mathbb{R}}_{+}$, define the set $\mathbb{Q}_{\lambda}$ as the set of minimizers for the functional $\mapsto D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})+\lambda\Xi(\textrm{q})$ q ) in $\mathbb{P}_{2}(\Theta)$. If $\textrm{q}^{*}_{\infty}\in\mathbb{Q}_{\infty}$, there exists a sequence $\textrm{q}^{*}_{\lambda}\in\mathbb{Q}_{\lambda}$ such that $\lim_{\lambda\to\infty}W_{2}(\textrm{q}^{*}_{\infty},\textrm{q}^{*}_{\lambda})=0$. Furthermore, the $\Xi$-VI cost converges to the mean-field ELBO, i.e. $\lim_{\lambda\to\infty}\left|\mathcal{C}_{\lambda}-\mathcal{C}_{\infty}\right|=0$.
The result shows that every mean-field variational posterior is an accumulation point of some sequence of $\Xi$-variational posteriors. This type of result is called ”large-time limits” in the optimal transport literature. It is possible to prove an exponential rate of convergence for $C_{\lambda}$ under more restrictive conditions and when the likelihood is quadratic [20]. However, this setting is uninteresting for Bayesian inference and we do not pursue it in this paper.
As $\lambda$ tends to zero, we provide analog results to show that $\Xi$-variational posterior converges to the exact posterior in the Wasserstein metric.
Theorem 6.
Assume that $\Xi(\textrm{q}^{*}_{0})<\infty$ $[P_{\theta_{0}}]$-almost surely. For each $\lambda\in\bar{\mathbb{R}}_{+}$, define the set $\mathbb{Q}_{\lambda}$ as the set of minimizers for the functional $\mapsto D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})+\lambda\Xi(\textrm{q})$ q ) in $\mathbb{P}_{2}(\Theta)$. If $\textrm{q}^{*}_{\lambda}\in\mathbb{Q}_{\lambda}$ converges as $\lambda\to 0$ in the Wasserstein metric, then $\lim_{\lambda\to 0}W_{2}(\textrm{q}^{*}_{0},\textrm{q}^{*}_{\lambda})=0$. Furthermore, the $\Xi$-VI cost converges to the true posterior ELBO, i.e. $\lim_{\lambda\to 0}\left|\mathcal{C}_{\lambda}-\mathcal{C}_{0}\right|=0$.
Algorithmic Stability.
Let $m^{*}_{\lambda}$ be the product of marginals of $\textrm{q}^{*}_{\lambda}$. In Section 2.2, we produce to replace idealized Algorithm 1 with a simple, efficient approximate Algorithm 2. A natural question to ask is whether the solution is stable after we replace $m^{*}_{\lambda}$ with pseudomarginals $\tilde{m}$. To answer this question, we leverage the tools of quantitative stability from OT theory [24]. We make two assumptions: a Lipschitz cost assumption, and transportation cost inequality.
Assumption 7 (Lipschitz Cost Assumption).
We assume that there exists a constant $L\geq 0$ and $\phi_{i}:\Theta_{i}\to\mathbb{R}$ such that
[…]
for all $\textrm{q}\in\mathcal{C}(m^{*}_{\lambda})$ and $\tilde{\textrm{q}}\in\mathcal{C}(\tilde{m})$.
The Lipschitz constant in Assumption 7 $L$ can depend implicitly on $n$. Moreover, this assumption is more general than the Lipschitzness of $\ell(\bm{x}^{(n)};\cdot)$ minus some additive correction factors. One sufficient condition is the existence of $\phi_{i}:\Theta_{i}\to\mathbb{R}$ such that $\ell(\bm{x}^{(n)};\theta)-\sum_{i=1}^{D}\phi_{i}(\theta_{i})=f(\theta)g(\theta)$ where $f,g$ are Lipschitz functions of $\theta$. Then Assumption 7 holds with constant $L$ that depends on the Lipschitz constants of $f,g$, the second moments of $m^{*}_{\lambda}$ and $\textrm{q}^{*}_{\infty}$, and $n$. For example, the Gaussian likelihood satisfies Assumption 7 (Lemma 3.5, [24]).
Assumption 8 (Transportion Cost Inequality).
A product distribution $m$ over $\Theta$ satisfies the transportation cost inequality if there exists a constant $C$ such that
$W_{2}(\textrm{q}_{1},\textrm{q}_{2})\leq CD_{\text{KL}}(\textrm{q}_{1}\parallel\textrm{q}_{2})^{1/2},\quad\text{for all }\textrm{q}_{1},\textrm{q}_{2}\in\mathcal{C}(m).$
Assumption 8 is a standard assumption in high dimensional statistics [68]. When $\Theta$ is bounded, the assumption is automatically satisfied by Pinsker’s inequality. Otherwise, this assumption holds when each marginal has a finite exponential moment.
We now state the main stability theorem. The result upper bounds the approximation error of Algorithm 2 in terms of the approximation error of the pseudomarginals.
Theorem 7 (Stability of Algorithm 2).
Let Assumption 7 hold with a Lipschitz constant $L$. Let $m^{*}_{\lambda}$ be the optimal marginals in Eq. 2.16 and $\tilde{m}\in\mathcal{M}\left(\Theta\right)$ be another product distribution. Let $m^{*}_{\lambda}$ satisfy Assumption 8 with a fixed constant $C$. Then for the one-step approximation $\tilde{\textrm{q}}_{\lambda}$ defined in Algorithm 2 with pseudomarginals $\tilde{m}$, the following upper bound holds:
$W_{2}(\textrm{q}^{*}_{\lambda},\tilde{\textrm{q}}_{\lambda})\leq W_{2}(m^{*}_{\lambda},\tilde{m})+2CL^{\frac{1}{4}}W_{2}^{\frac{1}{4}}(m^{*}_{\lambda},\tilde{m}).$ (3.15)
The proof uses an OT technique called shadowing. See Proofs of Section 3.3 for details.
The result highlights the stability of Algorithm 2. Specifically, the approximation error of $\textrm{q}^{*}_{\lambda}$ is only Lipschitz in the approximation error of the pseudomarginals. If $\tilde{m}$ is close enough to $m^{*}_{\lambda}$ in terms of the $W_{2}$ metric, the output of Algorithm Algorithm 2 is guaranteed to well approximate true variational posterior $\textrm{q}^{*}_{\lambda}$.
The setting of Theorem 7 is satisfied by a broad class of models, including the Gaussian linear model with compactly supported prior for the coefficients.
Finally, we present a result to guide the choice of pseudomarginals.
Corollary 5.
Assume Assumption 7 with Lipschitz constant $L$, and Assumption 8 for the pseudomarginals. Then the following limits hold:
  1. 1.
    Let $\textrm{q}^{*^{(\infty)}}_{\lambda}$ be the optimizer of Eq. 2.6 with marginals $\{\textrm{q}^{*}_{\infty,i}\}_{i\in[D]}$. Then $\lim_{\lambda\to\infty}W_{2}(\textrm{q}^{*^{(\infty)}}_{\lambda},\textrm{q}^{*}_{\lambda})=0$.
  2. 2.
    Let $\textrm{q}^{*^{(0)}}_{\lambda}$ be the optimizer of Eq. 2.6 with marginals $\{\textrm{q}^{*}_{0,i}\}_{i\in[D]}$. Then $\lim_{\lambda\to 0}W_{2}(\textrm{q}^{*^{(0)}}_{\lambda},\textrm{q}^{*}_{\lambda})=0$.
The Corollary is a consequence of Theorem 5, Theorem 6, and Theorem 7. As $\lambda$ tends to 00 or $\infty$, the error of replacing the idealized Algorithm 1 with Algorithm 2 vanishes when we use exact posterior marginals or mean-field variational posteriors, respectively. If we plug in a consistent estimate of the exact posterior marginals (e.g. TAP approximation of a linear model with i.i.d. Gaussian design [16]), then Algorithm 2 asymptotically recovers the exact posterior as $\lambda$ tends to zero.
4 Examples
We apply $\Xi$-VI to three statistical models: a multivariate Gaussian , a high-dimensional Bayesian linear regression, and a hierarchical Bayesian model on the eight schools data ([27], Section 5.5). The multivariate Gaussian illustrates the limitations of mean-field VI in estimating the posterior covariance [13], and how $\Xi$-VI can improve on it. In high-dimensional Bayesian linear regression, mean-field approximations accurately produce valid inference under weak covariate interactions [47, 48], but fail with stronger interactions [56, 16] and again $\Xi$-VI improves on the classical approach. Finally, the analysis of the Bayesian hierarchical model shows how $\Xi$-VI provides more accurate posterior inferences on a real-world dataset.
4.1 Multivariate Gaussian distributions
We first apply $\Xi$-VI to approximating a multivariate Gaussian with a Gaussian variational family. It is well known that mean-field VI underestimates the marginal covariance of a multivariate Gaussian distribution [13]. Here we show that $\Xi$-VI interpolates between the mean-field and the target posterior, and strictly outperforms mean-field VI in estimating the covariance.
Assume that the exact posterior is a multivariate normal, $\textrm{q}_{0}^{*}:=N(\mu_{0},\Sigma_{0})$ with $D$-dimensional mean vector $\mu_{0}$ and a $D\times D$ full-rank covariance matrix $\Sigma_{0}$.
We posit a $D$-dimensional Gaussian variational family, $\textrm{q}^{*}_{\lambda}=N(\mu,\Sigma)$, equivalent to the Bures-Wasserstein space defined in Section 3. The $\Xi$-VI formulation is
$^{*}_{\lambda}=\mathop{\mathrm{argmin}}_{\textrm{q}\in\text{BW}(\mathbb{R}^{D})}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{0}^{*})+\lambda\Xi(\textrm{q}).$ q ) . (4.1)
Let $f(\mu,\Sigma)$ be the objective function Eq. 4.1 parameterized by the variational mean $\mu$ and covariance $\Sigma$. A direct calculation shows
$f(\mu,\Sigma)=\frac{1}{2}\left[(\mu_{0}-\mu)^{T}\Sigma_{0}^{-1}(\mu_{0}-\mu)+\lambda\sum_{K=1}^{D}\log\Sigma_{KK}+\text{tr}\left\{\Sigma_{0}^{-1}\Sigma\right\}-(\lambda+1)\log|\Sigma|\right].$
First, we confirm that the optimal $\mu^{*}$ is equal to the true $\mu_{0}$,
$\frac{\partial}{\partial\mu}f(\mu,\Sigma)=0\implies\mu^{*}=\mu_{0}.$
Now we turn to the covariance or precision. Let the true precision matrix be $\Lambda_{0}$. The optimal variational precision matrix $\Lambda^{*}$ is defined by a fixed point equation,
$\frac{\partial}{\partial\Sigma}f(\mu,\Sigma)=0\implies\Lambda^{*}=\frac{1}{\lambda+1}\Lambda_{0}+\frac{\lambda}{\lambda+1}(\Sigma_{\text{diag}}^{*})^{-1},$ (4.2)
where $\Sigma_{\text{diag}}^{*},\Sigma_{\text{off}}^{*}$ denote the diagonal and off-diagonal minor of $\Sigma^{*}$, respectively.
Eq. 4.2 shows that $\Lambda^{*}$ is a convex combination of the true precision $\Lambda_{0}$ and the inverse of the variational marginal variances. The off-diagonal terms of $\Lambda^{*}$ are $\frac{1}{\lambda+1}$ times the off-diagonal components of $\Lambda_{0}$, thus the conditional covariance in the variation posterior is exactly down-weighted by a factor of $\lambda+1$. As the regularizer $\lambda\to\infty$, the off-diagonal elements of $\Lambda^{*}$ converge to 00 while the diagonal elements approach those of $\Lambda_{0}$.
The next result refines this result with upper and lower bounds:
Proposition 2.
[…]
$\left\|\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{\lambda}{\lambda+1}\Sigma_{0,\text{diag}}^{-1}\right]^{-1}-\Sigma_{0}\right\|\leq\|\Sigma^{*}-\Sigma_{0}\|\leq\left\|\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{\lambda}{\lambda+1}\Lambda_{0,\text{diag}}\right]^{-1}-\Sigma_{0}\right\|.$
The weight $\lambda$ controls the approximation error of a variational posterior covariance by balancing the true posterior marginal variances and the mean-field precision with weights $\frac{1}{\lambda+1}$ and $\frac{\lambda}{\lambda+1}$, respectively. For any $\lambda<\infty$, the $\Xi$-variational posterior offers a tighter approximation than the naive mean field.
As a concrete demonstration of these ideas, we study a bivariate Gaussian posterior. Here the $\Xi$-variational posterior has an analytical solution.
Proposition 3.
Consider bivariate Gaussian distribution with precision matrix $\Lambda_{0}=\begin{pmatrix}a_{0}&b_{0}\\ b_{0}&c_{0}\end{pmatrix}$. Then the $\Xi$-variational posterior has the following precision and covariance matrices:
$\Lambda^{*}=\begin{pmatrix}\frac{a_{0}}{2}+\sqrt{\frac{a_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{a_{0}b_{0}^{2}}{c_{0}}}&\frac{1}{\lambda+1}b_{0}\\ \frac{1}{\lambda+1}b_{0}&\frac{c_{0}}{2}+\sqrt{\frac{c_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{c_{0}b_{0}^{2}}{a_{0}}}\end{pmatrix},\quad\Sigma^{*}=\frac{1}{a_{0}c_{0}-\psi^{-1}(\lambda)b_{0}^{2}}\begin{pmatrix}c_{0}&-\frac{b_{0}}{\psi(\lambda)}\\ -\frac{b_{0}}{\psi(\lambda)}&a_{0}\end{pmatrix},$
where $\psi(\lambda)=\frac{1}{2}\left(\lambda+1+\sqrt{\left(\lambda-\frac{2b_{0}^{2}-a_{0}c_{0}}{a_{0}c_{0}}\right)^{2}+\frac{4b_{0}^{2}}{a_{0}^{2}c_{0}^{2}}\left(a_{0}c_{0}-b_{0}^{2}\right)}\right)$.
Compared to the exact covariance, the variational covariance matrix is adjusted by a factor depending on the regularizer $\lambda$. The adjusting function $\psi:[0,\infty)\mapsto[0,\infty)$ is strictly increasing. Thus, as $\lambda$ increases, we have element-wise strictly decreasing convergence to the mean-field covariance, i.e. $\lim_{\lambda\to\infty}\Sigma^{*}=\begin{pmatrix}a_{0}^{-1}&0\\ 0&c_{0}^{-1}\end{pmatrix}$.
Figure 1 demonstrates this analysis. It shows $\hat{\textrm{q}}_{\lambda}$ fitted to a bivariate Gaussian, for different values of $\lambda$. The left panel shows $\hat{\textrm{q}}_{\lambda}$ as a smooth interpolation between the true posterior and the mean-field variational posterior. The right panel paints a quantitative picture of this interpolation: when $\lambda\leq 10^{-1}$, the $\Xi$-variational posterior closely approximates the covariance values of the exact bivariate Gaussian posterior. For $\lambda\geq 10^{1}$, the covariance is close to zero, which indicates proximity to the mean-field variational posterior. Both plots suggest that $\hat{\textrm{q}}_{\lambda}$ undergoes a ”phase transition” phenomenon at $\lambda\in[10^{-1},10^{1}]$.
Figure 1: $\Xi$-VI for a bivariate Gaussian posterior for varying $\lambda$. The left panel illustrates the transition of the variational posterior $\textrm{q}^{*}_{\lambda}$ from closely approximating the exact posterior (at low $\lambda$) to resembling the mean-field approximation (at high $\lambda$). The right panel shows the covariance between the two normal coordinates versus $\lambda$ on a log scale. Note that the $\Xi$-variational approximation to the covariance is very accurate up to a critical $\lambda$ ($\approx 10^{-1}$), after which it degrades rapidly to 00.
4.2 High Dimensional Linear Regression
$\Xi$-VI implies a statistical-computational tradeoff. To study this, we consider a a Bayesian linear model with Laplace prior,
$\bm{y}=\bm{X}\theta+\epsilon,\quad\epsilon\sim N(0,\sigma^{2}I_{n}),\quad\theta_{i}\sim\text{Laplace}(0,1).$ (4.3)
The Laplace prior has density $\pi(\theta_{i})=\frac{1}{2b}\exp\left(-\frac{|\theta_{i}|}{b}\right)$.
We simulate a dataset consisting of $n=100$ observations and $d=12$ features. The true regression coefficients is drawn randomly from a $12$-dimensional standard Gaussian distribution, and $\sigma^{2}=1$. Columns $(1,2,3,8,9)$ of $\bm{X}$ are generated from a standard Gaussian distribution. Then we set each of features $(4,5,6,11,12)$ equal to each of features $(1,2,3,8,9)$ plus a standard Gaussian noise. This setup aims to simulate realistic multicollinearity. Finally, we generate the response $\bm{y}$ using model (4.3). With this simulated data, we calculate an “exact” posterior with a long-run MCMC algorithm of 3,000 iterations. The MCMC draws produce an $\hat{R}$ of below $1.01$ across coefficients [27], which is well below the typical threshold of $1.1$ for satisfactory mixing.
Since coupling all $12$ coefficients is computationally expensive, we couple groups of coefficients in the EOT step. We adopt a naive grouping approach where features $(1,2,3)$, $(4,5,6)$, $(7,8,9)$, $(10,11,12)$ are grouped together. This effectively reduces the computational cost by reducing a twelve-dimensional coupling problem into a four-dimensional one. While it is beneficial to use informed grouping, any choice of grouping will improve the approximation accuracy of MFVI. For each dimension, we use $M=20$ support points to represent the marginal distributions, as was explained in the implementation Section 2.4.
With this simulated data, we use Algorithm 2 to compute the $\Xi$-VI approximation. In the first step, we use expectation propagation (EP) to compute the pseudomarginals. For the analysis, we chose 100 $\lambda$ values on a logarithmic scale from $10^{-4}$ to $10^{6}$, and represented the variational posterior for each $\lambda$ by $2,000$ sample points.
(a) Accuracy of $\Xi$-VI for Laplace linear regression. The accuracy is measured in $W_{2}$ across values of $\lambda$.
(b) Runtime of $\Xi$-VI for Laplace linear regression. Runtime is measured in the number of iterations to reduce the Sinkhorn error to $10^{-4}$, across values of $\lambda$.
Figure 2(a) shows the approximation errors of $\Xi$-VI as a function of $\lambda$, measured using the Wasserstein distance ($W_{2}$). These distances are computed between the posterior distributions sampled via MCMC and those obtained from $\Xi$-VI. The $\Xi$-VI approximation errors are benchmarked against the baseline errors of EP at $\lambda=\infty$, mean-field VI, and the theoretical lower bound at $\lambda=0$. A vertical line at $\lambda=D$, the number of features, marks an inflection point where the posterior variational approximation error transitions from rapidly converging to the EP error ($\lambda\leq D$) to relatively stable ($\lambda>D$).
Figure 2(b) shows the runtime of the approximate coordinate ascent algorithm for Laplace linear regression, measured in the number of iterations until convergence. The $\lambda$ values are shown on a logarithmic scale to highlight the performance over several orders of magnitude. The plot shows a sharp decline for $\lambda\leq D$ before it becomes stable at $\lambda>D$. The inflection in both the approximation error and runtime plots suggests that a regularization strength around $\lambda=D$ offers an optimal balance in the tradeoff between approximation accuracy and computational complexity.
4.3 Hierarchical Model
The Eight-school Model ([27], Section 5.5) is an important example of a hierarchical Bayesian model. Each of the eight schools provided separate estimates for the mean $y_{i}$ and standard deviation $\sigma_{i}$ of their respective treatment effects. Let $\theta_{j}$ be the treatment effect in school $j$. Given the absence of prior evidence favoring the effectiveness of any particular treatment, we treat the outcomes from each school as independent:
$y_{j}|\theta_{j}$ $\sim N(\theta_{j},\sigma_{j}^{2}),\quad\theta_{j}|\mu,\tau\sim\mathcal{N}(\mu,\tau^{2}),\quad 1\leq j\leq 8,$ (4.4)
$\mu$ $\sim N(0,5),\quad\tau\sim\text{halfCauchy}(0,5).$
where $\mu$ and $\tau$ are the global parameters common to all schools, $\theta_{j}$ is a local parameter specific to school $j$. The target of posterior inference are $\theta_{j}$’ s, $\mu$ and $\tau^{2}$.
Eq. 4.4 does not match the $\Xi$-VI formulation in Section 2.2, because the prior for $\theta_{j}$ are dependent on $\mu$ and $\tau$. Instead, we consider the reparameterization $z_{j}:=(\theta_{j}-\mu)/\tau$,
$y_{j}$ $\mid\mu,z_{j},\tau\sim N(\mu+\tau z_{j},\sigma_{j}^{2}),$ (4.5)
$z_{j}$ $\sim N(0,1),\quad\mu\sim N(0,5),\quad\tau\sim\text{halfCauchy}(0,5).$
This reparameterization transforms the joint prior of $z_{j}$’s, $\mu$, and $\tau$ into a product distribution. Despite the reparametrization, the mean-field posterior still greatly underestimates the correlations between parameters observed in the true posterior(See Figure 9 in Section 4.3).
$\Xi$-VI has the following formulation:
$\min_{\textrm{q}(z,\mu,\tau)}\mathbb{E}\left[\sum_{j=1}^{8}\frac{(y_{j}-\tau z_{j}-\mu)^{2}}{2\sigma_{j}^{2}}\right]+D_{\text{KL}}(\textrm{q}\parallel\pi)+\lambda\Xi(\textrm{q}).$ (4.6)
We use Algorithm 2 to solve Eq. 4.6. In the first step, we use Automatic Differentiation Variational Inference (ADVI) to compute a set of pseudomarginals. In the second step, we solve the following EOT problem:
$\min_{\textrm{q}(z,\mu,\tau)\in\mathcal{C}(\textrm{q}^{*}_{\infty})}\mathbb{E}_{\textrm{q}}\left[\sum_{j=1}^{8}\frac{\tau^{2}z_{j}^{2}+2(\mu-y_{j})\tau z_{j}}{2\sigma_{j}^{2}}\right]+(\lambda+1)D_{\text{KL}}(\textrm{q}\parallel\hat{m}_{\lambda}).$ (4.7)
Given that cost tensor can be decomposed into rank-three tensors, Eq. 4.7 is polynomial-time solvable by Proposition 1. In our implementation, we store the cost tensor as a set of third-order tensors, with values $(\tau^{2}z_{j}^{2}+2(\mu-y_{j})\tau z_{j})/(2\sigma_{j}^{2})$ for every $j$. This reduces the memory cost from $M^{10}$ to $8M^{3}$, where $M$ is the number of support points. We then apply the sum-product trick to implement the marginalization oracle in the Sinkhorn algorithm.
We recover the joint distribution for $(\theta_{1},\cdots,\theta_{8},\mu,\tau)$ by reparametrizing the optimal coupling in Eq. 4.7. To benchmark the performance, we compute the true posterior using MCMC draws with 4 chains for 1000 tune and 5000 draw iterations. Figure 3 shows the joint distribution of $\theta_{1},\theta_{7}$ under the true posterior, mean-field variational posterior and $\Xi$-variational posterior when $\lambda=0,1,10,1000$. We represent each of the approximate posteriors with $10,000$ sample points. The true posterior shows a strong positive correlation between $\theta_{1}$ and $\theta_{7}$, which is effectively captured by the $\Xi$-variational posterior at small $\lambda$. However, as $\lambda$ increases, the correlation decreases and the $\Xi$-VI results approach the MFVI where the correlation between $\theta_{1},\theta_{7}$ is underestimated.
Figure 3: Contour plots for the joint distribution of $\theta_{1}$ and $\theta_{7}$ across varying various variational approximations. The subplots compare the true posterior distribution with $\Xi$-variational posteriors for varying $\lambda$ values, and the MFVI approximation. A linear regression fitted slope of $\theta_{7}$ over $\theta_{1}$ is provided for each subplot. Each subplot includes a linear regression line showing the fitted slope of $\theta_{7}$ over $\theta_{1}$.
In inference that involves high-order posterior interactions, $\Xi$-VI outperforms not only MFVI but other VI methods. Figure 4 shows the posterior credible intervals for the maximum and minimum treatment effects across schools, comparing $\Xi$-VI with MFVI, normalizing flow variational inference (NFVI), Stein variational gradient descent (SVGD), and full-rank ADVI. $\Xi$-VI provides more accurate interval width and coverage accuracy for both maximum and minimum treatment effects when benchmarked against the true posterior. For the maximum treatment effect, while MFVI, NFVI, and full-rank ADVI produce excessively large credible intervals and SVGD produces excessively small intervals, $\Xi$-VI results closely approximate the true 95% posterior credible interval. For the minimum treatment effect, none of the VI methods precisely capture the true posterior interval. MFVI, NFVI, and full-rank ADVI produce smaller intervals with a noticeably downward-shifted center, SVGD offers a considerably undersized interval, whereas $\Xi$-VI generates reasonably-sized intervals with less downward shift compared to MFVI.
Figure 4: Comparison of the 95% posterior credible intervals for the maximum and minimum treatment effects across schools. The sequence from left to right includes the true posterior, MFVI, $\Xi$-VI with $\lambda=0,1,10,1000$, normalizing flow, full-rank ADVI, and Stein variational gradient descent.
Next, we consider the task of comparing treatment effects between schools. We compute the posterior credible intervals for the differences in treatment effects $\theta_{i}-\theta_{j}$ between schools $i$ and $j$. Table 1 (in Additional Analysis of the Eight School Example) shows credible intervals for $\theta_{i}-\theta_{j}$ across ten randomly chosen school pairs, calculated under the true posterior, MFVI, and $\Xi$-VI with $\lambda=0,1,10,1000$. The results show $\Xi$-VI, especially with lower $\lambda$ values, yields intervals that more accurately reflect those derived from the true posterior, while MFVI produces the most inaccurate intervals.
Figure 5 shows that $\Xi$-VI significantly reduces approximation errors in the standard deviation of the difference in treatment effect compared to MFVI. Also unsurprisingly, smaller values of $\lambda$ lead to a closer approximation of the true posterior standard deviation. This corrective effect of the $\Xi$-variational posterior is consistently observed across all $25$ randomly sampled parameter pairs. Remarkably, $\Xi$-VI provides more accurate estimates than all other VI methods, including normalizing flow and full-rank ADVI. Stein variational gradient descent (SVGD), however, provides results comparable to $\Xi$-VI with $\lambda=10$. It is also important to note that SVGD requires a longer runtime compared to our method. The comparison between $\Xi$-VI and other methods is presented Figure 10 in Additional Analysis of the Eight School Example.
Figure 5: Comparison of the approximation error of the standard deviation of the difference in treatment effect under MFVI and $\Xi$-VI with $\lambda=0,1,10,1000$.
Finally, we study the computation-statistical tradeoff of $\Xi$-VI. We evaluate our procedure on 100 $\lambda$ values logarithmically spaced from $10^{-3}$ to $10^{5}$. Figure 7 illustrates the approximation errors of the $\Xi$-variational posterior relative to the true posterior, measured using KL divergence and $W_{2}$ distance. These errors are benchmarked against those of MFVI at $\lambda=\infty$ and a theoretical lower bound at $\lambda=0$. Notably, a vertical line at $\lambda=D=10$ marks a critical transition: errors remain relatively stable for $\lambda<1$ and approach MFVI for $\lambda\geq 100$. Interestingly, this phase transition is not unique to the Wasserstein distance but is inherent to the model and the $\Xi$-VI method.
Figure 7 shows the runtime of the Algorithm 2 for Eight School model, measured in the number of iterations to reduce the Sinkhorn error (Algorithm 3) below $10^{-4}$. The regularization strength is shown on a logarithmic scale to highlight the performance over several orders of magnitude of $\lambda$. The plot shows a sharp decline right before and right after $\lambda=D$. The phase transition in both plots confirms that a regularization strength of $\lambda=D$ offers an optimal balance in the tradeoff between computational efficiency and approximation accuracy. Interestingly, a computational-statistical gap exists in this model: while $\lambda<1$ yields a closer approximation to the true posterior, optimal runtime is only achieved for $\lambda>10$.
Figure 6: $\Xi$-VI approximation accuracy for the Eight School model across varying $\lambda$, measured in KL divergence and Wasserstein-2 distance.
Figure 7: Runtime for the Eight School model as a function of varying $\lambda$, measured in the number of iterations to reduce the Sinkhorn error to $10^{-4}$.
Finally, we note that for the Eight-school model, the reparametrized MFVI is overdispersed. Generally, we recommend using overdispersed pseudomarginals in Algorithm 2. The advantage comes from an intuitive understanding of the one-step EOT correction: it seeks overlaps between the pseudomarginals and the true posterior to effectively capture the dependency information present in the true posterior. When the pseudomarginals are underdispersed, the one-step EOT correction still leads to underdispersed samples. With overdispersed pseudomarginals, the one-step EOT coupling compensates for the overdispersion by subsampling points from the marginals that reflect the dependency structure of the true posterior distributions, as seen in Figure 3.
5 Discussion
In this work, we introduced $\Xi$-VI, a new way of doing variational inference that extends MFVI through entropic regularization. We characterize the asymptotic normality of $\Xi$-variational posteriors in lower-dimensional scenarios and the tradeoff between computational complexity and statistical fidelity in higher-dimensional settings. Demonstrations on both simulated and empirical datasets underscore the superiority of our method over traditional MFVI. Importantly, our analysis reveals a deep connection between $\Xi$-VI and entropic optimal transport, allowing for the application of optimal transport techniques, such as the Sinkhorn algorithm, in the resolution of VI challenges, and vice versa.
A key question prompted by our work concerns the fundamental limits of high-dimensional Bayesian models. It is now well known that many high-dimensional problems exhibit a gap between what is statistically achievable (in a minimax sense) and what is achievable via a polynomial-time algorithm, such as sparse PCA [71] and denoising problems [18]. However, characterization of a statistical-computational gap remains unexplored in probabilistic machine learning. This is partly due to the gap between sampling methods such as MCMC and optimization methods such as VI. However, recent advancements are bridging the gap, with optimal transport playing a major role [40, 23, 19]. Our contribution aligns with this evolving line of work, offering fresh perspectives on the statistical-computational tradeoff inherent in VI and charting a course for further exploration.
Our theoretical results contributed novel insights to Bayesian statistics by identifying distinct asymptotic regimes corresponding to the true posterior and naive mean field. The transition between these regimes relates to classical phase-transition results on spin glass models [46]. While we derive these results in linear regression contexts, similar analyses could extend to models like the Ising model or the quadratic interaction model. Future work may explore relaxing assumptions of compact parameter space or log-concave posteriors and refining stability guarantees for VI algorithms under specific conditions.
The application of $\Xi$-VI in large-scale datasets presents computational challenges, yet distributed computing and stochastic optimization present can help circumvent these hurdles. But there is no readily available software for implementing multimarginal Sinkhorn algorithms at scale. Future efforts might focus on developing accessible, efficient software tools to facilitate the application of $\Xi$-VI across diverse models and data types.
While this paper implemented the examples using EP and BBVI for approximating posterior marginals, advanced mean-field methods such as the TAP approach may be preferable in certain contexts, such as spiked covariance models [26] and high-dimensional Bayesian linear regression [16]. Exploring $\Xi$-VI combined with the TAP method presents a promising avenue for future research, potentially providing a more accurate approximation to the true posterior.
Acknowledgements
We thank Sumit Mukherjee, Marcel Nutz, Yixin Wang, Eli Weinstein, Kaizheng Wang and Chenyang Zhong for their helpful discussion.
Appendix A Details of polynomial-time conditions for Algorithm 3
The two sufficient criteria for polynomial-time Sinkhorn algorithms are 1) the graphical model underlying the cost has bounded junction treewidth and 2) the cost tensor is low rank and sparse.
Definition 2 (Graphical Model Structure).
Let $S\subset 2^{[k]}$. The graphical model corresponding to $S$ is a graph $G_{S}=(V,E)$ with vertices $V=[k]$ and edges $E=\{(i,j):i,j\in S,\text{ for some }S\in S\}$.
Definition 3 (Graphical Model).
Let $S\subset 2^{[k]}$. A probability distribution $P$ over $\{X_{i}\}_{i\in[k]}$ is a graphical model with structure $S$ if there exist functions $\{\psi_{S}\}_{S\in S}$ and a normalizing constant $Z$ such that
$P(\{x_{i}\}_{i\in[k]})=\frac{1}{Z}\prod_{S\in S}\psi_{S}(\{x_{i}\}_{i\in S}).$
The junction treewidth of a graphical model captures the storage complexity and the computational complexity for performing basic inference tasks such as computing the mode, log-partition function, and marginals of a joint distribution (See Section 2, [69]).
Definition 4 (Junction Tree, Treewidth).
A junction tree $T=(V_{T},E_{T},\{B_{u}\}_{u\in V_{T}})$ for a graph $G=(V,E)$ consists of a tree $(V_{T},E_{T})$ and a set of bags $\{B_{u}\subseteq V\}_{u\in V_{T}}$ satisfying:
  • For each variable $i\in V$, the set of nodes $U_{i}=\{u\in V_{T}:i\in B_{u}\}$ induces a subtree of $T$.
  • For each edge $e\in E$, there is some bag $B_{u}$ containing both endpoints of $e$.
The width of the junction tree is $\max_{u\in V_{T}}|B_{u}|-1$. The treewidth of a graph is the width of its minimum-width junction tree.
Next, we review the concept of tensor rank, which is a natural analog of matrix rank.
Definition 5 (Tensor Rank).
A rank-$r$ factorization of a tensor $\bm{R}\in\mathbb{R}^{n\times\cdots\times n}$ is a collection of $r\times k$ vectors $\{u_{i,l}\}_{i\in[k],l\in[r]}\subset\mathbb{R}^{n}$ satisfying
$\bm{R}=\sum_{l=1}^{r}\bigotimes_{i=1}^{k}u_{i,l}.$
The rank of a tensor is the minimal $r$ for which there exists a rank-$r$ factorization.
We now consider these definitions in the context of optimal transport.
Definition 6 (Graphical Structure for EOT).
A cost tensor $C\in(R^{n})^{\otimes k}$ has graphical structure with treewidth $\omega$ if there is a graphical model structure $S\subset 2^{[k]}$ and functions $\{f_{S}\}_{S\in S}$ such that
$C_{\mathbf{j}}=\sum_{S\in S}f_{S}(\{j_{i}\}_{i\in S}),\quad\forall\mathbf{j}=(j_{1},\ldots,j_{k})\in[n]^{k},$
and the graph $G_{S}$ has treewidth $\omega$.
The functions $\{f_{S}\}_{S\in S}$ can be arbitrary so long as the corresponding graphical model structure has treewidth at most $\omega$.
Definition 7 (Low-Rank Plus Sparse Structure for EOT).
A cost tensor $\mathbf{C}\in\mathbb{R}^{n\times\cdots\times n}$ has a low-rank plus sparse structure of rank $r$ and sparsity $s$ if it decomposes as
$\bm{C}=\bm{R}+\bm{S},$
where $\bm{R}$ is a rank-$r$ tensor and $\bm{S}$ is an $s$-sparse tensor.
In implementation, we could represent $\bm{S}$ through its $s$ non-zero entries, and $\bm{R}$ through a rank-$r$ factorization. Under the bounded cost assumption, the entries of both $\bm{R}$ and $\bm{S}$ are of size $O(C_{\text{max}})$. This rules out the case of having extremely large entries for $\bm{R}$ and $\bm{S}$, one positive and one negative, which cancel each other to yield a small entry of $\mathbf{C}=\bm{R}+\bm{S}$.
Appendix B Algorithms for computing posterior marginals
Many algorithms exist for computing the pseudomarginals in Algorithm 2. Here we present the expectation propagation (EP) algorithm and the black-box variational inference (BBVI) algorithm. EP [45] is a generalized loopy belief propagation algorithm that targets the reverse VI by minimizing $\min_{m\in\mathcal{M}(\Theta)}D_{\text{KL}}(\textrm{q}^{*}_{0}\parallel m)$ locally. BBVI is a MFVI algorithm that minimizes $D_{\text{KL}}(\cdot\parallel\textrm{q}^{*}_{0})$ over a parametrized family of product measures by computing black-box gradients of the log-joint.
In both algorithms, the marginals $m(\cdot|\eta)$ are parametrized using a finite- dimensional parameter $\eta$ to make the computation tractable. The parametric class $m(\cdot|\eta)$ is typically chosen to match the support of the exact posterior. For example, we approximate distributions with continuous unbounded support as Gaussian distributions where $\eta$ includes the mean and diagonal covariance matrix.
Expectation Propagation (EP).
Expectation Propagation approximates the posterior marginals by iteratively matching the moments of factors of the approximating distribution to the exact posterior. It assumes that the posterior decomposes as a product of factors, $\textrm{q}^{*}_{0}\propto\prod_{k=0}^{n}f^{k}(\theta)$. One possible choice is $f^{0}(\theta)=\pi(\theta)$ and $f^{k}(\theta)=\ell(x_{k};x_{1:k-1},\theta)$.
We posit a mean-field variational family with the same factor structure.
$m(\theta)\propto\prod_{k=0}^{n}m^{k}(\theta).$
EP updates the individual components $m^{k}=(m^{k}_{1},\ldots,m^{k}_{D})$ via iterative moment-matching operations, which minimizes the KL divergence from each factor of the posterior to its variational approximation.
In each iteration, for $k=0,\ldots,n$, we update the global approximation $m(\theta)$ by incorporating the exact posterior factor $f_{k}(\theta)$, which leads to to the construction of the cavity distribution $m^{-k}(\theta)$ and the tilted distribution $m^{\backslash k}(\theta)$,
$m^{-k}(\theta)$ $=\prod_{j\neq k}m^{j}(\theta),$
$m^{\backslash k}(\theta)$ $\propto f_{k}(\theta)m^{-k}(\theta).$
Computing $m^{-k}$ involves computing a normalizing constant. The next step optimizes $m_{\text{new}}^{k}(\theta)$ to best approximate $m^{\backslash k}(\theta)$ within the mean-field variational family, according to the following optimization problem:
$m_{\text{new}}^{k}(\theta)=\mathop{\mathrm{argmin}}_{m\in\mathcal{M}(\Theta)}D_{\text{KL}}(m^{\backslash k}(\theta)\parallel m(\theta)).$ (B.1)
We can solve this problem efficiently via moment matching. For instance, if $m(\theta\mid\eta)$ is parametrized as a multivariate Gaussian, then Eq. B.1 is equivalent to performing the Laplace approximation on $m$ and matching its mean and the diagonal covariance matrix with the approximate distribution.
We update the local factor $m^{k}(\theta)$ that to reflect the new information, according to:
$m^{k}(\theta)=\frac{m_{\text{new}}^{k}(\theta)}{m^{-k}(\theta)}.$ (B.2)
After updating the local factors $m^{0},\ldots,m^{n}$, we update the global distribution $m$ independently across each dimension:
$m_{i}(\theta_{i})=\frac{1}{Z_{i}}\prod_{k=0}^{n}m^{k}_{i}(\theta_{i}),$ (B.3)
[…]
The iteration repeats until convergence and returns a distribution that closely approximates the exact posterior in terms of its moments.
For Gaussian approximations, moment matching sets the mean and covariance of the tilted distribution $m^{\backslash k}(\theta)$ to the updated factor $m_{\text{new}}^{k}(\theta)$. For exponential family distributions, moment matching matches the sufficient statistics of the tilted distribution to those of the updated local factor. The updated local factor $m_{\text{new}}^{k}(\theta)$ is then the distribution in the approximating family that has these matched sufficient statistics.
Let $m(\cdot\mid\eta)$ be the parametrized mean-field family. The algorithmic implementation of EP is summarized in Algorithm 4.
Input: Set of true posterior factors $\{f^{k}(\theta)\}_{k=0}^{n}$, initial parameter $\eta^{k}_{0}$ for $k=0,\ldots,n$.
Output: Approximate posterior distribution $m(\theta)$.
Initialize: Set initial parameters $m^{k}(\theta)=m(\theta\mid\eta=\eta_{0}^{k})$ for $k=0,\ldots,n$, and define the global approximation $m(\theta)\propto\prod_{k=0}^{n}m^{k}(\theta)$, set iteration $t=0$.
while convergence not met do
       for $k=0$ to $n$ do
             Compute the cavity distribution $m^{-k}(\theta)=\prod_{j\neq k}m^{j}(\theta\mid\eta=)$.
             Compute the tilted distribution $m^{\backslash k}(\theta)\propto f_{k}(\theta)\cdot m^{-k}(\theta)$.
             Update $\eta^{k}_{\text{new}}=\mathop{\mathrm{argmin}}_{\eta}D_{\text{KL}}(m^{\backslash k}(\theta)\parallel m(\theta\mid\eta))$ via moment matching.
             Update $m^{k}(\theta)=\frac{m(\theta\mid\eta=\eta^{k}_{\text{new}})}{m^{-k}(\theta)}$.
       end for
      Update $m(\theta)\propto\prod_{k=0}^{n}m^{k}(\theta)$ with normalization.
       Increment $t=t+1$.
end while
Algorithm 4 Expectation Propagation
We check the convergence of Algorithm 4 by examining the change in the parameters of the approximating factors. The algorithm terminates when the change is small.
The EP algorithm is well suited when the exact posterior decomposes into factors that allow for direct moment calculation. The models with conjugate exponential family distributions fall into this category. For non-conjugate models, EP can adapt through numerical integration or Monte Carlo methods to approximate the required moment matching. Though EP typically approximates the posterior marginals more accurately than MFVI, EP converges more slowly and overestimate the posterior variance, because the EP approximation relies on a convex outer bound on the set of mean parameters as opposed to the mean-field approximation which constructs a nonconvex inner bound on the same set [69].
Black Box Variational Inference.
MFVI solves the following variational problem:
$\tilde{m}=\mathop{\mathrm{argmin}}_{m\in\mathcal{M}(\Theta)}D_{\text{KL}}(m\parallel\textrm{q}^{*}_{0}).$ (B.4)
Let the approximate posterior $m(\theta\mid\eta)$ be parametrized by $\eta$, called the mean-field parameter. MFVI can be cast as an optimization problem for finite-dimensional parameters $\eta$, where objective is the ELBO,
$\mbox{ELBO}(\eta):=\mathbb{E}_{m(.\mid\eta)}\left[\ell(\bm{x};\theta)+\log\pi(\theta)-\log m(\theta\mid\eta)\right].$
While other techniques are available, we use a gradient-based optimization technique called Black Box Variational Inference (BBVI) to optimize the ELBO.
The procedure requires only the evaluation of the joint log-likelihood of the model [57], making it useful for a wide range of possible models.
In BBVI, we maximize the ELBO as a function of $\eta$ using stochastic gradient descent (SGD). The gradient of ELBO is computed as an expectation with respect to the approximate posterior. We calculate a noisy but unbiased estimate of the gradient using draws from the approximate posterior, which is then used to update the mean-field parameters via a gradient-descent step.
Mathematically, we can write the gradient of the ELBO as follows:
$\nabla_{\eta}\text{ELBO}=\mathbb{E}_{m(.\mid\eta)}\left[\nabla_{\eta}m(\theta\mid\eta)\left(\ell(\bm{x};\theta)+\log\pi(\theta)-\log m(\theta\mid\eta)\right)\right].$ (B.5)
We use Monte Carlo approximations to compute a noisy, unbiased estimate of the gradients:
$\nabla_{\eta}\text{ELBO}\approx\frac{1}{S}\sum_{s=1}^{S}\nabla_{\eta}m(\theta[s]\mid\eta)\left(\ell(\mathbf{x};\theta[s])+\log\pi(\theta[s])-\log m(\theta[s]\mid\eta)\right),$ (B.6)
where $\theta[s]\sim m(\theta\mid\eta)$ are drawn independently.
Algorithm 5 outlines the Black-Box Variational Inference (BBVI) procedure for computing the mean-field variational posterior. This algorithm is based on the Monte-Carlo approximation provided by Eq. B.6.
Input: Log-likelihood $\ell(\bm{x};\theta)$, prior $\pi$, initial parameter $\eta_{0}$.
Initialize: $\eta=\eta_{0},t=1$;
while change of $\eta$ is greater than 0.01 do
       for $s\in[S]$ do
             Draw $\theta[s]\stackrel{{\scriptstyle iid}}{{\sim}}m(.\mid\eta)$;
       end for
      Compute $\rho_{t}=t^{th}$ value of a Robbins Monro sequence;
       Compute $\eta=\eta+\rho_{t}\frac{1}{S}\sum_{s=1}^{S}\nabla_{\eta}m(\theta[s]|\eta)(\tilde{\ell}(\bm{x};\theta[s])+\log\pi(\theta[s])-\log m(\theta[s]|\eta))$;
       Increment $t=t+1$;
end while
Output: $m(\theta\mid\eta)$.
Algorithm 5 Black Box Variational Inference
We choose the learning rate $\rho_{t}$ to follow the Robbins-Monro conditions
$\sum_{t=1}^{\infty}\rho_{t}=\infty,\qquad\sum_{t=1}^{\infty}\rho_{t}^{2}<\infty,$
This condition guarantees convergence of the stochastic optimization to a local maximizer of the ELBO.
Based on the Bernstein-von Mises theorem and Proposition 2, it is natural to set $\tilde{m}_{\lambda}$ as a convex combination of the Expectation Propagation (EP) and mean-field approximations. Though not implemented in our examples, this approach could potentially better approximate the marginals of $\textrm{q}^{*}_{\lambda}$. For a mean-field normal approximation $\tilde{m}_{\text{MF}}:=N(\mu_{\text{MF}},\Sigma_{\text{MF}})$ and an EP approximation $\tilde{m}_{\text{EP}}:=N(\mu_{\text{EP}},\Sigma_{\text{EP}})$, we consider:
$\tilde{m}_{\lambda}=N(\mu_{\text{MF}},\Sigma_{\lambda}),\quad\Sigma_{\lambda}=\left(\frac{1}{\lambda+1}\Sigma_{\text{EP}}^{-1}+\frac{\lambda}{\lambda+1}\Sigma_{\text{MF}}^{-1}\right)^{-1},$ (B.7)
where $\Sigma_{\lambda}$ is a diagonal matrix. This formulation suggests $\Sigma_{\lambda}$ as a weighted inverse combination of $\Sigma_{\text{EP}}$ and $\Sigma_{\text{MF}}$, adjusting the covariance structure based on the regularization parameter $\lambda$. When EP accurately approximates the marginals of $\textrm{q}^{*}_{0}$ and MFVI produces $\textrm{q}^{*}_{\infty}$, this approach could leverage the strengths of both approaches to accurately approximate the marginals of $\textrm{q}^{*}_{\lambda}$.
Appendix C Support Results
Lemma 1 (Gibbs variational principle).
For probability measures $\mu$ on $\Theta$,
$\log\mathbb{E}_{\mu}[\exp(f(\theta))]=\sup_{\nu\in\mathbb{P}(\Theta)}\left\{\mathbb{E}_{\nu}[f(\theta)]-D_{\text{KL}}(\nu\parallel\mu)\right\}.$ (C.1)
Lemma 2.
Let $\Sigma$ be a symmetric, positive definite matrix of size $D$ with eigenvalues $\lambda_{1},\cdots,\lambda_{n}$ and eigendecomosition $\Sigma=Q\Lambda Q^{-1}$ where $\Lambda$ is a diagonal matrix such that $\Lambda_{jj}=\lambda_{j}$ for all $1\leq j\leq n$. Also, let $\sigma_{j}^{2}=\Sigma_{jj}$ for all $j$. Then,
$\begin{pmatrix}\sigma_{1}^{2}\\ \sigma_{2}^{2}\\ \cdots\\ \sigma_{D}^{2}\end{pmatrix}=(Q\circ Q)\begin{pmatrix}\lambda_{1}\\ \lambda_{2}\\ \cdots\\ \lambda_{D}\end{pmatrix},$
where $\circ$ denotes the Hadamard product, i.e. $(A\circ B)_{ij}=[A]_{ij}[B]_{ij}$.
Proof.
Let $\lambda$ be the vector of $\lambda_{i}$’s. Let’s show $\sigma_{i}^{2}=(Q\circ Q\lambda)_{i}$.
$(Q\circ Q\lambda)_{i}$ $=\sum_{k=1}^{D}(Q\circ Q)_{ik}\lambda_{k}=\sum_{k=1}^{D}Q_{ik}Q_{ik}\lambda_{k}=\sum_{k=1}^{D}Q_{ik}\lambda_{k}Q_{ik}=\sum_{k=1}^{D}Q_{ik}\lambda_{k}Q^{T}_{ki}$
$=\sum_{k=1}^{D}Q_{ik}\Lambda_{kk}Q^{T}_{ki}=\sum_{j=1}^{D}\sum_{k=1}^{D}Q_{ik}\Lambda_{jk}Q^{T}_{ji}=[Q\Lambda Q^{T}]_{ii}=\Sigma_{ii}=\sigma_{i}^{2}.$
Lemma 3.
Let $A$ be a symmetric, positive definite matrix. For all $j\in[D]$, it holds that $(A^{-1})_{jj}\geq\frac{1}{A_{jj}}$.
Proof.
Given that $A$ is symmetric and positive definite, there exists another symmetric positive definite matrix $B$ such that $B^{2}=A$. We note that $A_{jj}=e_{j}^{T}Ae_{j}=e_{j}^{T}B^{T}Be_{j}=\|Be_{j}\|_{2}^{2}$ and similarly, $(A^{-1})_{jj}=\|B^{-1}e_{j}\|_{2}^{2}$.
By the Cauchy-Schwarz inequality, we have
$\langle Be_{j},B^{-1}e_{j}\rangle^{2}\leq\|Be_{j}\|_{2}^{2}\|B^{-1}e_{j}\|_{2}^{2}=A_{jj}(A^{-1})_{jj}.$
However,
$\langle Be_{j},B^{-1}e_{j}\rangle=e_{j}^{T}(B^{-1})^{T}Be_{j}=e_{j}^{T}B^{-1}Be_{j}=e_{j}^{T}e_{j}=1.$
Therefore, we have $A_{jj}(A^{-1})_{jj}\geq 1$, which simplifies to $(A^{-1})_{jj}\geq\frac{1}{A_{jj}}$. This completes the proof. ∎
Lemma 4.
For $\textrm{q}\in\mathbb{P}(\Theta)$, the following variational characterization of its expressivity holds:
$\Xi(\textrm{q})=\sup_{f}\mathbb{E}_{\textrm{q}}[f]-\log\mathbb{E}_{\textrm{q}_{i}}\mathbb{E}_{\textrm{q}_{-i}}[\exp(f(\theta_{i},\theta_{-i}))\mid\theta_{i}].$
Proof.
Apply Donsker-Vardhan lemma. We obtain
$\Xi(\textrm{q})=\sup_{f}\mathbb{E}_{\textrm{q}}[f]-\log\mathbb{E}_{\textrm{q}_{i}}\mathbb{E}_{\textrm{q}_{-i}}[\exp(f(\theta_{i},\theta_{-i}))].$
Apply Donsker-Vardhan lemma again to $D_{\text{KL}}(\textrm{q}(\theta_{i},\theta_{-i})$
Theorem 8 (Theorem 5.11, [63]).
In the space $\mathbb{P}_{p}(\mathbb{R}^{d})$, we have $W_{p}(\mu_{n},\mu)\to 0$ if and only if $\mu_{n}\to\mu$ weakly and
[…]
where $p>0$ is a given exponent.
Lemma 5.
Let $\textrm{q}_{n}$ be a sequence of measures in $\mathbb{P}_{2}(\Theta)$. If $W_{2}(\textrm{q}_{n},\textrm{q})\to 0$ for some $\textrm{q}\in\mathbb{P}_{2}(\Theta)$, then
$\liminf_{n\to\infty}\Xi(\textrm{q}_{n})\geq\Xi(\textrm{q}).$
Let $\textrm{q}_{0}$ be another measure in $\mathbb{P}_{2}(\Theta)$. We have
$\liminf_{n\to\infty}D_{\text{KL}}(\textrm{q}_{n}\parallel\textrm{q}_{0})\geq D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{0}).$
Proof.
The second property follows from the fact that functional $D_{\text{KL}}(\cdot\parallel\textrm{q}^{*}_{0})$ is continuous in the Wasserstein metric (Proposition 7.1, [63]). For any $_{n}\overset{W_{2}}{\to}\textrm{q}$ q, Theorem 8 implies that $\textrm{q}_{n}$ weakly converge to $\textrm{q}_{0}$. The convergence $W_{2}(\textrm{q}_{n},\textrm{q}_{0})\to 0$ implies the convergence $W_{2}(\textrm{q}_{n,i},\textrm{q}_{0,i})\to 0$ for each $i\in[D]$. Since $D_{\text{KL}}$ is lower semicontinuous in both arguments (Theorem 4.8, [55]), we get
$\liminf_{n\to\infty}\Xi(\textrm{q}_{n})=\liminf_{n\to\infty}D_{\text{KL}}\left(\textrm{q}_{n}\parallel\prod_{i=1}^{D}\textrm{q}_{n,i}\right)\geq D_{\text{KL}}\left(\textrm{q}_{0}\parallel\prod_{i=1}^{D}\textrm{q}_{0,i}\right)=\Xi(\textrm{q}),$ (C.2)
where $D$ is fixed with respect to $n$. ∎
In the optimal transport theory, the concept of a shadow is used to establish quantitative results for transport between distributions with different marginals. A shadow, denoted by $\textrm{q}^{s}$, is constructed by gluing optimal transports such that it ”mimics” the optimal coupling q, given another set of marginals.
Definition 8 (Shadow).
Let $p\in[1,\infty]$ and $m,\tilde{m}$ be product measures within $\mathbb{P}_{p}(\Theta)$. Assume $\kappa_{i}\in\mathcal{C}(m_{i},\tilde{m}_{i})$ is a coupling that achieves $W_{p}(m_{i},\tilde{m}_{i})$ and let $\kappa_{i}=m_{i}\otimes K_{i}$ represent a disintegration. For a given $\textrm{q}\in\mathcal{C}(m)$, its shadow $\textrm{q}^{s}\in\mathcal{C}(\tilde{m})$ is defined as the second marginal of $\textrm{q}\otimes K\in\mathbb{P}(\Theta\times\Theta)$, where the kernel $K:\Theta\rightarrow\mathbb{P}(\Theta)$ is constructed as a direct sum $K(x)=K_{1}(x_{1})\otimes\ldots\otimes K_{D}(x_{D})$.
Shadow is defined through kernels that map the original spaces to probability measures. The uniqueness of these optimal kernels is not guaranteed, and any suitable kernel that satisfies the optimality conditions is a valid shadow. In this context, we exclusively focus on shadow couplings induced by the $W_{2}$ distance.
Given a coupling $\textrm{q}\in\mathcal{C}(m)$, its shadow $\textrm{q}^{s}$ satisfied the following properties.
Lemma 6 (Lemma 3.2, [24]).
For product distributions $m,\tilde{m}\in\mathbb{P}_{2}(\Theta)$ and coupling $\textrm{q}\in\mathcal{C}(m)$, its shadow $\textrm{q}^{s}\in\mathcal{C}(\tilde{m})$ satisfies
$W_{2}(\textrm{q},\textrm{q}^{s})=W_{2}(m,\tilde{m}),\quad D_{f}(\textrm{q}^{s}\parallel\tilde{m})\leq D_{f}(\textrm{q}\parallel m),$
where $D_{f}(\cdot)$ is any $f$-divergence.
Theorem 9 (Theorem 12, [41]).
Let $\{\phi^{t}\}_{t\geq 0}$ be the iterates generated by Algorithm 3. The number of iterations $t$ required to reach the stopping criterion $E\leq\epsilon^{\prime}$ is upper bounded by:
$t\leq 2+\frac{2D^{2}\left[\|\bm{C}\|_{\infty}/(\lambda+1)-\log\left(\min_{1\leq i\leq M,1\leq j\leq D}m_{ij}\right)\right]}{\epsilon^{\prime}}.$
Proofs of Section 2
Proof of Eq. 2.6.
Let $m$ be given. Then to optimize $\textrm{q}^{*}$, we have
$\textrm{q}^{*}(\theta)$ $=\mathop{\mathrm{argmin}}\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}[-\ell(\bm{x};\theta)]+\lambda D_{\text{KL}}(\textrm{q}\parallel m)+D_{\text{KL}}(\textrm{q}\parallel\pi)$
$=\mathop{\mathrm{argmin}}\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}[-\ell(\bm{x};\theta)]+\lambda D_{\text{KL}}(\textrm{q}\parallel m)+D_{\text{KL}}(\textrm{q}\parallel m)+D_{\text{KL}}(m\parallel\pi)$
$=\mathop{\mathrm{argmin}}\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}[-\ell(\bm{x};\theta)]+(\lambda+1)D_{\text{KL}}(\textrm{q}\parallel m).$
The first line uses the fact that $m$ is a product distribution. The third line drops $D_{\text{KL}}(m\parallel\pi)$ as it does not depend on q. ∎
Proof of Theorem 1.
Define an auxiliary probability distribution $\textrm{q}_{\text{aux}}\in\mathbb{P}(\Theta)$.
$\textrm{q}_{\text{aux}}=\mathcal{Z}_{n}(\lambda)^{-1}\exp\left(\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)m(\theta).$ (C.3)
Since $\sup_{\theta\in\Theta}\bm{l}\left(\bm{x};\theta\right)<\infty$, the normalizing constant $\mathcal{Z}_{n}(\lambda)$ must be finite, thus $\textrm{q}_{\text{aux}}$ is well-defined and absolutely continuous with respect to $m$.
Minimizing the objective function in Eq. 2.6 is equivalent to minimizing the KL loss to $\textrm{q}_{\text{aux}}$.
$\min\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}\left[-\ell(\bm{x};\theta)\right]+(\lambda+1)D_{\text{KL}}(\textrm{q}\parallel m)$
$=\min\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}\left[-\frac{1}{\lambda+1}\ell(\bm{x};\theta)\right]+D_{\text{KL}}(\textrm{q}\parallel m)$
$=\min\limits_{\textrm{q}\in\mathcal{C}(m)}\int_{\Theta}\textrm{q}(\theta)\log\frac{\textrm{q}(\theta)}{\exp\left(\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)m(\theta)}d\theta$
$=\min\limits_{\textrm{q}\in\mathcal{C}(m)}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{\text{aux}})-\log\mathbb{E}_{m}\exp\left(\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)$
[…]
Since $\mathcal{C}(m)$ is displacement convex and the KL loss is displacement convex [67], the solution is unique.
Conversely, let $\textrm{q}^{*}$ be the solution to the EOT primal problem. Since the problem is convex, we can write the minimizer $\textrm{q}^{*}$ using Langragian formulation.
$\textrm{q}^{*}=\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathbb{P}(\Theta)}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{\text{aux}})+\sum_{i=1}^{D}[\mathbb{E}_{m_{i}}\phi_{i}-\mathbb{E}_{\textrm{q}_{i}}\phi_{i}],$
where $\phi_{i}\in L^{\infty}(m_{i})$ are the dual variables. This is equivalent to
$\textrm{q}^{*}=\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathbb{P}(\Theta)}D_{\text{KL}}(\textrm{q}\parallel C^{-1}e^{\sum_{i=1}^{D}\phi_{i}}\textrm{q}_{\text{aux}})+\sum_{i=1}^{D}\mathbb{E}_{m_{i}}\phi_{i},$ (C.5)
[…]
Since $\phi_{i}$ are uniformly bounded, $C$ is positive and finite. We can absorb $C$ into the potentials $\phi_{1},\cdots,\phi_{D}$ to obtain a desired form:
$\textrm{q}^{*}=\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)m(\theta).$
”only if” direction: Assume that the optimal coupling $\textrm{q}^{*}$ is given by
$\textrm{q}^{*}=\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)m(\theta).$
where $\phi=(\phi_{1},\cdots,\phi_{D})\in\prod_{i=1}^{D}L^{1}(m_{i})$ are some potential functions.
Plugging the solution in the EOT primal problem, for each $i$ and $[m_{i}]$-a.s.$\theta_{i}$, the potentials satisfy a set of fixed point equations called the Schrödinger system:
[…]
The Schrödinger system (Eq. C.6) satisfies the Euler-Lagrange optimality condition for the primal EOT problem [15]. Precisely, the EOT potentials solve
$\max_{\phi\in\prod_{i=1}^{D}L^{1}(m_{i})}\sum_{i=1}^{D}\mathbb{E}_{m_{i}}\phi_{i}-\mathbb{E}_{m}\left[\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)\right],$
which is the dual problem to the multimarginal EOT problem (Eq. 2.6). Since the EOT problem is convex [50], the primal-dual gap closes, which means the probability measure q defined under $\phi$ solves Eq. 2.6.
To see that $\phi_{i}\in L^{\infty}(\Theta_{i})$ for $i\in[D]$. Assume that $\mathbb{E}_{m_{i}}[\phi_{i}]\geq 0$, which is possible under the Euler-Langrange condition:
$\sum_{i=1}^{D}\mathbb{E}_{m_{i}}\phi_{i}=\min\limits_{\textrm{q}\in\mathcal{C}(m)}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{\text{aux}})\geq 0\$
By Eq. C.6, we apply Jensen’s inequality to obtain that
[…]
$\leq-\mathbb{E}_{m_{-i}}\left[\sum_{j\neq i}\phi_{j}(\theta_{j})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right]\leq-\frac{1}{\lambda+1}\mathbb{E}_{m_{-i}}\left[\ell\left(\bm{x};\theta\right)\right],$
thus $\sup_{\theta_{i}\in\Theta_{i}}\phi_{i}(\theta_{i})\leq-\sup_{\theta\in\Theta}|\ell\left(\bm{x};\theta\right)|/(\lambda+1)$ for all $i\in[D]$.
For the other direction, since $\sup_{\theta\in\Theta}\bm{l}\left(\bm{x};\theta\right)<\infty$, we have
[…]
Since the right-hand side of the inequality does not depend on $\theta_{i}$, $\inf_{\theta_{i}\in\Theta_{i}}\phi_{i}(\theta_{i})>-\infty$ as long as $\sum_{j\neq i}\phi_{j}(\theta_{j})<\infty$ holds $[m_{-i}]$-almost surely. Since $\sup_{\theta_{i}\in\Theta_{i}}\phi_{i}(\theta_{i})\leq-\sup_{\theta\in\Theta}\ell\left(\bm{x};\theta\right)/(\lambda+1)$, we have that $\inf_{\theta_{i}\in\Theta_{i}}\phi_{i}(\theta_{i})>-\infty$ for all $i\in[D]$. ∎
Proof of Eq. 2.13.
We make the following derivation,
$\min\limits_{m\in\mathcal{M}(\Theta)}\min\limits_{\textrm{q}\in\mathcal{C}(m)}\mathbb{E}_{\textrm{q}}[-\ell(\bm{x};\theta)]+(\lambda+1)D_{\text{KL}}(\textrm{q}\parallel m)+D_{\text{KL}}(m\parallel\pi)$
$=\min\limits_{m\in\mathcal{M}(\Theta)}(\lambda+1)\int_{\Theta}\textrm{q}^{*}(\theta)\log\frac{\textrm{q}^{*}(\theta)}{\exp(\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right))m(\theta)}d\theta+D_{\text{KL}}(m\parallel\pi)$
$=\min\limits_{m\in\mathcal{M}(\Theta)}(\lambda+1)\int_{\Theta}\textrm{q}^{*}(\theta)\log\frac{\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)}{\exp(\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right))}d\theta+D_{\text{KL}}(m\parallel\pi)$
$=\min\limits_{m\in\mathcal{M}(\Theta)}\mathbb{E}_{m}\left[\underbrace{(\lambda+1)\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})\right)\exp\left(\sum_{i=1}^{D}\phi_{i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x};\theta\right)\right)}_{\text{surrogate loss}}\right]+D_{\text{KL}}(m\parallel\pi).$
Proofs of Section 2.4
Proof of Proposition 1.
By Theorem 14 of [41], Algorithm 3 reaches the stopping criterion $\bm{E}\leq\epsilon$ in $t$ iterations, where $t$ satisfies
$t\leq 2+\frac{2D^{2}\|\bm{C}\|_{\infty}-\log(\max_{ij}\bm{M}_{ij})}{\epsilon(\lambda+1)}.$
This implies that
$t\asymp\text{poly}(D,\bm{C}_{\max}/\epsilon,\frac{1}{\lambda+1}).$ (C.7)
Algorithm 3 calls the following oracle $D$ times:
$\text{Compute}\quad r_{i}(\bm{Q})=\frac{\sum_{1\leq k_{j}\leq M,\forall j\neq i}\exp\left[\sum_{j=1}^{D}\bm{F}_{jk_{j}}-\frac{1}{\lambda+1}\bm{C}_{k_{1}\cdots k_{D}}\right]\bm{M}_{k_{1}\cdots k_{D}}}{\sum_{1\leq k_{j}\leq M,\forall j\in[D]}\exp\left(\sum_{i=1}^{D}\bm{F}_{ik_{i}}-\frac{1}{\lambda+1}\bm{C}_{k_{1}\cdots k_{D}}\right)\bm{M}_{k_{1}\cdots k_{D}}}.$ (C.8)
The other steps are computed in linear time.
By Theorem 5.5 and Theorem 7.4 of [5], the oracle can be computed in $\text{poly}(M,D)$ iterations. Repeating the oracle complexity $Dt$ times, by Eq. C.7, the algorithm terminates in $\text{poly}(M,D,\bm{C}_{\max}/\epsilon,\frac{1}{\lambda+1})$ time. ∎
Proof of Corollary 1.
We consider Algorithm 1 from [25]. The algorithm implements the marginalization in Algorithm 3 using the sum-product method. Consider a graph $G=([D],E,K)$, where $[D],E,K$ represent the set of nodes, edges, and maximal cliques. If the log-likelihood $\ell(\bm{x}^{(n)};\theta)$ factorizes according to $G$, by the Hammersley-Clifford theorem, we get
$\ell\left(\bm{x};\theta\right)=\sum_{\alpha\in K}\ell_{\alpha}\left(\theta_{\alpha}\right),$
with $\ell_{\alpha}$ is defined over $\prod_{j\in\alpha}\Theta_{j}$.
Define $\bm{C}_{k_{\alpha}}$ as the tensor of $\ell_{\alpha}(\theta_{\alpha})$ values at support points $(\theta_{i}^{(s)},i\in\alpha)_{s\in[M]}$. The cost tensor decomposes as follows:
$\bm{C}_{k_{1},\cdots,k_{D}}=\sum_{\alpha\in K}\bm{C}_{k_{\alpha}}.$
Let $t$ be the iteration count for Algorithm 1 of [25] to terminate with criterion $\epsilon$. By Theorem 1 of [25], we get
$\mathbb{E}[t]=O(D^{2}\max_{\alpha\in K}\|\bm{C}_{k_{\alpha}}\|_{\infty}(\lambda+1)^{-1}\epsilon^{-1}).$
With ${\mathcal{T}}$ as the minimal junction tree for $G$, marginalizing over each factor in ${\mathcal{T}}$ takes $O(M^{\omega(G)})$ iterations, and message passing takes $O(d({\mathcal{T}})M^{\omega(G)})$ iterations, where $d({\mathcal{T}})$ is the average leaf distance in ${\mathcal{T}}$.
Since $\max_{\alpha\in K}\|\bm{C}_{k_{\alpha}}\|_{\infty}$ is uniformly bounded, we conclude that sum-product implementation of the Sinkhorn algorithm $O(d({\mathcal{T}})M^{\omega(G)}D^{2}(\lambda+1)^{-1}\epsilon^{-1})$ iterations. Since $d({\mathcal{T}})\leq D$, the complexity is also $O(M^{\omega(G)}D^{3}(\lambda+1)^{-1}\epsilon^{-1})$
Proofs of Section 3
Proofs of Section 3.1
[…]
Lemma 7 (Transformation Identities).
For $h:=\delta_{n}^{-1}(\theta-\theta_{0}-\delta_{n}\Delta_{n,\theta_{0}})$ where $\theta\sim\textrm{q}$, we have
$\textrm{q}(\theta)={|\text{det}(\delta_{n})|}^{-1}\tilde{\textrm{q}}(h),\quad\text{and}\quad\textrm{q}_{i}(\theta)=\delta_{n,ii}^{-1}\tilde{\textrm{q}}_{i}(h),i\in[D],$
Moreover, we have
$\mathbb{H}(\textrm{q})=\mathbb{H}(\tilde{\textrm{q}})-\log|\text{det}(\delta_{n})|,\quad\text{and}\quad\mathbb{H}(\textrm{q}_{i})=\mathbb{H}(\tilde{\textrm{q}}_{i})-\log\delta_{n,ii},$
and
$\Xi(\textrm{q})=\Xi(\tilde{\textrm{q}})+\log|\text{det}(\delta_{n})|-\sum_{i=1}^{D}\log\delta_{n,ii},$
and for any distribution $\textrm{q}_{1},\textrm{q}_{2}$ over $\Theta$, we have
$D_{\text{KL}}(\textrm{q}_{1}\parallel\textrm{q}_{2})=D_{\text{KL}}(\tilde{\textrm{q}}_{1}\parallel\tilde{\textrm{q}}_{2}),$
where $\tilde{\textrm{q}}_{1},\tilde{\textrm{q}}_{2}$ are densities defined via Eq. 3.2.
Proof of Lemma 7.
We obtain the first equality by applying the change of variable formul to Eq. 3.2.
$\textrm{q}(\theta)={|\text{det}(\delta_{n})|}^{-1}\tilde{\textrm{q}}(h),\quad\text{and}\quad\textrm{q}_{i}(\theta)=\delta_{n,ii}^{-1}\tilde{\textrm{q}}_{i}(h),i\in[D],$
For the second equality, we have
[…]
The univariate case follows from this.
For the third equality, we can write
$\Xi(\tilde{\textrm{q}})$ $=\mathbb{H}(\tilde{\textrm{q}})-\sum_{i=1}^{D}\mathbb{H}(\tilde{\textrm{q}}_{i})$
$=\mathbb{H}(\textrm{q})-\sum_{i=1}^{D}\mathbb{H}(\tilde{\textrm{q}}_{i})-\log|\text{det}(\delta_{n})|+\sum_{i=1}^{D}\log\delta_{n,ii}$
$=\Xi(\textrm{q})-\log|\text{det}(\delta_{n})|+\sum_{i=1}^{D}\log\delta_{n,ii}.$
For the fourth equality, we have
[…]
$=D_{\text{KL}}(\textrm{q}_{1}\parallel\textrm{q}_{2}).$
This concludes the proof. ∎
To establish the Bernstein von-Mises theorem, we introduce the tool of $\Gamma$-convergence [14].
Definition 9 ($\Gamma$-Convergence).
Let $X$ be a metric space and consider a set of functionals $F_{\varepsilon}:X\to\mathbb{R}$ indexed by $\varepsilon>0$. A limiting functional $F_{0}$ exists and is called the $\Gamma$-limit of $F_{\varepsilon}$ as $\varepsilon\to 0$, if the following conditions are met:
  1. 1.
    Liminf Inequality: For all $x\in X$ and for every sequence $x_{\varepsilon}\to x$,
    $F_{0}(x)\leq\liminf_{\varepsilon\to 0}F_{\varepsilon}(x_{\varepsilon}).$
  2. 2.
    Limsup Inequality / Existence of a Recovery Sequence: For each $x\in X$, there exists a sequence $\bar{x}_{\varepsilon}\to x$ such that
    $F_{0}(x)\geq\limsup_{\varepsilon\to 0}F_{\varepsilon}(\bar{x}_{\varepsilon}).$
The first condition requires $F_{0}$ to be asymptotically upper bounded by $F_{\varepsilon}$. When paired with the second condition, it ensures that $F_{0}(x)=\lim\limits_{\varepsilon\to 0}F_{\varepsilon}(\bar{x}_{\varepsilon})$, thereby confirming that the lower bound is tight.
Definition 10 (Equi-Coerciveness of Functionals).
A sequence of functionals $F_{\varepsilon}:X\to\mathbb{R}$ is said to be equi-coerciv if for every bounded sequence $x_{\varepsilon}$ with $F_{\varepsilon}(x_{\varepsilon})\leq t$, there exists a subsequence $x_{j}$ of $x_{\epsilon}$ and a converging sequence $x_{j}^{\prime}$ satisfies $F_{\varepsilon_{j}}(x_{j}^{\prime})\leq F_{\varepsilon_{j}}(x_{j})+o(1)$.
Equi-coerciveness ensures the existence of a precompact minimizing sequence for $F_{\varepsilon}$, which helps establish the convergence $x_{\varepsilon}\to x$.
Theorem 10 (Fundamental Theorem of $\Gamma$-Convergence).
Let $X$ be a metric space and $F_{\varepsilon}$ an equi-coercive sequence of functionals. If $F=\Gamma$-$\lim\limits_{\varepsilon\to 0}F_{\varepsilon}$, then
$\arg\min_{x\in X}F=\lim\limits_{\varepsilon\to 0}\arg\min_{x\in X}F_{\varepsilon}.$
This theorem implies that if minimizers $x_{\varepsilon}$ for all $F_{\varepsilon}$ exist, the sequence converges, potentially along a subsequence, to a minimizer of $F$. We note that the converse is not necessarily true; there may exist minimizers for $F$ which are not limits of minimizers for $F_{\varepsilon}$.
Note that when $\delta_{n}=\lambda_{n}^{1/2}\delta_{n}$, we have $|\text{det}(\delta_{n})|=\lambda_{n}^{D/2}|\text{det}(\delta_{n})|$ and $\delta_{n,ii}=\lambda_{n}^{1/2}\delta_{n,ii}$.
We can explicitly characterize the transformed variational posterior:
$\tilde{\textrm{q}}_{\lambda_{n}}(h):=|\text{det}(\delta_{n})|\textrm{q}^{*}_{\lambda_{n}}(\theta_{0}+\delta_{n}h+\delta_{n}\Delta_{n,\theta_{0}}),$ (C.9)
where $\textrm{q}^{*}_{\lambda}$ is the original $\Xi$-variational posterior.
Lemma 8.
Under Definition Eq. C.9, the distribution $\tilde{\textrm{q}}_{\lambda}$ solves the following variational problem
$_{\lambda}=\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathbb{P}_{2}(\Theta)}D_{\text{KL}}(\textrm{q}\parallel\tilde{\textrm{q}}_{0})+\lambda_{n}\Xi(\textrm{q}).$ q ) .
This Lemma is a direct consequence of the transformation identities (Lemma 7) and Eq. 3.1, thus the proof is omitted.
Proof of Theorem 2.
WLOG, we assume that $\Theta=\mathbb{R}^{D}$. Otherwise, we use the same proof by adding an indicator of the minimizing set to the sequence of functionals.
Regime 1: $\lambda_{n}\to\infty$. It suffices to show
$F_{n}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel\tilde{\textrm{q}}_{0})+\lambda_{n}\Xi(\textrm{q}),$
$\Gamma$-converge to
$F_{0}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\infty\Xi(\textrm{q}),$
in $[P_{\theta_{0}}]$-probability as $n\to\infty$.
By Theorem 10, $\Gamma$ convergence implies $W_{2}(\tilde{\textrm{q}}_{\lambda_{n}},\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathbb{P}_{2}(\Theta)}F_{0}(\textrm{q}))\overset{P_{\theta_{0}}}{\to}0$, where $\textrm{q}_{0}$ is the minimizer of $F_{0}$.
To prove the $\Gamma$-convergence, we rewrite $F_{n}$.
$F_{n}(\textrm{q})$ $:=D_{\text{KL}}(\textrm{q}\parallel\tilde{\textrm{q}}_{0})+\lambda_{n}\Xi(\textrm{q})$
$=\mathbb{E}_{\textrm{q}}\left[-\ell\left(\bm{x}^{(n)};\theta_{0}+\delta_{n}h+\delta_{n}\Delta_{n,\theta_{0}}\right)\right]+D_{\text{KL}}(\textrm{q}\parallel\tilde{\pi})+\log|\text{det}(\delta_{n})|+\lambda_{n}\Xi(\textrm{q})$
[…]
$=-\ell(\bm{x}^{(n)};\theta_{0})+\mathbb{E}_{\textrm{q}}\left[\frac{1}{2}h^{T}V_{\theta_{0}}h\right]+\mathbb{H}(\textrm{q})-\mathbb{E}_{\textrm{q}}\left[\log\pi(\theta_{0}+\delta_{n}h+\delta_{n}\Delta_{n,\theta_{0}})\right]+\lambda_{n}\Xi(\textrm{q})$
[…]
Applying LAN expansion and Laplace approximation to the log-normalizer, we have
[…]
$=\frac{D}{2}\log 2\pi-\frac{1}{2}\log\text{det}(V_{\theta_{0}})+\log\pi(\theta_{0})+\ell(\bm{x}^{(n)};\theta_{0})+o_{P}(1).$
After cancellation, we have
$F_{n}(\textrm{q})$ $=\mathbb{E}_{\textrm{q}}\left[\frac{1}{2}h^{T}V_{\theta_{0}}h\right]+\mathbb{H}(\textrm{q})+\lambda_{n}\Xi(\textrm{q})+(\frac{D}{2}\log 2\pi-\frac{1}{2}\log\text{det}(V_{\theta_{0}}))$
$-\left\{\mathbb{E}_{\textrm{q}}\left[\log\pi(\theta_{0}+\delta_{n}h+\delta_{n}\Delta_{n,\theta_{0}})\right]-\log\pi(\theta_{0})\right\}+o_{P}(1).$
Using Assumption 1 to bound the prior tail via Taylor expansion, we have an expression for $F_{n}$
$F_{n}(\textrm{q})$ $=\mathbb{E}_{\textrm{q}}\left[\frac{1}{2}h^{T}V_{\theta_{0}}h\right]+\mathbb{H}(\textrm{q})+\lambda_{n}\Xi(\textrm{q})+\frac{D}{2}\log 2\pi-\frac{1}{2}\log\text{det}(V_{\theta_{0}})+o_{P}(1).$ (C.10)
$=D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q})+o_{P}(1).$
Now we rewrite $F_{0}(\textrm{q})$.
$F_{0}(\textrm{q})$ $:=D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\infty\Xi(\textrm{q})$
$=\mathbb{E}_{\textrm{q}}\left[\frac{1}{2}h^{T}V_{\theta_{0}}h\right]+\mathbb{H}(\textrm{q})+\infty\Xi(\textrm{q})+\frac{D}{2}\log 2\pi-\frac{1}{2}\log\text{det}(V_{\theta_{0}}).$
$=F_{n}(\textrm{q})+\infty\Xi(\textrm{q})+o_{P}(1).$
Now we prove the $\Gamma$ convergence.
First, we verify the liminf inequality. Let $_{n}\overset{W_{2}}{\to}\textrm{q}$ q. When q is not mean-field, we have:
$\liminf_{n\to\infty}F_{n}(\textrm{q}_{n})$ $\geq\liminf_{n\to\infty}\left\{D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q}_{n})\right\}-\epsilon$
$\geq\liminf_{n\to\infty}D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))+\liminf_{n\to\infty}\lambda_{n}\liminf_{n\to\infty}\Xi(\textrm{q}_{n})-\epsilon$
$\geq D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\infty\Xi(\textrm{q})-\epsilon=\infty\geq F_{0}(\textrm{q}).$
The second inequality follows from the definition of liminf. The third line is due to Lemma 5, which states that the KL functional and $\Xi$ functional are lower semicontinuous.
$\liminf_{n\to\infty}F_{n}(\textrm{q}_{n})$ $\geq\liminf_{n\to\infty}\left\{D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q}_{n})\right\}-\epsilon$
$\geq\liminf_{n\to\infty}D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))-\epsilon$
$\geq D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))-\epsilon.$
Since this holds for all $\epsilon$, we verified that $\liminf_{n\to\infty}F_{n}(\textrm{q}_{n})\geq F_{0}(\textrm{q})$.
Next, we show the existence of a recovery sequence. When q is not mean-field, $F_{0}(\textrm{q})=+\infty$, and the limsup inequality is automatically satisfied. When q is mean-field, choose $_{n}:=\textrm{q}$ q, then:
$\limsup_{n\to\infty}F_{n}(\textrm{q}_{n})=\limsup_{n\to\infty}D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+o_{P}(1)\leq F_{0}(\textrm{q}).$
Thus, $F_{0}$ is the $\Gamma$-limit of the sequence $F_{n}$.
Next we prove that the sequence $F_{n}$ is eqi-coercive. Take $n_{j}\to\infty$ and $\textrm{q}_{n_{j}}$ such that $F_{n_{j}}(\textrm{q}_{n_{j}})\leq t$ for all $j$. Then $\lambda_{n_{j}}\Xi(\textrm{q}_{n_{j}})$ is bounded as $\lambda_{n_{j}}\to\infty$, thus $\Xi(\textrm{q}_{n_{j}})=o(1)$. Using this and Eq. C.10, we have
$D_{\text{KL}}\left(\textrm{q}_{n_{j}}\parallel N(0,V_{\theta_{0}}^{-1})\right)\leq t+1,\quad\text{for sufficiently large }j.$
Since $D_{\text{KL}}\left(.\parallel N(0,V_{\theta_{0}}^{-1})\right)$ is a Wasserstein (geodastically) convex functional, it is coercive by Lemma 2.4.8 of [6]. This implies that the set $\{\textrm{q}\in\mathbb{P}_{2}(\Theta)\mid D_{\text{KL}}\left(\textrm{q}\parallel\tilde{\textrm{q}}_{0}\right)\leq t+1\}$ is compact under the Wasserstein metric, thus $\textrm{q}_{n_{j}}$ has a subsequence $\textrm{q}_{n_{j}}^{\prime}$ that converges to $\textrm{q}^{*}$ in the Wasserstein metric of and $D_{\text{KL}}\left(\textrm{q}^{*}\parallel\tilde{\textrm{q}}_{0}\right)\leq t+1$. Thus we have $F_{n_{j}}(\textrm{q}_{n_{j}}^{\prime})\leq F_{n_{j}}(\textrm{q}_{n_{j}}^{\prime})+o(1)$ by Eq. C.10 where $\textrm{q}_{n_{j}}^{\prime}$ is a converging subsequence of $\textrm{q}_{n_{j}}$. This verifies the equi-coercivity of $F_{n}$.
Lastly, we note that $F_{0}$ attains its minimum at $N(\Delta_{n,\theta_{0}},V_{\theta_{0}}^{{}^{\prime}-1})$ where $V_{\theta_{0}}^{{}^{\prime}-1}$ is the MFVI covariance. As a result of Theorem 10, we conclude that the desired convergence takes place:
$D_{\text{KL}}\left(\textrm{q}_{n_{j}}\parallel N(0,V_{\theta_{0}}^{-1})\right)\leq t+1,\quad\text{for sufficiently large }j.$
Regime 2: $\lambda_{n}\to 0$.
In this regime, we show that the functionals
$F_{n}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel\tilde{\textrm{q}}_{0})+\lambda_{n}\Xi(\textrm{q}),$
$\Gamma$-converge to
$F_{0}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1})),$
in $[P_{\theta_{0}}]$-probability as $n\to\infty$.
Given that $F_{n}$ is defined analogous to Regime 1, we will skip the derivation:
$F_{n}(\textrm{q})$ $=\mathbb{E}_{\textrm{q}}\left[\frac{1}{2}h^{T}V_{\theta_{0}}h\right]+\mathbb{H}(\textrm{q})+\lambda_{n}\Xi(\textrm{q})+\frac{D}{2}\log 2\pi-\frac{1}{2}\log\text{det}(V_{\theta_{0}})+o_{P}(1).$
$=D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q})+o_{P}(1).$
Now we prove the $\Gamma$ convergence. First, we verify the liminf inequality. Let $_{n}\overset{W_{2}}{\to}\textrm{q}$ q. We have:
$\liminf_{n\to\infty}F_{n}(\textrm{q}_{n})$ $\geq\liminf_{n\to\infty}\left\{D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q}_{n})\right\}-\epsilon$
$\geq\liminf_{n\to\infty}D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))+\liminf_{n\to\infty}\lambda_{n}\liminf_{n\to\infty}\Xi(\textrm{q}_{n})-\epsilon$
$\geq D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))-\epsilon.$
Since this holds for all $\epsilon$, we verified that $\liminf_{n\to\infty}F_{n}(\textrm{q}_{n})\geq F_{0}(\textrm{q})$.
For the recovery sequence, we take $_{n}:=\textrm{q}$ q. Since q is absolutely continuous with respect to the product of its marginals, $\Xi(\textrm{q})$ is finite. Then we have:
$\limsup_{n\to\infty}F_{n}(\textrm{q}_{n})=\limsup_{n\to\infty}D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+o_{P}(1)\leq F_{0}(\textrm{q}).$
The equicoercivity of $F_{n}$ follows from the argument in regime 1. By Theorem 10, we conclude with the desired convergence:
$W_{2}(\tilde{\textrm{q}}_{\lambda_{n}},N(0,V_{\theta_{0}}^{{}^{\prime}-1}))\to 0.$
Regime 3: $\lambda_{n}\to\lambda_{\infty}\in(0,\infty)$.
In this regime, we show that the functionals
$F_{n}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel\tilde{\textrm{q}}_{0})+\lambda_{n}\Xi(\textrm{q}),$
$\Gamma$-converge to
$F_{0}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{\infty}\Xi(\textrm{q}),$
in $[P_{\theta_{0}}]$-probability as $n\to\infty$.
Recall that
$F_{n}(\textrm{q})$ $=\mathbb{E}_{\textrm{q}}\left[\frac{1}{2}h^{T}V_{\theta_{0}}h\right]+\mathbb{H}(\textrm{q})+\lambda_{n}\Xi(\textrm{q})+\frac{D}{2}\log 2\pi-\frac{1}{2}\log\text{det}(V_{\theta_{0}})+o_{P}(1).$
$=D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q})+o_{P}(1).$
Now we prove the $\Gamma$ convergence. First, we verify the liminf inequality. Let $_{n}\overset{W_{2}}{\to}\textrm{q}$ q. We have:
$\liminf_{n\to\infty}F_{n}(\textrm{q}_{n})$ $\geq\liminf_{n\to\infty}\left\{D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q}_{n})\right\}-\epsilon$
$\geq\liminf_{n\to\infty}D_{\text{KL}}(\textrm{q}_{n}\parallel N(0,V_{\theta_{0}}^{-1}))+\liminf_{n\to\infty}\lambda_{n}\liminf_{n\to\infty}\Xi(\textrm{q}_{n})-\epsilon$
$\geq D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{\infty}\Xi(\textrm{q})-\epsilon.$
The second inequality follows from the definition of liminf, and the last inequality is due to Lemma 5, which states that the KL functional and $\Xi$ functional are lower semicontinuous.
For the recovery sequence, we take $_{n}:=\textrm{q}$ q. As long as $\Xi(\textrm{q})$ is finite, we have:
$\limsup_{n\to\infty}F_{n}(\textrm{q}_{n})=\limsup_{n\to\infty}D_{\text{KL}}(\textrm{q}\parallel N(0,V_{\theta_{0}}^{-1}))+\lambda_{n}\Xi(\textrm{q})+o_{P}(1)=F_{0}(\textrm{q}).$
The equicoercivity of $F_{n}$ follows from the argument in regime 1. By Theorem 10, we have the convergence:
$W_{2}(\tilde{\textrm{q}}_{\lambda_{n}},\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathbb{P}_{2}(\Theta)}F_{0}(\textrm{q}))\to 0.$
Proof of Corollary 2.
Recall the definition of Wasserstein distance:
$W_{2}(\textrm{p},\textrm{q})=(\inf_{\pi\in\mathcal{C}(\textrm{p},\textrm{q})}\mathbb{E}_{\pi}[\|X-Y\|^{2}])^{1/2}.$
Given the change of variable definition (Eq. 3.2), we have
$W_{2}(\tilde{\textrm{q}}_{\lambda_{n}},N(\mu,\Sigma))$ $=(\inf_{\pi\in\mathcal{C}(\tilde{\textrm{q}}_{\lambda_{n}},N(\mu,\Sigma))}\mathbb{E}_{\pi}[\|h-h^{\prime}\|^{2}])^{1/2}$
$=|\text{det}(\delta_{n})|^{-1}(\inf_{\pi\in\mathcal{C}(\textrm{q}^{*}_{\lambda_{n}},N(\delta_{n}\mu+\theta_{0}+\delta_{n}\Delta_{n,\theta_{0}},\delta_{n}^{T}\Sigma\delta_{n}))}\mathbb{E}_{\pi}[\|\theta-\theta^{\prime}\|^{2}])^{1/2}$
$=|\text{det}(\delta_{n})|^{-1}W_{2}(\textrm{q}^{*}_{\lambda_{n}},N(\delta_{n}\mu+\theta_{0}+\delta_{n}\Delta_{n,\theta_{0}},\delta_{n}^{T}\Sigma\delta_{n})).$
If $W_{2}(\tilde{\textrm{q}}_{\lambda_{n}},N(\mu,\Sigma))$ tends to 00, then $W_{2}(\textrm{q}^{*}_{\lambda_{n}},N(\delta_{n}\mu+\theta_{0}+\delta_{n}\Delta_{n,\theta_{0}},\delta_{n}^{T}\Sigma\delta_{n}))$ tends to 00. Since $N(\delta_{n}\mu+\theta_{0}+\delta_{n}\Delta_{n,\theta_{0}},\delta_{n}^{T}\Sigma\delta_{n})$ weakly converge to $\delta_{\theta_{0}}$, it converges to $\delta_{\theta_{0}}$ in Wasserstein metric. By Theorem 2, we have $\textrm{q}^{*}_{\lambda_{n}}$ converges in Wasserstein metric to $\delta_{\theta_{0}}$, as desired. ∎
Proofs of Section 3.2
We first prove a useful proposition.
Proposition 4 (Optimality to fixed point).
Let Assumption 4 hold. Let $m^{*}_{\lambda}(\theta)=\prod_{i=1}^{D}m^{*}_{\lambda,i}(\theta_{i})$ be the product of optimal marginals, and $\phi^{*}_{\lambda}$ be the optimal EOT potentials. Then $m^{*}_{\lambda}$ and $\phi^{*}_{\lambda}$ satisfy the fixed point equations:
$m^{*}_{\lambda,i}(\theta_{i})$ $=Z_{i}^{-1}\exp(-(\lambda+1)\phi^{*}_{\lambda,i}(\theta_{i}))\pi_{i}(\theta_{i}),\quad\text{and }\quad$ (C.11)
$\hat{\phi}_{\lambda,i}(\theta_{i})$ $=-\log\int_{\Theta_{-i}}\exp\left(\frac{1}{\lambda+1}\ell\left(\bm{x}^{(n)};\theta\right)-\lambda\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})\right)\prod_{j\neq i}\pi_{j}(\theta_{j})d\theta_{-i}$
[…]
where $Z_{i}$’s are the normalizing constants.
Proof of Proposition 4.
Define $f_{\lambda}(\theta)$ as follows,
$f_{\lambda}(\theta):=-(\lambda+1)\left(\sum_{i=1}^{D}\phi^{*}_{\lambda,i}(\theta_{i})\right)\exp\left(\sum_{i=1}^{D}\phi^{*}_{\lambda,i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x}^{(n)};\theta\right)\right)+\sum_{i=1}^{D}\log\pi_{i}(\theta_{i}).$ (C.12)
By Theorem 1, Assumption 4 and the uniform boundedness of $\ell\left(\bm{x}^{(n)};\cdot\right)$, the function $f_{\lambda}$ is integrable with respect to $m^{*}_{\lambda}$. From the derivation in Section 2.2, the distribution $m^{*}_{\lambda}$ attain the minimum,
$\min\limits_{m\in\mathcal{M}(\Theta)}-\mathbb{E}_{m}\left[f_{\lambda}(\theta)\right]+\mathbb{H}(m).$ (C.13)
Define $\hat{f}_{\lambda,i}(\theta_{i})=\mathbb{E}_{m^{*}_{\lambda}}[f_{\lambda}(\theta)\mid\theta_{i}]$ for $[m^{*}_{\lambda,i}]$-a.s. $\theta_{i}\in\Theta_{i}$. Since $m\in\mathcal{M}(\Theta)$, $\mathbb{H}(m)=\sum_{i=1}^{D}\mathbb{H}(m_{i})$. From this and the tower property, we have that for $i\in[D]$, $m^{*}_{\lambda,i}$ attains the minimum.
$m^{*}_{\lambda,i}(\theta_{i}):=\mathop{\mathrm{argmin}}_{m_{i}\in\mathcal{M}(\Theta_{i})}\left(\mathbb{E}_{m_{i}}[-\hat{f}_{\lambda,i}(\theta_{i})]+\mathbb{H}(m_{i})\right).$ (C.14)
By the Gibbs variational principle (Lemma 1), the minimum is uniquely attained by
$m^{*}_{\lambda,i}(\theta_{i})\propto\exp(\hat{f}_{\lambda,i}(\theta_{i})).$ (C.15)
Recall that the optimal EOT potentials $\phi^{*}_{\lambda,i}$’s satisfy the Schrödinger system, which means for $i\in[D]$,
$\hat{\phi}_{\lambda,i}(\theta_{i})=-\log\int_{\Theta_{-i}}\exp\left(\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})+\frac{1}{\lambda+1}\ell\left(\bm{x}^{(n)};\theta\right)\right)\prod_{j\neq i}m^{*}_{\lambda,j}(\theta_{j})d\theta_{-i}.$ (C.16)
This allows us to simplify $\hat{f}_{\lambda,i}(\theta_{i})$:
$\hat{f}_{\lambda,i}(\theta_{i})=\mathbb{E}_{m^{*}_{\lambda}}[f_{\lambda}(\theta)\mid\theta_{i}]$
$=\mathbb{E}_{m^{*}_{\lambda}}\left[-(\lambda+1)\left(\sum_{i=1}^{D}\phi^{*}_{\lambda,i}(\theta_{i})\right)\exp\left(\sum_{i=1}^{D}\phi^{*}_{\lambda,i}(\theta_{i})+\frac{1}{\lambda+1}\ell\left(\bm{x}^{(n)};\theta\right)\right)+\sum_{i=1}^{D}\log\pi_{i}(\theta_{i})\mid\theta_{i}\right]$
$=-(\lambda+1)\frac{\mathbb{E}_{m^{*}_{\lambda,-i}}\left[\left(\sum_{i=1}^{D}\phi^{*}_{\lambda,i}(\theta_{i})\right)\exp\left(\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})+\frac{1}{\lambda+1}\ell\left(\bm{x}^{(n)};\theta\right)\right)\right]}{\exp\left(-\phi^{*}_{\lambda,i}(\theta_{i})\right)}$
$+\mathbb{E}_{m^{*}_{\lambda,-i}}\left[\sum_{j\neq i}\log\pi_{j}(\theta_{j})\right]+\log\pi_{i}(\theta_{i})$
$=-(\lambda+1)\phi^{*}_{\lambda,i}(\theta_{i})+\log\pi(\theta_{i})-(\lambda+1)\mathbb{E}_{\hat{h}_{\lambda}}\left[\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})\right]+C.$
where $\hat{h}_{\lambda}(\theta_{-i})\propto\exp\left(\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})+\frac{1}{\lambda+1}\ell\left(\bm{x}^{(n)};\theta\right)\right)\prod_{j\neq i}m^{*}_{\lambda,i}(\theta_{i})$. Since $\hat{h}_{\lambda}(\theta_{-i})\propto\textrm{q}^{*}_{\lambda}(\theta_{-i},\theta_{i})$, we have for all $\theta_{i}$,
$\mathbb{E}_{\hat{h}_{\lambda}(\theta_{-i})}\left[\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})\right]=\mathbb{E}_{\textrm{q}^{*}_{\lambda}(\theta_{-i},\theta_{i})}\left[\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})\right]=\sum_{j\neq i}\mathbb{E}_{m^{*}_{\lambda,i}}\left[\phi^{*}_{\lambda,j}(\theta_{j})\right].$ (C.17)
The last equality uses the fact that $m^{*}_{\lambda,i}$ is the $i^{\text{th}}$ marginal of $\textrm{q}^{*}_{\lambda}$.
Since $\sum_{j\neq i}\mathbb{E}_{m^{*}_{\lambda,j}}\left[-(\lambda+1)\phi^{*}_{\lambda,j}(\theta_{j})+\log\pi_{j}(\theta_{j})\right]$ does not depend on $\theta_{i}$, we obtain
$m^{*}_{\lambda,i}(\theta_{i})\propto\exp(-(\lambda+1)\phi^{*}_{\lambda,i}(\theta_{i}))\pi_{i}(\theta_{i}).$ (C.18)
Using Eq. C.18, we conclude
$\hat{\phi}_{\lambda,i}(\theta_{i})$ $=-\log\int_{\Theta_{-i}}\exp\left(\frac{1}{\lambda+1}\ell\left(\bm{x}^{(n)};\theta\right)-\lambda\sum_{j\neq i}\phi^{*}_{\lambda,j}(\theta_{j})\right)\prod_{j\neq i}\pi_{j}(\theta_{j})d\theta_{-i}$
[…]
Proof of Theorem 3.
We define constants $u_{i},v_{i},$ and $w$ based on the partial derivatives of the log-likelihood $\ell(\bm{x}^{(n)};\theta)$. Specifically, $u_{i}:=\frac{1}{2}\inf_{\theta\in\Theta}\partial_{ii}\ell(\bm{x}^{(n)};\theta)$, $v_{i}$ is chosen such that $\sup_{\theta\in\Theta}|\partial_{i}\ell(\bm{x}^{(n)};\theta)-v_{i}-2u_{i}\theta_{i}|=b_{i}$, and $w$ is chosen such that $\sup_{\theta\in\Theta}|\ell(\bm{x}^{(n)};\theta)-w-\sum_{i}v_{i}\theta_{i}-\sum_{i}u_{i}\theta_{i}^{2}|=a$. These terms, as constants in $\mathbb{R}$, are well-defined under Assumption 5. We define a new log-likelihood ~$\tilde{\ell}(\bm{x}^{(n)};\theta)$ that shift $\ell(\bm{x}^{(n)};\theta)$ by a quadratic function:
~$\tilde{\ell}(\bm{x}^{(n)};\theta)$ $:=\ell(\bm{x}^{(n)};\theta)-w-\sum_{i=1}^{D}v_{i}\theta_{i}-\sum_{i=1}^{D}u_{i}\theta_{i}^{2}.$
Calculation yields that
$\sup_{\theta\in\Theta}|\tilde{\ell}(\bm{x}^{(n)};\theta)|$ $=a,\quad\sup_{\theta\in\Theta}|\partial_{i}\tilde{\ell}(\bm{x}^{(n)};\theta)|=b_{i},\quad\sup_{\theta\in\Theta}|\partial_{ii}\tilde{\ell}(\bm{x}^{(n)};\theta)|=c_{i}.$
Given the optimal $m^{*}_{\lambda_{n}}$, the inner variational (EOT) problem has the following formulation.
$\textrm{q}^{*}_{\lambda_{n}}=\mathop{\mathrm{argmin}}_{\textrm{q}\in\mathcal{C}(m^{*}_{\lambda_{n}})}-\mathbb{E}_{\textrm{q}}[\tilde{\ell}(\bm{x}^{(n)};\theta)]+(\lambda_{n}+1)D_{\text{KL}}(\textrm{q}\parallel m^{*}_{\lambda_{n}}),$
where we use ~$\tilde{\ell}(\bm{x}^{(n)};\theta)$ instead of $\ell(\bm{x}^{(n)};\theta)$ because the subtracting a tensorized function $w+\sum_{i=1}^{D}v_{i}\theta_{i}+\sum_{i=1}^{D}u_{i}\theta_{i}^{2}$ from the cost does not change the optimal EOT coupling.
By Theorem 1, we can write $\textrm{q}^{*}_{\lambda_{n}}$ using the EOT solution structure.
$\textrm{q}^{*}_{\lambda_{n}}(\theta)=\exp\left(\frac{1}{\lambda_{n}+1}\tilde{\ell}(\bm{x}^{(n)};\theta)+\sum_{i=1}^{D}\phi^{*}_{\lambda_{n},i}(\theta_{i})\right)\prod_{i=1}^{D}m^{*}_{\lambda_{n},i}(\theta)$
where $m^{*}_{\lambda_{n},i}$’s are the marginals of $\textrm{q}^{*}_{\lambda_{n}}$ and $\phi^{*}_{\lambda_{n},i}$’s are the EOT potentials.
Define another product distribution $\tilde{m}_{\lambda_{n}}(\theta)\propto\exp\left(\sum_{i=1}^{D}\phi^{*}_{\lambda_{n},i}(\theta_{i})\right)m^{*}_{\lambda_{n},i}(\theta)$. We can rewrite $\textrm{q}^{*}_{\lambda_{n}}$ as the product of a tempered likelihood and a $\tilde{m}_{\lambda_{n}}$.
$\textrm{q}^{*}_{\lambda_{n}}(\theta)=\frac{1}{\mathcal{Z}_{D}(\lambda_{n})}\exp\left(\frac{1}{\lambda_{n}+1}\tilde{\ell}(\bm{x}^{(n)};\theta)\right)\tilde{m}_{\lambda_{n}}(\theta),$ (C.19)
where the normalizing constant is given by
[…]
First, we want to show that
$\lim\limits_{n\to\infty}\frac{1}{D}\left[\log\mathcal{Z}_{D}(\lambda_{n})-\sup_{m\in\mathcal{M}(\Theta),m\ll\tilde{m}_{\lambda_{n}}}\left\{\frac{1}{\lambda_{n}+1}\mathbb{E}_{m}[\tilde{\ell}(\bm{x}^{(n)};\theta)]-\sum_{i=1}^{D}D_{\text{KL}}(m_{i}\parallel\tilde{m}_{\lambda_{n},i})\right\}\right]=0.$ (C.21)
Let $\|f\|_{\infty}$ denote the supremum norm of a function $f$. Fix some $\epsilon>0$. Let ${\mathcal{S}}_{\lambda_{n}}(\epsilon)\subset\Theta$ be a finite set such that for any $\theta\in\Theta$, there exists $s\in{\mathcal{S}}_{\lambda_{n}}(\epsilon)$ satisfying
$\sum_{i=1}^{D}\|\frac{1}{\lambda_{n}+1}\partial_{i}\tilde{\ell}(\bm{x}^{(n)};\theta)-s_{i}\|_{\infty}^{2}\leq\epsilon^{2}D.$ (C.22)
Denote by $|{\mathcal{S}}_{\lambda_{n}}(\epsilon)|$ the cardinality of ${\mathcal{S}}_{\lambda_{n}}(\epsilon)$. Theorem 1.1 of [75] implies that
$\log\mathcal{Z}_{D}(\lambda_{n})-\sup_{m\in\mathcal{M}(\Theta),m\ll\tilde{m}_{\lambda_{n}}}\left[\frac{1}{\lambda_{n}+1}\mathbb{E}_{m}[\tilde{\ell}(\bm{x}^{(n)};\theta)]-\sum_{i=1}^{D}D_{\text{KL}}(m_{i}\parallel\tilde{m}_{\lambda_{n},i})\right]\leq$ (C.23)
$4\left(\frac{4}{(\lambda_{n}+1)^{2}}\left(a\sum_{i=1}^{D}c_{ii}+\sum_{i=1}^{D}b_{i}^{2}\right)+\frac{8}{(\lambda_{n}+1)^{2}}\sum_{i=1}^{D}\sum_{j=1}^{D}b_{i}c_{ij}+\frac{16}{(\lambda_{n}+1)^{3/2}}\left(a\sum_{i=1}^{D}\sum_{j=1}^{D}c_{ij}^{2}+\sum_{i=1}^{D}\sum_{j=1}^{D}b_{i}b_{j}c_{ij}\right)\right)^{1/2}$
$+4\left(\frac{1}{(\lambda_{n}+1)^{2}}\sum_{i=1}^{D}b_{i}^{2}+\epsilon^{2}D\right)^{1/2}\left(\frac{8}{\lambda_{n}+1}\left(\sum_{i=1}^{D}c_{ii}^{2}\right)^{1/2}+4D^{1/2}\epsilon\right)+\frac{4}{\lambda_{n}+1}\sum_{i=1}^{D}c_{ii}+2D\epsilon$
$+\log 2+\log|{\mathcal{S}}_{\lambda_{n}}(\epsilon)|.$
Consider $\lambda_{n}\succ D^{-1/2}\max\left(\sqrt{a\sum_{i=1}^{D}c_{ii}},\sqrt{\sum_{i=1}^{D}b_{i}^{2}},\sqrt{\sum_{i=1}^{D}\sum_{j=1}^{D}c_{ij}^{2}},D^{1/2}\right)$. Then,
$\frac{\sum_{i=1}^{D}\sum_{j=1}^{D}b_{i}c_{ij}}{(\lambda_{n}+1)^{2}}\leq\frac{\sqrt{\sum_{i=1}^{D}b_{i}^{2}}\sqrt{\sum_{i=1}^{D}\sum_{j=1}^{D}c_{ij}^{2}}}{(\lambda_{n}+1)^{2}}=o(D),$
$\frac{\sum_{i=1}^{D}\sum_{j=1}^{D}b_{i}b_{j}c_{ij}}{(\lambda_{n}+1)^{3}}\leq\frac{\sqrt{\sum_{i=1}^{D}\sum_{j=1}^{D}c_{ij}^{2}}\sum_{i=1}^{D}b_{i}^{2}}{(\lambda_{n}+1)^{3}}=o(D^{3/2}).$
by the Cauchy-Schwarz inequality. With the other terms being $o(D)$, we have
$\log\mathcal{Z}_{D}(\lambda_{n})-\sup_{m\in\mathcal{M}(\Theta),m\ll\tilde{m}_{\lambda_{n}}}\left[\frac{1}{\lambda_{n}+1}\mathbb{E}_{m}[\tilde{\ell}(\bm{x}^{(n)};\theta)]-\sum_{i=1}^{D}D_{\text{KL}}(m_{i}\parallel\tilde{m}_{\lambda_{n},i})\right]$ (C.24)
$\leq o(D)+2D\epsilon+\log 2+\log|{\mathcal{S}}_{\lambda_{n}}(\epsilon)|.$
To upper bound $|{\mathcal{S}}_{\lambda_{n}}(\epsilon)|$, we can construct an $\epsilon$-covering by covering $[-\frac{b_{i}}{\lambda_{n}+1},\frac{b_{i}}{\lambda_{n}+1}]$ with balls of size $2\epsilon$. We consider a candidate set $\tilde{\mathcal{S}}_{\lambda_{n}}(\epsilon)$ as the product of these coverings. Since $|\tilde{\mathcal{S}}_{\lambda_{n}}(\epsilon)|=\frac{\prod_{i=1}^{D}b_{i}}{(\lambda_{n}+1)^{D}\epsilon^{D}}$, we have
$\log|{\mathcal{S}}_{\lambda_{n}}(\epsilon)|\leq\sum_{i=1}^{D}\log b_{i}-D\log(\lambda_{n}+1)-D\log\epsilon.$
Define $\bar{b}:=\sum_{i=1}^{D}b_{i}$. Since $\lambda_{n}\succ D^{-1/2}\sqrt{\sum_{i=1}^{D}b_{i}^{2}}$, we have $D\lambda_{n}\succ D^{1/2}\sqrt{\sum_{i=1}^{D}b_{i}^{2}}\geq D\bar{b}$. By Jensen’s inequality,
$D\log(\lambda_{n}+1)\succ D\log(\bar{b}+1)\geq\sum_{i=1}^{D}\log(b_{i}).$
To complete the bound of $\log|{\mathcal{S}}_{\lambda_{n}}(\epsilon_{n})|$, we take $\epsilon_{n}:=\sqrt{\frac{\bar{b}+1}{\lambda_{n}+1}}$. The inequality above shows that $\epsilon_{n}=o(1)$.
Then we have
$\log|{\mathcal{S}}_{\lambda_{n}}(\epsilon_{n})|$ $\leq D\log(\bar{b}+1)-D\log(\lambda_{n}+1)-D\log\epsilon_{n}$ (C.25)
$=\frac{1}{2}D\log\left(\frac{\bar{b}+1}{\lambda_{n}+1}\right)\to-\infty.$
Plugging the definition of $\epsilon_{n}$ into Eq. C.24, we get
$\log\mathcal{Z}_{D}(\lambda_{n})-\sup_{m\in\mathcal{M}(\Theta),m\ll\tilde{m}_{\lambda_{n}}}\left[\frac{1}{\lambda_{n}+1}\mathbb{E}_{m}[\tilde{\ell}(\bm{x}^{(n)};\theta)]-\sum_{i=1}^{D}D_{\text{KL}}(m_{i}\parallel\tilde{m}_{\lambda_{n},i})\right]=o(D).$ (C.26)
For any $m\in\mathcal{M}(\Theta)$, we have
[…]
$=\log\mathcal{Z}_{D}(\lambda_{n})-\frac{1}{\lambda_{n}+1}\mathbb{E}_{m}[\tilde{\ell}(\bm{x}^{(n)};\theta)]+\sum_{i=1}^{D}D_{\text{KL}}(m_{i}\parallel\tilde{m}_{\lambda_{n},i}(\theta_{i})).$
Result Eq. C.24 implies that
$\inf_{m\in\mathcal{M}(\Theta)}D_{\text{KL}}(m\parallel\textrm{q}^{*}_{\lambda_{n}})=o(D).$ (C.27)
For any $1$-Lipschitz function $f$ under the $L_{1}$ norm, consider the random variable $f(\theta)$, where $\theta\sim\textrm{q}^{*}_{\lambda_{n}}$. This variable satisfies the inequality $\log\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\exp(\langle t,f(\theta)-\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[f(\theta)]\rangle)]\leq 2D\|t\|_{2}^{2}$, which is derived from the assumption that $\Theta=[-1,1]^{D}$. Thus, $\textrm{q}^{*}_{\lambda_{n}}$ is $4D$-subGaussian. By the $T_{1}$-transportation inequality (Theorem 4.8, [66]), for any $m\in\mathcal{M}(\Theta)$, the following upper bound holds:
$W_{1}(m,\textrm{q}^{*}_{\lambda_{n}})\leq\sqrt{8D\cdot D_{\text{KL}}(m\parallel\textrm{q}^{*}_{\lambda_{n}})}=o(D),$ (C.28)
where $W_{1}$ is the $1-$Wasserstein distance defined with respect to the $L_{1}$ norm.
Let $m_{\lambda_{n}}^{*}$ denote the minimizer of the left hand side Eq. C.27. Consider a function $\psi$ that is $1$-Lipschitz on $\mathbb{R}$. The function $\theta\mapsto\sum_{i=1}^{D}\psi(\theta_{i})$ is also $1$-Lipschitz with respect to the $L_{1}$ norm. This follows from the inequality:
$|\sum_{i=1}^{D}\psi(\theta_{i})-\sum_{i=1}^{D}\psi(\theta_{i}^{\prime})|\leq\sum_{i=1}^{D}|\psi(\theta_{i})-\psi(\theta_{i}^{\prime})|\leq\sum_{i=1}^{D}|\theta_{i}-\theta^{\prime}_{i}|\leq\|\theta-\theta^{\prime}\|_{1}.$ (C.29)
Applying Kantorovich duality, we obtain the bound:
$\sup_{\psi\in\text{Lip}(\mathbb{R})}\left|\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{m_{\lambda_{n}}^{*}}[\psi(\theta_{i})]\right)\right|\leq W_{1}(\textrm{q}^{*}_{\lambda_{n}},m_{\lambda_{n}}^{*})=o(D).$ (C.30)
Since the bound in Eq. C.23 does not depend on the value of $\bm{x}^{(n)}$, we have
$\sup_{\bm{x}^{(n)}\in\mathbb{X}^{n}}\left|\frac{1}{D}\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{m_{\lambda_{n}}^{*}}[\psi(\theta_{i})]\right)\right|\leq\frac{1}{D}W_{1}(\textrm{q}^{*}_{\lambda_{n}},m_{\lambda_{n}}^{*})\overset{P_{\theta_{0}}}{\to}0.$ (C.31)
Consider the regime $\lambda_{n}\prec D\Xi^{-1}(\textrm{q}^{*}_{0})$. Recall that the the $\Xi$-VI has the Lagrangian formulation as $\min_{\Xi(\textrm{q})\leq t(\lambda_{n})}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})$ for some constant $t(\lambda_{n})$ depending on $\lambda_{n}$. If $\Xi(\textrm{q}^{*}_{0})\leq t(\lambda_{n})$, then $\textrm{q}^{*}_{\lambda_{n}}=\textrm{q}^{*}_{0}$, which implies $\Xi(\textrm{q}^{*}_{0})\geq\Xi(\textrm{q}^{*}_{\lambda_{n}})$ for all $\lambda_{n}$. For fixed $n$, we have
$D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel\textrm{q}^{*}_{0})-D_{\text{KL}}(\textrm{q}^{*}_{0}\parallel\textrm{q}^{*}_{0})\leq\lambda_{n}(\Xi(\textrm{q}^{*}_{0})-\Xi(\textrm{q}^{*}_{\lambda_{n}}))\leq\lambda_{n}\Xi(\textrm{q}^{*}_{0})=o(D).$
By the $T_{1}$-transportation inequality and Kantorovich duality, we have
$\sup_{\bm{x}^{(n)}\in\mathbb{X}^{n}}\left|\frac{1}{D}\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{\textrm{q}_{0}^{*}}[\psi(\theta_{i})]\right)\right|\leq\frac{1}{D}W_{1}\left(\textrm{q}^{*}_{\lambda_{n}},\textrm{q}_{0}^{*}\right)\lesssim\sqrt{D\cdot D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel\textrm{q}^{*}_{0})}\overset{P_{\theta_{0}}}{\to}0.$ (C.32)
Proof of Corollary 3.
Under the assumptions,
$\sqrt{\sum_{i=1}^{D}ac_{ii}/D}\lesssim D,\quad\sqrt{\sum_{i=1}^{D}b_{i}^{2}/D}\lesssim D,\quad\sqrt{\sum_{i,j}c_{ij}^{2}}\lesssim D.$
When we plug these terms in the upper bounds (3), Eq. 3.4 follows as the desired result. ∎
For the linear regression model, we denote
$\bm{w}:=\sigma^{-2}\bm{X}^{T}\bm{y},\quad\text{and}\quad d_{i}=[\bm{B}_{\text{diag}}]_{ii},\quad\forall i\in[D],$ (C.33)
where $[\bm{B}_{\text{diag}}]_{ii}$ denotes the $i^{th}$ diagonal entry of matrix $\bm{B}_{\text{diag}}$.
The next result shows $\Xi$-VI respects log-concavity of the exact posterior.
Lemma 9.
Let Assumption 6 hold. For $\lambda_{n}\in\bar{\mathbb{R}}_{+}$, the solution $\hat{\textrm{q}}_{\lambda_{n}}$ to Eq. 3.1 is $(\kappa_{1}+\kappa_{2})$-log-concave. Moreover, for each $i$, the optimal EOT potential $\phi^{*}_{\lambda_{n},i}$ is $\kappa_{2}/(\lambda_{n}+1)$-convex and marginal $m^{*}_{\lambda_{n},i}$ is $(\kappa_{1}+\kappa_{2})$-log-concave.
Proof of Lemma 9.
We first prove existence. By Lagragian duality, $\Xi$-VI (Eq. 3.1) is equivalent to $\min_{\Xi(\textrm{q})\leq t(\lambda_{n})}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})$. An optimizer of the latter problem exists because $\Xi(\cdot)$ has weakly closed sublevel set in $\mathbb{P}_{2}(\Theta)$ and because $D_{\text{KL}}(\cdot\parallel\textrm{q}^{*}_{0})$ has weakly compact sub-level sets.
Recall the $\Xi$-variational posterior be represented in term of optimal marginals $m^{*}_{\lambda_{n}}$ and optimal EOT potentials $\phi^{*}_{\lambda_{n}}$:
$\textrm{q}^{*}_{\lambda_{n}}(\theta)=\exp\left(\sum_{i=1}^{D}\phi^{*}_{\lambda_{n},i}(\theta_{i})+\frac{1}{\lambda_{n}+1}\ell(\bm{x}^{(n)};\theta)\right)m^{*}_{\lambda_{n}}(\theta).$ (C.34)
By Proposition 4, $m^{*}_{\lambda_{n}}$ and $\phi^{*}_{\lambda_{n}}$ satisfy the following fixed point equations:
$m^{*}_{\lambda_{n},i}(\theta_{i})$ $=Z_{i}^{-1}\exp(-(\lambda_{n}+1)\phi^{*}_{\lambda_{n},i}(\theta_{i}))\pi_{i}(\theta_{i}),\quad\text{and }\quad$ (C.35)
$\hat{\phi}_{\lambda_{n},i}(\theta_{i})$ $=-\log\int_{\Theta_{-i}}\exp\left(\frac{1}{\lambda_{n}+1}\ell\left(\bm{x}^{(n)};\theta\right)-\lambda_{n}\sum_{j\neq i}\phi^{*}_{\lambda_{n},j}(\theta_{j})\right)\prod_{j\neq i}\pi_{j}(\theta_{j})d\theta_{-i}$
[…]
Using equations Eq. C.35 to replace $m^{*}_{\lambda_{n}}$ in Eq. C.34, the variational posterior $\textrm{q}^{*}_{\lambda_{n}}$ satisfies
$\textrm{q}^{*}_{\lambda_{n}}(\theta)\propto\exp\left(\frac{1}{\lambda_{n}+1}\ell(\bm{x}^{(n)};\theta)-\lambda_{n}\sum_{i=1}^{D}\phi^{*}_{\lambda_{n},i}(\theta_{i})\right)\pi(\theta).$ (C.36)
We now establish the log-concavity of $\textrm{q}^{*}_{\lambda_{n}}$. Applying equations Eq. C.35 to Eq. C.36, we get
$\textrm{q}^{*}_{\lambda_{n}}(\theta)\propto\exp\left(\frac{1}{\lambda_{n}+1}\ell(\bm{x}^{(n)};\theta)+\lambda_{n}\sum_{i=1}^{D}\log\mathbb{E}_{\pi_{-i}}\exp\left(\frac{1}{\lambda_{n}+1}\ell\left(\bm{x}^{(n)};\theta\right)-\lambda_{n}\sum_{j\neq i}\phi^{*}_{\lambda_{n},j}(\theta_{j})\right)\right)\pi(\theta).$ (C.37)
For $\alpha\in[0,1]$ and $\theta_{i}^{0},\theta_{i}^{1}\in\Theta_{i}$, we have
$-\hat{\phi}_{\lambda_{n},i}(\alpha\theta_{i}^{0}+(1-\alpha)\theta_{i}^{1})=\log\mathbb{E}_{\pi_{-i}}\left[\exp\left(\frac{1}{\lambda_{n}+1}\ell\left(\bm{x}^{(n)};\alpha\theta_{i}^{0}+(1-\alpha)\theta_{i}^{1},\theta_{-i}\right)-\lambda_{n}\sum_{j\neq i}\phi^{*}_{\lambda_{n},j}(\theta_{j})\right)\right]+C.$
Under the log-concavity assumption, the log-likelihood is $\kappa_{2}$-concave. This means
$\ell\left(\bm{x}^{(n)};\alpha\theta_{i}^{0}+(1-\alpha)\theta_{i}^{1},\theta_{-i}\right)\geq\alpha\ell\left(\bm{x}^{(n)};\theta_{i}^{0},\theta_{-i}\right)+(1-\alpha)\ell\left(\bm{x}^{(n)};\theta_{i}^{1},\theta_{-i}\right)+\frac{\kappa_{2}\alpha(1-\alpha)}{2}(\theta_{i}^{0}-\theta_{i}^{1})^{2}.$
By the Prékopa–Leindler inequality (Theorem 19.16, [67]), we have
$\mathbb{E}_{\pi_{-i}}\left[\exp\left(\frac{\alpha\ell\left(\bm{x}^{(n)};\theta_{i}^{0},\theta_{-i}\right)+(1-\alpha)\ell\left(\bm{x}^{(n)};\theta_{i}^{1},\theta_{-i}\right)}{\lambda_{n}+1}-\lambda_{n}\sum_{j\neq i}\phi^{*}_{\lambda_{n},j}(\theta_{j})\right)\right]$
$\geq\mathbb{E}_{\pi_{-i}}\left[\exp\left(\frac{\ell\left(\bm{x}^{(n)};\theta_{i}^{0},\theta_{-i}\right)}{\lambda_{n}+1}-\lambda_{n}\sum_{j\neq i}\phi^{*}_{\lambda_{n},j}(\theta_{j})\right)\right]^{\alpha}\mathbb{E}_{\pi_{-i}}\left[\exp\left(\frac{\ell\left(\bm{x}^{(n)};\theta_{i}^{1},\theta_{-i}\right)}{\lambda_{n}+1}-\lambda_{n}\sum_{j\neq i}\phi^{*}_{\lambda_{n},j}(\theta_{j})\right)\right]^{1-\alpha}.$
Using this and the concavity of the logarithmic function, we have
$-\hat{\phi}_{\lambda_{n},i}(\alpha\theta_{i}^{0}+(1-\alpha)\theta_{i}^{1})$
[…]
$+\frac{\kappa_{2}\alpha(1-\alpha)}{2(\lambda_{n}+1)}(\theta_{i}^{0}-\theta_{i}^{1})^{2}$
$\geq-\alpha\phi_{\lambda_{n},i}(\theta_{i}^{0})-(1-\alpha)\phi_{\lambda_{n},i}(\theta_{i}^{1})+\frac{\kappa_{2}\alpha(1-\alpha)}{2(\lambda_{n}+1)}(\theta_{i}^{0}-\theta_{i}^{1})^{2}.$
Thus, the function $-\hat{\phi}_{\lambda_{n},i}(\cdot)$ is $\kappa_{2}/(\lambda_{n}+1)$-concave. By the fixed point representation (C.35), $m^{*}_{\lambda_{n}}$ is $(\kappa_{2}+\kappa_{1})$-log-concave. Using the representation (C.36), we conclude that the distribution $\textrm{q}^{*}_{\lambda_{n}}$ is $(\kappa_{2}+\kappa_{1})$-log-concave. ∎
We introduce some notations to streamline the next two proofs.
Definition 11 (Nonlinear quadratic tilt).
Let $\mu$ be a probability measure on $\mathbb{R}$. For any $(\phi,\gamma)\in L_{1}(\mathbb{R})\in(0,\infty)$, set
[…]
and define the probability distribution $\mu_{\phi,\gamma}$ on $\mathbb{R}$ by setting
$\mu_{\phi,\gamma}(\theta):=\exp\left(-\phi(\theta)-\frac{\gamma}{2}\theta^{2}-c_{\mu}(\phi,\gamma)\right)\mu(\theta),\quad\forall\theta\in\mathbb{R}.$ (C.39)
For any probability measure $\mu$, we have $c_{\mu}(\phi,\gamma)<\infty$ for any $(\phi,\gamma)\in L_{1}(\mathbb{R})\in(0,\infty)$. Given the base measure $\mu$, the tilted measure $\mu_{\phi,\gamma}(\theta)$ has an exponential family density that has $(\phi(\theta),\theta^{2})$ as the sufficient statistics. We call $\mu_{\phi,\gamma}$ a nonlinear quadratic tilt of $\mu$.
Using Theorem 1 and Proposition 4, we have
$\textrm{q}^{*}_{\lambda_{n}}(\theta)=\frac{1}{\mathcal{Z}_{D}(\lambda_{n})}\exp\left(-\frac{\theta^{T}\bm{B}_{\text{off}}\theta}{2(\lambda_{n}+1)}+\bm{w}^{T}\theta+\sum_{i=1}^{D}c_{\pi_{i}}\left(\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1}\right)\right)\prod_{i=1}^{D}\pi_{i,\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1}}(\theta_{i}),$ (C.40)
Here $c_{\pi_{i}}\left(\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1}\right)$ is defined in eq. C.38, and $\pi_{i,\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1}}$ is the nonlinear quadratic tilt of $\pi_{i}$ with parameters $(\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1})$. The constant $\mathcal{Z}_{D}(\lambda_{n})$ is defined as:
$\mathcal{Z}_{D}(\lambda_{n}):=\int_{\Theta}\exp\left(-\frac{\theta^{T}\bm{B}_{\text{off}}\theta}{2(\lambda_{n}+1)}+\bm{w}^{T}\theta\right)\prod_{i=1}^{D}\pi_{i,\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1}}(\theta_{i})d\theta.$
When $\lambda_{n}=0$, $\mathcal{Z}_{D}(0)$ is the normalizing constant of the exact posterior. When $\lambda_{n}>0$, we can view $\mathcal{Z}_{D}(\lambda_{n})$ as an approximation to $\mathcal{Z}_{D}(0)$.
The log-concavity of $\textrm{q}^{*}_{\lambda_{n}}$ induces an upper bound of $\Xi(\textrm{q}^{*}_{\lambda_{n}})$ using the covariance matrix, the design matrix, and the regularization parameter.
Lemma 10.
Let Assumption 6 hold. the solution $\hat{\textrm{q}}_{\lambda_{n}}$ to Eq. 3.1 satisfies
$\Xi(\textrm{q}^{*}_{\lambda_{n}})\leq\frac{\|\bm{B}_{\text{off}}+\bm{B}_{\text{diag}}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\text{Cov}_{\textrm{q}^{*}_{\lambda_{n}}}(\theta))}{(\kappa_{1}+\kappa_{2})(\lambda_{n}+1)^{2}}$
Proof of Lemma 10.
Shifting $\theta\sim\textrm{q}^{*}_{\lambda_{n}}$ by a constant is equivalent to shifting $\theta\sim m^{*}_{\lambda_{n}}$ by the same constant. Since the KL divergence is invariant to constant shift, $\Xi(\textrm{q}^{*}_{\lambda_{n}})$ is the same if we shift $\textrm{q}^{*}_{\lambda_{n}}$ by a constant. This allows us to assume WLOG that $\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\theta]=0$.
Define $\tilde{m}_{i}(\theta_{i})=\pi_{i,\lambda_{n}\phi^{*}_{\lambda_{n},i},d_{i}^{\prime}/(\lambda_{n}+1)}(\theta_{i})$, and $\tilde{m}(\theta)=\prod_{i=1}^{D}\tilde{m}_{i}(\theta_{i})$. By the variational representation of mutual information, we have:
$\Xi(\textrm{q}^{*}_{\lambda_{n}})=D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel m^{*}_{\lambda_{n}})\leq D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel\tilde{m}).$
By Lemma 9, $\phi^{*}_{\lambda_{n},i}$ is $\kappa_{2}/(\lambda_{n}+1)$-convex. Since $\pi$ is $\kappa_{1}$-log-concave by assumption, $\tilde{m}$ is $(\kappa_{1}+\kappa_{2})$-log-concave. By the log-Sobolev inequality, we have:
$\Xi(\textrm{q}^{*}_{\lambda_{n}})$ $=D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel m^{*}_{\lambda_{n}})$
[…]
$=\frac{1}{(\kappa_{1}+\kappa_{2})(\lambda_{n}+1)^{2}}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\|(\bm{B}_{\text{off}}+\bm{B}_{\text{diag}}-\bm{B}_{\text{diag}}^{\prime})\theta\|_{2}^{2}\right].$
Using the matrix identity $\|A\theta\|_{2}\leq\|A\|_{2}\|\theta\|_{2}$, we have:
$\Xi(\textrm{q}^{*}_{\lambda_{n}})$ $\leq\frac{\|\bm{B}_{\text{off}}+\bm{B}_{\text{diag}}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\|\theta\|_{2}^{2}]}{(\kappa_{1}+\kappa_{2})(\lambda_{n}+1)^{2}}=\frac{\|\bm{B}_{\text{off}}+\bm{B}_{\text{diag}}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\text{Cov}_{\textrm{q}^{*}_{\lambda_{n}}}(\theta))}{(\kappa_{1}+\kappa_{2})(\lambda_{n}+1)^{2}}.$
The last equality follows from the assumption that $\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\theta]=0$. ∎
Proof of Theorem 4.
Define $\tilde{\bm{B}}_{\text{off}}:=\frac{\bm{B}_{\text{off}}}{\lambda_{n}+1}$ and $\tilde{\pi}_{i}:=\pi_{i,\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1}}(\theta_{i})$. We can write
$\textrm{q}^{*}_{\lambda_{n}}(\theta)=\frac{1}{\mathcal{Z}_{D}(\lambda_{n})}\exp\left(-\frac{\theta^{T}\bm{B}_{\text{off}}\theta}{2(\lambda_{n}+1)}+\bm{w}^{T}\theta+\sum_{i=1}^{D}c_{\pi_{i}}(\lambda_{n}\phi^{*}_{\lambda_{n},i},\frac{d_{i}}{\lambda_{n}+1})\right)\prod_{i=1}^{D}\tilde{\pi}_{i}(\theta_{i}).$
By Lemma 9, $\textrm{q}^{*}_{\lambda_{n}}$ is a $(\kappa_{1}+\kappa_{2})$-log-concave. By Theorem 1 of [39], we have:
$\log\mathcal{Z}_{D}(\lambda_{n})-\sup_{m\in\mathcal{M}(\Theta)}\left[-\frac{1}{2}\mathbb{E}_{m}[\theta]^{T}\tilde{\bm{B}}_{\text{off}}\mathbb{E}_{m}[\theta]+\bm{w}^{T}\mathbb{E}_{m}[\theta]-\sum_{i=1}^{D}D_{\text{KL}}(m_{i}\parallel\tilde{\pi}_{i})\right]\leq\frac{\sum_{j=1}^{D}\sum_{i=1}^{D}[\bm{B}_{\text{off}}]_{ij}^{2}}{(\kappa_{1}+\kappa_{2})^{2}(\lambda_{n}+1)^{2}},$ (C.41)
where $\mathbb{E}_{m}[\theta]$ is the mean vector of $m$.
For any $m\in\mathcal{M}(\Theta)$, we have
[…]
$=\log\mathcal{Z}_{D}(\lambda_{n})+\frac{1}{2}\mathbb{E}_{m}[\theta]^{T}\tilde{\bm{B}}_{\text{off}}\mathbb{E}_{m}[\theta]-\bm{w}^{T}\mathbb{E}_{m}[\theta]+\sum_{i=1}^{D}D_{\text{KL}}(m_{i}\parallel\tilde{\pi}_{i}).$
We invokve the upper bound on the log normalizer Eq. C.41:
$\inf_{m\in\mathcal{M}(\Theta)}D_{\text{KL}}(m\parallel\textrm{q}^{*}_{\lambda_{n}})\leq\frac{\sum_{j=1}^{D}\sum_{i=1}^{D}[\bm{B}_{\text{off}}]_{ij}^{2}}{(\kappa_{1}+\kappa_{2})^{2}(\lambda_{n}+1)^{2}}.$ (C.42)
By the $T_{2}$-transportion inequality (Theorem 1 and 2, [52]), we upper bound the Wasserstein metric with the square root of KL divergence:
$\inf_{m\in\mathcal{M}(\Theta)}W_{2}(\textrm{q}^{*}_{\lambda_{n}},m)\leq\sqrt{\frac{2}{\kappa_{2}+\kappa_{1}}\inf_{m\in\mathcal{M}(\Theta)}D_{\text{KL}}(m\parallel\textrm{q}^{*}_{\lambda_{n}})}\leq\sqrt{\frac{2\sum_{j=1}^{D}\sum_{i=1}^{D}[\bm{B}_{\text{off}}]_{ij}^{2}}{(\kappa_{1}+\kappa_{2})^{3}(\lambda_{n}+1)^{2}}}.$ (C.43)
For $\lambda_{n}\succ\sqrt{\text{tr}(\bm{B}_{\text{off}}^{2})}$, we have $\inf_{m\in\mathcal{M}(\Theta)}W_{2}(\textrm{q}^{*}_{\lambda_{n}},m)\overset{P_{\theta_{0}}}{\to}0$.
Consider the second regime $\lambda_{n}\succ\sqrt{\text{tr}(\bm{B}_{\text{off}}^{2})/D}$. By the triangle inequality, we have
$\sup_{\bm{y}\in\mathbb{R}^{n}}\inf_{m\in\mathcal{M}(\Theta)}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})-\frac{1}{D}\sum_{i=1}^{D}\mathbb{E}_{m}[\psi(\theta_{i})]\right)^{2}\right]^{1/2}$ (C.44)
$\leq\sup_{\bm{y}\in\mathbb{R}^{n}}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})-\frac{1}{D}\sum_{i=1}^{D}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]\right)^{2}\right]^{1/2}+\inf_{m\in\mathcal{M}(\Theta)}\left|\frac{1}{D}\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{m}[\psi(\theta_{i})]\right)\right|.$
Since $\psi$ is $1$-Lipschitz, we apply Katorovich duality to bound the second term.
$\left(\inf_{m\in\mathcal{M}(\Theta)}\left|\frac{1}{D}\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{m}[\psi(\theta_{i})]\right)\right|\right)^{2}$
$\leq\inf_{m_{1},\cdots m_{D}}\frac{1}{D}\sum_{i=1}^{D}W_{1}^{2}(\textrm{q}^{*}_{\lambda_{n},i},m_{i})\leq\inf_{m_{1},\cdots m_{D}}\frac{1}{D}\sum_{i=1}^{D}W_{2}^{2}(\textrm{q}^{*}_{\lambda_{n},i},m_{i}).$
By the subadditivity inequality of Wasserstein distance and Eq. C.43, we have
$\inf_{m_{1},\cdots m_{D}}\sum_{i=1}^{D}W_{2}^{2}(\textrm{q}^{*}_{\lambda_{n},i},m_{i})\leq\inf_{m\in\mathcal{M}(\Theta)}W_{2}^{2}(\textrm{q}^{*}_{\lambda_{n}},m)\leq\frac{2\sum_{j=1}^{D}\sum_{i=1}^{D}[\bm{B}_{\text{off}}]_{ij}^{2}}{(\kappa_{1}+\kappa_{2})^{3}(\lambda_{n}+1)^{2}}$ (C.45)
Thus,
$\inf_{m\in\mathcal{M}(\Theta)}\left|\frac{1}{D}\sum_{i=1}^{D}\left(\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]-\mathbb{E}_{m}[\psi(\theta_{i})]\right)\right|\leq\sqrt{\frac{2\sum_{j=1}^{D}\sum_{i=1}^{D}[\bm{B}_{\text{off}}]_{ij}^{2}}{D(\kappa_{1}+\kappa_{2})^{3}(\lambda_{n}+1)^{2}}}.$ (C.46)
The Lipschitzness implies $\|\nabla\psi\|_{2}\leq 1$. To bound the first term, we apply Poincaré inequality to the function $x\mapsto\frac{1}{D}\sum_{i=1}^{D}\psi(x_{i})$.
$\sup_{\bm{y}\in\mathbb{R}^{n}}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})-\frac{1}{D}\sum_{i=1}^{D}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\psi(\theta_{i})]\right)^{2}\right]\leq\sup_{\bm{y}\in\mathbb{R}^{n}}\text{Var}_{\textrm{q}^{*}_{\lambda_{n}}}\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})\right)$ (C.47)
$\leq\frac{1}{(\kappa_{1}+\kappa_{2})D^{2}}\sum_{i=1}^{D}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}[\|\nabla\psi\|_{2}^{2}]\leq\frac{1}{D(\kappa_{1}+\kappa_{2})}.$
Combining bounds Eq. C.46 and Eq. C.47, we have
$\sup_{\bm{y}\in\mathbb{R}^{n}}\inf_{m\in\mathcal{M}(\Theta)}\mathbb{E}_{\textrm{q}^{*}_{\lambda_{n}}}\left[\left(\frac{1}{D}\sum_{i=1}^{D}\psi(\theta_{i})-\frac{1}{D}\sum_{i=1}^{D}\mathbb{E}_{m}[\psi(\theta_{i})]\right)^{2}\right]$ (C.48)
$\leq\frac{(\kappa_{1}+\kappa_{2})^{2}(\lambda_{n}+1)^{2}+2\sum_{j=1}^{D}\sum_{i=1}^{D}[\bm{B}_{\text{off}}]_{ij}^{2}}{D(\kappa_{1}+\kappa_{2})^{3}(\lambda_{n}+1)^{2}}$
For $\lambda_{n}\succ\sqrt{\text{tr}(\bm{B}_{\text{off}}^{2})/D}$, the bounds implies the Eq. 3.10.
Consider the third regime $\lambda_{n}\prec\left[\|\bm{B}_{\text{off}}+\bm{B}_{\text{diag}}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))\right]^{-1}$. Recall that $\Xi$-VI is equivalent to $\min_{\Xi(\textrm{q})\leq t(\lambda_{n})}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})$ for some constant $t(\lambda_{n})$ depending on $\lambda_{n}$. If $\Xi(\textrm{q}^{*}_{0})\leq t(\lambda_{n})$, then $\textrm{q}^{*}_{\lambda_{n}}=\textrm{q}^{*}_{0}$, hence $\Xi(\textrm{q}^{*}_{0})\geq\Xi(\textrm{q}^{*}_{\lambda_{n}})$ for all $\lambda_{n}$. For $t<\lambda_{n}$ and fixed $n$, we apply Lemma 10 to obtain an upper bound,
$D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel\textrm{q}^{*}_{0})-D_{\text{KL}}(\textrm{q}^{*}_{t}\parallel\textrm{q}^{*}_{0})\leq\lambda_{n}(\Xi(\textrm{q}^{*}_{t})-\Xi(\textrm{q}^{*}_{\lambda_{n}}))\leq\lambda_{n}\Xi(\textrm{q}^{*}_{0})\overset{P_{\theta_{0}}}{\to}0.$
Finally, consider the fourth regime $\lambda_{n}\prec D^{-1}\left[\|\bm{B}_{\text{off}}+\bm{B}_{\text{diag}}-\bm{B}_{\text{diag}}^{\prime}\|_{2}^{2}\text{tr}(\text{Cov}_{\textrm{q}^{*}_{0}}(\theta))\right]^{-1}$. We follow an analogous derivation as the third regime:
$D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel\textrm{q}^{*}_{0})-D_{\text{KL}}(\textrm{q}^{*}_{t}\parallel\textrm{q}^{*}_{0})\leq\lambda_{n}(\Xi(\textrm{q}^{*}_{t})-\Xi(\textrm{q}^{*}_{\lambda_{n}}))\leq\lambda_{n}\Xi(\textrm{q}^{*}_{0})=o(D).$
Since $\textrm{q}^{*}_{0}$ is $(\kappa_{1}+\kappa_{2})$-log-concave, we invoke the $T_{2}$-transportion inequality:
$W_{2}(\textrm{q}^{*}_{\lambda_{n}},\textrm{q}^{*}_{0})\leq\sqrt{\frac{2}{\kappa_{2}+\kappa_{1}}D_{\text{KL}}(\textrm{q}^{*}_{\lambda_{n}}\parallel\textrm{q}^{*}_{0})}.$
The remaining proofs for the third and fourth regimes are the same as the first two regimes, where we plug in the upper bounds for the KL divergence to upper-bound the Wasserstein distance. We skip repeating the details. ∎
Proof of Corollary 4.
Given $\text{tr}(\bm{B}_{\text{off}}^{2})=\sum_{i=1}^{D}\eta_{i}^{2}$, Theorem 4 ensures that the convergence of $W_{2}(\textrm{q}^{*}_{\lambda_{n}},m_{\lambda_{n}}^{*})$ holds for $\lambda_{n}\succ\sqrt{\sum_{i=1}^{D}\eta_{i}^{2}}$. Since $\sum_{i=1}^{D}\eta_{i}^{2}\lesssim D^{2}$, we have $W_{2}(\textrm{q}^{*}_{\lambda_{n}},m_{\lambda_{n}}^{*})$ converges in probability to zero, for any choice of $\lambda_{n}\succ D$. ∎
Proofs of Section 3.3
We first state an auxiliary lemma to Theorem 5.
Lemma 11.
Let $\textrm{q}^{*}_{\lambda}$ be the $\Xi$-variational posterior. Then $\text{ELBO}(\textrm{q}^{*}_{\lambda})$ and $\mathcal{C}_{\lambda}$ are monotonically decreasing function of $\lambda$.
Proof of Lemma 11.
Since the variational posterior $\textrm{q}^{*}_{\lambda}$ is a maximizer of $\text{ELBO}(\textrm{q})-\lambda\Xi(\textrm{q})$, we have
$\mathcal{C}_{\lambda}=\text{ELBO}(\textrm{q}^{*}_{\lambda})-\lambda\Xi(\textrm{q}^{*}_{\lambda}).$
For $\lambda_{1}<\lambda_{2}$, we have
$\mathcal{C}_{\lambda_{1}}=\text{ELBO}(\textrm{q}^{*}_{\lambda_{1}})-\lambda_{1}\Xi(\textrm{q}^{*}_{\lambda_{1}})\geq\text{ELBO}(\textrm{q}^{*}_{\lambda_{2}})-\lambda_{1}\Xi(\textrm{q}^{*}_{\lambda_{2}})\geq\text{ELBO}(\textrm{q}^{*}_{\lambda_{2}})-\lambda_{2}\Xi(\textrm{q}^{*}_{\lambda_{2}})=\mathcal{C}_{\lambda_{2}}.$
By Langragian duality, we have $\text{ELBO}(\textrm{q}^{*}_{\lambda})=\max\limits_{\Xi(\textrm{q})\leq t(\lambda)}\text{ELBO}(\textrm{q})$ for $t(\lambda)$ monotonically decreasing in $\lambda$.
For $\lambda_{1}<\lambda_{2}$, $t(\lambda_{1})\geq t(\lambda_{2})$ hence
$\text{ELBO}(\textrm{q}^{*}_{\lambda_{1}})=\max\limits_{\Xi(\textrm{q})\leq t(\lambda_{1})}\text{ELBO}(\textrm{q})\geq\max\limits_{\Xi(\textrm{q})\leq t(\lambda_{2})}\text{ELBO}(\textrm{q})=\text{ELBO}(\textrm{q}^{*}_{\lambda_{2}}).$
Proof of Theorem 5.
Let $(P_{2}(\Theta),W_{2})$ be the metric space. We want to show that the functionals
$F_{\lambda}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})+\lambda\Xi(\textrm{q}).$
$\Gamma$-converge to
$F_{\infty}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})+\infty\Xi(\textrm{q}),$
as $\lambda\to\infty$.
To verify $\Gamma$ convergence, we make use of the property that the KL divergence functional $D_{\text{KL}}(\cdot\parallel\textrm{q}^{*}_{0})$ and $\Xi(.)$ functional are lower semicontinuous (l.s.c.) in Wasserstein metric. This is provided in Lemma 5.
Let $\textrm{q}\in\mathbb{P}_{2}(\Theta)$ and $W_{2}(\textrm{q}_{\lambda},\textrm{q})\to 0$. If q is a product measure, then
$F_{\infty}(\textrm{q})$ $=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})\leq\liminf_{\lambda\to\infty}D_{\text{KL}}(\textrm{q}_{\lambda}\parallel\textrm{q}^{*}_{0})\leq\liminf_{\lambda\to\infty}F_{\lambda}(\textrm{q}_{\lambda}).$
The first inequality holds because $D_{\text{KL}}(.\parallel\textrm{q}^{*}_{0})$ is l.s.c.
If q is not a product measure, we have $\liminf_{n\to\infty}\Xi(\textrm{q}_{n})\geq\Xi(\textrm{q})>0$ by the lower semicontinuity of $\Xi$. Since the KL term is nonnegative, we have
$F_{\infty}(\textrm{q})=\infty=\liminf_{\lambda\to\infty}F_{\lambda}(\textrm{q}_{\lambda}).$
Thus the liminf inequality is verified.
Next we show the existence of a recovery sequence. For any $\textrm{q}\in\mathbb{P}_{2}(\Theta)$, we take $_{\lambda}=\textrm{q}$ q. If q is a product measure, then
$F_{\infty}(\textrm{q})=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})\geq D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0}).$
Otherwise,
$F_{\infty}(\textrm{q})=\infty\geq\limsup_{\lambda\to\infty}F_{\lambda}(\textrm{q}_{\lambda}).$
This verifies the limsup inequality. Combining the liminf and limsup inequalities, we obtain that $F_{\infty}=\Gamma-\text{lim}_{\lambda\to\infty}F_{\lambda}$.
Next we prove that the sequence $F_{\lambda}$ is eqi-coercive. Take $\lambda_{j}\to\infty$ and $\textrm{q}_{\lambda_{j}}$ such that $F_{\lambda_{j}}(\textrm{q}_{\lambda_{j}})\leq t$ for all $j$. Then $\Xi(\textrm{q}_{\lambda_{j}})=o(1)$ because $\lambda_{j}\Xi(\textrm{q}_{\lambda_{j}})$ is bounded as $\lambda_{j}\to\infty$. Moreover, $D_{\text{KL}}(\textrm{q}_{\lambda_{j}}\parallel\textrm{q}^{*}_{0})$ is upper bounded by $t$. Since $D_{\text{KL}}(.\parallel\textrm{q}^{*}_{0})$ is Wasserstein (geodastically) convex, it is coercive by Lemma 2.4.8 of [6]. Thus, there exists a converging sequence $\textrm{q}_{\lambda_{j}}^{\prime}$ such that $D_{\text{KL}}(\textrm{q}_{\lambda_{j}}^{\prime}\parallel\textrm{q}^{*}_{0})\leq D_{\text{KL}}(\textrm{q}_{\lambda_{j}}\parallel\textrm{q}^{*}_{0})+o(1)$. Since $\Xi(\textrm{q}_{\lambda_{j}})=o(1)$, we obtain that $F_{\lambda_{j}}(\textrm{q}_{j}^{\prime})\leq F_{\lambda_{j}}(\textrm{q}_{j})+o(1)$. This verifies the equi-coercivity of $F_{\lambda}$.
Finally, by the fundamental theorem of $\Gamma$ convergence (Theorem 10), we conclude that
$W_{2}(\textrm{q}^{*}_{\infty},\textrm{q}^{*}_{\lambda})\to 0,\quad\text{as}\quad\lambda\to\infty,$
and
$\left|\mathcal{C}_{\lambda}-\mathcal{C}_{\infty}\right|\to 0,\quad\text{as}\quad\lambda\to\infty.$
By Corollary 2.1 of [14], every minimizer of $F_{\infty}$ is the limit of some converging minimizing sequences of $F_{\lambda}$. For any $\textrm{q}^{*}_{\infty}\in\mathbb{Q}_{\infty}$, this implies the existence of a sequence $\textrm{q}^{*}_{\lambda}\in\mathbb{Q}_{\lambda}$ such that
$W_{2}(\textrm{q}^{*}_{\infty},\textrm{q}^{*}_{\lambda})\to 0,\quad\text{as}\quad\lambda\to\infty.$
Proof of Theorem 6.
We define $\mathbb{P}_{2}^{\prime}(\Theta)$ as $P_{2}^{\prime}(\Theta)=\{\textrm{q}\in\mathbb{P}_{2}(\Theta):\Xi(\textrm{q})<\infty\}$. The space $(\mathbb{P}_{2}^{\prime}(\Theta),W_{2})$ is a metric space. We want to show that the sequence of functionals
$F_{\lambda}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})+\lambda\Xi(\textrm{q}).$
$\Gamma$-converge to
$F_{0}(\textrm{q}):=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0}),$
as $\lambda\to 0$. Both $F_{\lambda}(\textrm{q})$ and $F_{0}(\textrm{q})$ are defined on $(\mathbb{P}_{2}^{\prime}(\Theta),W_{2})$.
We make use of Lemma 5 which shows that the KL divergence functional $D_{\text{KL}}(\cdot\parallel\textrm{q}^{*}_{0})$ and $\Xi(.)$ functional are lower semicontinuous (l.s.c.) in Wasserstein metric.
Let $\textrm{q}\in\mathbb{P}_{2}^{\prime}(\Theta)$ and $W_{2}(\textrm{q}_{\lambda},\textrm{q})\to 0$. We have
$F_{0}(\textrm{q})$ $=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})\leq\liminf_{\lambda\to 0}D_{\text{KL}}(\textrm{q}_{\lambda}\parallel\textrm{q}^{*}_{0})\leq\liminf_{\lambda\to 0}F_{\lambda}(\textrm{q}_{\lambda}).$
The first inequality holds because $D_{\text{KL}}(.\parallel\textrm{q}^{*}_{0})$ is l.s.c. The second inequality holds because $\Xi(\cdot)$ is nonnegative.
Next we show that the existence of a recovery sequence. For any $\textrm{q}\in\mathbb{P}_{2}(\Theta)$, we take $_{\lambda}=\textrm{q}$ q. Since $\Xi(\textrm{q})<\infty$, we have
$F_{0}(\textrm{q})=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})\geq\limsup_{\lambda\to 0}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})+\lambda\Xi(\textrm{q}).$
This verifies the limsup inequality. Combining the liminf and limsup inequalities, we obtain that $F=\Gamma-\text{lim}_{\lambda\to\infty}F_{\lambda}$.
We proceed to establish equi-coercivity of the sequence $F_{\lambda}$. Consider a sequence $\lambda_{j}\to 0$ and $\textrm{q}_{\lambda_{j}}\in\mathbb{P}_{2}^{\prime}(\Theta)$ for which $F_{\lambda_{j}}(\textrm{q}_{\lambda_{j}})\leq t$ holds for all $j$. Given that $\Xi(\textrm{q}_{\lambda_{j}})\geq 0$, it follows that $D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})\leq t$. Owing to the geodesic convexity of the Kullback-Leibler divergence $D_{\text{KL}}(\cdot\parallel\textrm{q}^{*}_{0})$ in the Wasserstein space, Lemma 2.4.8 from [6] ensures that it is coercive, implying that the set $\{\textrm{q}\in\mathbb{P}_{2}^{\prime}(\Theta)\mid D_{\text{KL}}(\textrm{q}\parallel\textrm{q}^{*}_{0})\leq t\}$ is compact in the metric space $(\mathbb{P}_{2}^{\prime}(\Theta),W_{2})$. Sequential compactness guarantees the existence of a convergent subsequence of $\textrm{q}_{\lambda_{j}}$, which converges to some $\textrm{q}_{0}$ in $\mathbb{P}_{2}^{\prime}(\Theta)$. Since $D_{\text{KL}}(\cdot\parallel\textrm{q}^{*}_{0})$ is lower semicontinuous (l.s.c.), we conclude that:
$F_{\lambda_{j}}(\textrm{q}_{0})$ $=D_{\text{KL}}(\textrm{q}_{0}\parallel\textrm{q}^{*}_{0})+\lambda_{j}\Xi(\textrm{q}_{0})\leq D_{\text{KL}}(\textrm{q}_{\lambda_{j}}\parallel\textrm{q}^{*}_{0})+\lambda_{j}\Xi(\textrm{q}_{0})$
$\leq F_{\lambda_{j}}(\textrm{q}_{\lambda_{j}})+\lambda_{j}\Xi(\textrm{q}_{0})=F_{\lambda_{j}}(\textrm{q}_{\lambda_{j}})+o(1),$
Finally, by the fundamental theorem of $\Gamma$ convergence, we conclude that
$W_{2}(\textrm{q}^{*}_{0},\textrm{q}^{*}_{\lambda})\to 0,\quad\text{as}\quad\lambda\to 0,$
where $\textrm{q}_{0}$ is a minimizer of $F_{0}$, and
$\left|\mathcal{C}_{\lambda}-\mathcal{C}_{0}\right|\to 0,\quad\text{as}\quad\lambda\to 0.$
Since $\textrm{q}^{*}_{0}$ is the unique minimizer of $F_{0}$, we conclude that $\textrm{q}_{0}=\textrm{q}^{*}_{0}$.
To prove the convergence of optimal cost, we note that
$F_{\lambda}(\textrm{q}^{*}_{\lambda})\leq F_{\lambda}(\textrm{q}^{*}_{0})=F_{0}(\textrm{q}^{*}_{0})+\lambda\Xi(\textrm{q}^{*}_{0}).$
Thus,
$|\mathcal{C}_{\lambda}-\mathcal{C}_{0}|=|F_{\lambda}(\textrm{q}^{*}_{\lambda})-F_{0}(\textrm{q}^{*}_{0})|\leq\lambda\Xi(\textrm{q}^{*}_{0}).$
Define a functional $\Phi_{\lambda}$ that combines the objective functional of the inner variational objective problem and Assumption 7:
$\Phi_{\lambda}(\textrm{q}):=\mathbb{E}_{\textrm{q}}\left[-\ell(\bm{x}^{(n)};\theta)+\sum_{i=1}^{D}\phi_{i}(\theta_{i})\right]+(\lambda+1)\Xi(\textrm{q}),$ (C.49)
where $\phi_{i}:\Theta_{i}\mapsto\mathbb{R}$ are the one-dimensional function in the Lipschitz cost assumption (Assumption 7). Since $\phi_{i}$ are tensorized, minimizing $\Phi_{\lambda}$ over $\mathcal{C}(m)$ is equivalent to solving the inner variational problem over $\mathcal{C}(m)$.
For proving Theorem 7, we introduce a Pythagorean theorem for the inner variational problem.
Lemma 12.
Let $\textrm{q}_{\lambda}\in\mathcal{C}(m)$ be a optimizer of $\Phi_{\lambda}$ over $\mathcal{C}(m)$. Then
$D_{\text{KL}}(\textrm{q},\textrm{q}_{\lambda})\leq\Phi_{\lambda}(\textrm{q})-\Phi_{\lambda}(\textrm{q}_{\lambda}),\quad\text{for all }\textrm{q}\in\mathcal{C}(m).$
Proof of Lemma 12.
We recall definition of the auxiliary measure $\textrm{q}_{\text{aux}}$ in the proof of Theorem 1, $\textrm{q}_{\text{aux}}(\theta)=\alpha^{-1}e^{\frac{\ell(\bm{x}^{(n)};\theta)-\sum_{i=1}^{D}\phi_{i}(\theta_{i})}{\lambda+1}}m(\theta)$, where $\alpha$ is the normalizing constant. Then
$\Phi_{\lambda}(\textrm{q})=D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{\text{aux}})-\log\alpha,$ (C.50)
so that the entropic optimal transport problem is equivalent to minimizing $D_{\text{KL}}(\cdot\parallel\textrm{q}_{\text{aux}})$. In particular, $\textrm{q}_{\lambda}=\arg\min_{\mathcal{C}(m)}D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{\text{aux}})$ and the Pythagorean theorem for relative entropy (Theorem 2.2, [21]) yields
$D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{\text{aux}})\geq D_{\text{KL}}(\textrm{q}_{\lambda}\parallel\textrm{q}_{\text{aux}})+D_{\text{KL}}(\textrm{q}\parallel\textrm{q}_{\lambda})\quad\text{for all }\textrm{q}\in\mathcal{C}(m).$
In view of Eq. C.50, the desired claim holds. ∎
The next Lemma is also auxiliary to the proof of Theorem 7.
Lemma 13.
Let $\textrm{q}^{*}_{\lambda}\in\mathcal{C}(m^{*})$ be a optimizer of $\Phi_{\lambda}$ over $\mathcal{C}(m^{*})$, and $\textrm{q}_{\lambda}^{s}\in\mathcal{C}(\tilde{m})$ be its shadow. Then
$\left|\Phi_{\lambda}(\textrm{q}^{*}_{\lambda})-\Phi_{\lambda}(\textrm{q}_{\lambda}^{s})\right|\leq LW_{2}(\textrm{q}^{*}_{\lambda},\textrm{q}_{\lambda}^{s}).$
Proof of Lemma 13.
Using the Lipschitz cost assumption and Lemma 6, we have
$\Phi_{\lambda}(\textrm{q}^{*}_{\lambda})$ $=\mathbb{E}_{\textrm{q}^{*}_{\lambda}}\left[-\ell(\bm{x}^{(n)};\theta)+\sum_{i=1}^{D}\phi_{i}(\theta_{i})\right]+(\lambda+1)\Xi(\textrm{q}^{*}_{\lambda}).$
$\geq\mathbb{E}_{\textrm{q}_{\lambda}^{s}}\left[-\ell(\bm{x}^{(n)};\theta)+\sum_{i=1}^{D}\phi_{i}(\theta_{i})\right]-LW_{2}(\textrm{q}^{*}_{\lambda},\textrm{q}_{\lambda}^{s})+(\lambda+1)\Xi(\textrm{q}_{\lambda}^{s})$
$=\Phi_{\lambda}(\textrm{q}_{\lambda}^{s})-LW_{2}(\textrm{q}^{*}_{\lambda},\textrm{q}_{\lambda}^{s}).$
The claim follows by a symmetric argument. ∎
Proof of Theorem 7.
Consider the optimizers $\tilde{\textrm{q}}_{\lambda}\in\mathcal{C}(\tilde{m})$ and $\textrm{q}^{*}_{\lambda}\in\mathcal{C}(m^{*}_{\lambda})$. Let $\textrm{q}^{s}_{\lambda}\in\mathcal{C}(\tilde{m})$ be the shadow of $\textrm{q}^{*}_{\lambda}$. By Lemma 6 and the Lipschitz cost assumption, we have:
[…]
$\leq LW_{2}(\textrm{q}_{\lambda}^{s},\textrm{q}^{*}_{\lambda})\leq LW_{2}(m^{*}_{\lambda},\tilde{m}).$
Lemma 13 implies $\Phi_{\lambda}(\tilde{\textrm{q}}_{\lambda})-\Phi_{\lambda}(\textrm{q}^{*}_{\lambda})\leq LW_{2}(m^{*}_{\lambda},\tilde{m})$. Adding the inequalities shows:
$|\Phi_{\lambda}(\tilde{\textrm{q}}_{\lambda})-\Phi_{\lambda}(\textrm{q}_{\lambda}^{s})|\leq 2LW_{2}(m^{*}_{\lambda},\tilde{m}).$
By Lemma 12, we have that $D_{\text{KL}}(\tilde{\pi},\pi^{*})\leq 2LW_{2}(m^{*}_{\lambda},\tilde{m})$, and the transport inequality assumption implies:
$W_{\rho}(\textrm{q}_{\lambda}^{s},\tilde{\textrm{q}}_{\lambda})\leq C_{\rho}(2LW_{2}(m^{*}_{\lambda},\tilde{m}))^{\frac{1}{2\rho}}.$
By Lemma 6, we get $W_{2}(\textrm{q}^{*}_{\lambda},\textrm{q}_{\lambda}^{s})=W_{2}(m^{*}_{\lambda},\tilde{m})$. We conclude the proof via the triangle inequality,
$W_{2}(\textrm{q}^{*}_{\lambda},\tilde{\textrm{q}}_{\lambda})\leq W_{2}(\textrm{q}^{*}_{\lambda},\textrm{q}_{\lambda}^{s})+W_{2}(\textrm{q}_{\lambda}^{s},\tilde{\textrm{q}}_{\lambda})\leq W_{2}(m^{*}_{\lambda},\tilde{m})+C_{\textrm{q}}(2LW_{2}(m^{*}_{\lambda},\tilde{m}))^{\frac{1}{2\textrm{q}}}.$
Proofs of Section 4.1
Proof of Proposition 2.
By Eq. 4.2, we have
$\Lambda^{*}_{\text{diag}}=\frac{1}{\lambda+1}\Lambda_{0,diag}+\frac{\lambda}{\lambda+1}(\Sigma^{*}_{\text{diag}})^{-1}$ $\implies(\lambda+1)\Lambda^{*}_{\text{diag}}-\lambda(\Sigma^{*}_{\text{diag}})^{-1}=\Lambda_{0,\text{diag}}$
By Lemma 3, we have the inequality
$\Sigma^{*^{-1}}_{\text{diag}}\leq\Lambda^{*}_{\text{diag}}\leq\Lambda_{0,\text{diag}}.$ (C.51)
By Hua’s identity, we have
$\Sigma^{*}$ $=\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{\lambda}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}\right]^{-1}=\frac{\lambda+1}{\lambda}\Sigma^{*}_{\text{diag}}-\left[\frac{\lambda}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}+\frac{\lambda^{2}}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}\Sigma_{0}\Sigma^{*^{-1}}_{\text{diag}}\right]^{-1}.$
Taking the diagonal elements on both sides, we have
$\frac{1}{\lambda}\Sigma^{*}_{\text{diag}}$ $=\left(\left[\frac{\lambda}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}+\frac{\lambda^{2}}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}\Sigma_{0}\Sigma^{*^{-1}}_{\text{diag}}\right]^{-1}\right)_{\text{diag}}$
$=\left(\left[\underbrace{\frac{\lambda}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}+\frac{\lambda^{2}}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}\Sigma_{0,\text{diag}}\Sigma^{*^{-1}}_{\text{diag}}}_{A}+\underbrace{\frac{\lambda^{2}}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}\Sigma_{0,\text{off}}\Sigma^{*^{-1}}_{\text{diag}}}_{B}\right]^{-1}\right)_{\text{diag}}.$
Note that $B$ is a matrix with zero diagonal entries. By Lemma 3, we have
$\frac{1}{\lambda}\Sigma^{*}_{\text{diag}}=\left([A+B]^{-1}\right)_{\text{diag}}\geq[A+B]_{\text{diag}}^{-1}=A^{-1}.$
This implies that
$\frac{1}{\lambda}\Sigma^{*}_{\text{diag}}\geq\Sigma^{*}_{\text{diag}}\left[\frac{\lambda}{\lambda+1}\Sigma^{*}_{\text{diag}}+\frac{\lambda^{2}}{\lambda+1}\Sigma_{0,\text{diag}}\right]^{-1}\Sigma^{*}_{\text{diag}},$
which after simplification yields
$\Sigma^{*}_{\text{diag}}\leq\Sigma_{0,\text{diag}}.$ (C.52)
By Hua’s identity, we have
$\Sigma^{*}$ $=\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{\lambda}{\lambda+1}\Sigma^{*^{-1}}_{\text{diag}}\right]^{-1}=(\lambda+1)\Sigma_{0}-\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{1}{\lambda(\lambda+1)}\Lambda_{0}\Sigma^{*}_{\text{diag}}\Lambda_{0}\right]^{-1}.$
It follows that
$\Sigma^{*}-\Sigma_{0}$ $=\lambda\Sigma_{0}-\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{1}{\lambda(\lambda+1)}\Lambda_{0}\Sigma^{*}_{\text{diag}}\Lambda_{0}\right]^{-1}.$
By Eq. C.52, the matrix $\Sigma^{*}-\Sigma_{0}$ is negative semidefinite. By Eq. C.51, we have $\Sigma^{*}_{\text{diag}}\geq\Lambda_{0,\text{diag}}^{-1}$. Then
$\|\Sigma^{*}-\Sigma_{0}\|$ $\leq\left\|\lambda\Sigma_{0}-\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{1}{\lambda(\lambda+1)}\Lambda_{0}\Lambda_{0,\text{diag}}^{-1}\Lambda_{0}\right]^{-1}\right\|.$
Since $\Sigma^{*}_{\text{diag}}\leq\Sigma_{0,\text{diag}}$, we obtain a lower bound with analogous techniques.
$\|\Sigma^{*}-\Sigma_{0}\|$ $\geq\left\|\lambda\Sigma_{0}-\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{1}{\lambda(\lambda+1)}\Lambda_{0}\Sigma_{0,\text{diag}}\Lambda_{0}\right]^{-1}\right\|.$
This lower bound holds when the matrix on the right hand side is negative semidefinite. To see that, we have
$\lambda\Sigma_{0}-\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{1}{\lambda(\lambda+1)}\Lambda_{0}\Sigma_{0,\text{diag}}\Lambda_{0}\right]^{-1}$ $=-\Sigma_{0}+\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{\lambda}{\lambda+1}\Sigma_{0,\text{diag}}^{-1}\right]^{-1},$
Since $\Lambda_{0,\text{diag}}\geq\Sigma_{0,\text{diag}}^{-1}$, we have
$\left(-\Sigma_{0}+\left[\frac{1}{\lambda+1}\Lambda_{0}+\frac{\lambda}{\lambda+1}\Sigma_{0,\text{diag}}^{-1}\right]^{-1}\right)_{\text{diag}}\leq-\Sigma_{0,\text{diag}}+\Sigma_{0,\text{diag}}=0.$
This completes the proof. ∎
Proof of Proposition 3.
Denote $\Lambda^{*}=\begin{pmatrix}a&b\\ b&c\end{pmatrix}$. The inverse is
$\Sigma^{*}=\frac{1}{ac-b^{2}}\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}.$
As shown in Section 4.1, we have
$\begin{pmatrix}a&b\\ b&c\end{pmatrix}=\frac{1}{\lambda+1}\begin{pmatrix}a_{0}&b_{0}\\ b_{0}&c_{0}\end{pmatrix}+\frac{\lambda}{\lambda+1}\begin{pmatrix}a-\frac{b^{2}}{c}&0\\ 0&d-\frac{b^{2}}{c}\end{pmatrix}.$
This implies
$\begin{pmatrix}a+\lambda\frac{b^{2}}{c}&(\lambda+1)b\\ (\lambda+1)b&c+\lambda\frac{b^{2}}{a}\end{pmatrix}=\begin{pmatrix}a_{0}&b_{0}\\ b_{0}&c_{0}\end{pmatrix}.$
This translates to a system of equations
$(\lambda+1)b$ $=b_{0}$
$a+\lambda\frac{b^{2}}{c}$ $=a_{0}$
$c+\lambda\frac{b^{2}}{a}$ $=c_{0}.$
The first Equation yields $b=\frac{b_{0}}{\lambda+1}$. The other two equations yield
$\frac{a_{0}c}{c_{0}a}=\frac{ac+b^{2}}{ac+b^{2}}=1\implies c=\frac{c_{0}}{a_{0}}a.$
Substituting $c$ gives us
$a+\lambda\frac{a_{0}b^{2}}{c_{0}a}=a_{0}.$
which yields $a=\frac{a_{0}}{2}\pm\sqrt{\frac{a_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{a_{0}b_{0}^{2}}{c_{0}}}$.
Similarly, substituting $a$ with $c$ yields $c=\frac{c_{0}}{2}\pm\sqrt{\frac{c_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{c_{0}b_{0}^{2}}{a_{0}}}$.
Finally, use the fact that $\Lambda^{*}=\Lambda_{0}$ when $\lambda=0$ to obtain the solution set
$a=\frac{a_{0}}{2}+\sqrt{\frac{a_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{a_{0}b_{0}^{2}}{c_{0}}},\quad b=\frac{b_{0}}{\lambda+1},\quad c=\frac{c_{0}}{2}+\sqrt{\frac{c_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{c_{0}b_{0}^{2}}{a_{0}}}.$
To obtain the covariance matrix, note that
$|\Lambda^{*}|=ac-b^{2}=\frac{a_{0}c_{0}}{2}\left(1+\sqrt{1-\frac{\lambda}{(\lambda+1)^{2}}\frac{4b_{0}^{2}}{a_{0}c_{0}}}\right)-\frac{1}{\lambda+1}b_{0}^{2}.$
By the matrix inversion formula,
$\Sigma^{*}=\frac{1}{|\Lambda^{*}|}\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}$ $=\begin{pmatrix}\frac{\frac{c_{0}}{2}+\sqrt{\frac{c_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{c_{0}b_{0}^{2}}{a_{0}}}}{\frac{a_{0}c_{0}}{2}\left(1+\sqrt{1-\frac{\lambda}{(\lambda+1)^{2}}\frac{4b_{0}^{2}}{a_{0}c_{0}}}\right)-\frac{1}{\lambda+1}b_{0}^{2}}&-\frac{b_{0}}{(\lambda+1)\left(\frac{a_{0}c_{0}}{2}\left(1+\sqrt{1-\frac{\lambda}{(\lambda+1)^{2}}\frac{4b_{0}^{2}}{a_{0}c_{0}}}\right)-\frac{1}{\lambda+1}b_{0}^{2}\right)}\\ -\frac{b_{0}}{(\lambda+1)\left(\frac{a_{0}c_{0}}{2}\left(1+\sqrt{1-\frac{\lambda}{(\lambda+1)^{2}}\frac{4b_{0}^{2}}{a_{0}c_{0}}}\right)-\frac{1}{\lambda+1}b_{0}^{2}\right)}&\frac{\frac{a_{0}}{2}+\sqrt{\frac{a_{0}^{2}}{4}-\frac{\lambda}{(\lambda+1)^{2}}\frac{a_{0}b_{0}^{2}}{c_{0}}}}{\frac{a_{0}c_{0}}{2}\left(1+\sqrt{1-\frac{\lambda}{(\lambda+1)^{2}}\frac{4b_{0}^{2}}{a_{0}c_{0}}}\right)-\frac{1}{\lambda+1}b_{0}^{2}}\\ \end{pmatrix}$
$=\begin{pmatrix}\frac{c_{0}}{a_{0}c_{0}-\frac{2}{\lambda+1+\sqrt{(\lambda+1)^{2}-\lambda\frac{4b_{0}^{2}}{a_{0}c_{0}}}}b_{0}^{2}}&-\frac{b_{0}}{\frac{a_{0}c_{0}}{2}\left(\lambda+1+\sqrt{(\lambda+1)^{2}-\lambda\frac{4b_{0}^{2}}{a_{0}c_{0}}}\right)-b_{0}^{2}}\\ -\frac{b_{0}}{\frac{a_{0}c_{0}}{2}\left(\lambda+1+\sqrt{(\lambda+1)^{2}-\lambda\frac{4b_{0}^{2}}{a_{0}c_{0}}}\right)-b_{0}^{2}}&\frac{a_{0}}{a_{0}c_{0}-\frac{2}{\lambda+1+\sqrt{(\lambda+1)^{2}-\lambda\frac{4b_{0}^{2}}{a_{0}c_{0}}}}b_{0}^{2}}\end{pmatrix}$
$=\frac{1}{a_{0}c_{0}-\psi^{-1}(\lambda)b_{0}^{2}}\begin{pmatrix}c_{0}&-\frac{b_{0}}{\psi(\lambda)}\\ -\frac{b_{0}}{\psi(\lambda)}&a_{0}\end{pmatrix}.$
where $\psi(\lambda)=\frac{1}{2}\left(\lambda+1+\sqrt{\left(\lambda-\frac{2b_{0}^{2}-a_{0}c_{0}}{a_{0}c_{0}}\right)^{2}+\frac{4b_{0}^{2}}{a_{0}^{2}c_{0}^{2}}\left(a_{0}c_{0}-b_{0}^{2}\right)}\right)$. ∎
Proof of Corollary 5.
We use the well known property of the $2$-Wasserstein metric that for any $\textrm{q}_{0},\textrm{q}_{1}\in\mathbb{P}_{2}(\Theta)$,
$W_{2}^{2}(\prod_{i=1}^{D}\textrm{q}_{0,i},\prod_{i=1}^{D}\textrm{q}_{1,i})\leq\sum_{i=1}^{D}W_{2}^{2}(\textrm{q}_{0,i},\textrm{q}_{1,i})\leq W_{2}^{2}(\textrm{q}_{0},\textrm{q}_{1}).$ (C.53)
By Theorem 5, $W_{2}(\textrm{q}^{*}_{\lambda},\textrm{q}^{*}_{\infty})\to 0$ as $\lambda\to\infty$, hence $W_{2}(m^{*}_{\lambda},\textrm{q}^{*}_{\infty})\to 0$ by the identity Eq. C.53 which implies $W_{2}(\textrm{q}^{*^{(\infty)}}_{\lambda},\textrm{q}^{*}_{\lambda})\to 0$ by Theorem 7. An analogous derivation holds for $\lambda\to 0$. ∎
Additional Simulation Results
Figure 8: Approximation errors for posterior covariance for Linear Regression for varying $\lambda$. The experiment uses a $\Xi$-VI with expectation propagation (EP) for approximating the marginals. Errors are quantified using the matrix L2 norm and contrasted across a spectrum of $\lambda$ values, including the theoretical lower bound at $\lambda=0$ and the diagonal EP approximation at $\lambda=\infty$. The vertical line marks the regularization parameter $\lambda=D$.
Additional Analysis of the Eight School Example
In this section, we present additional results for the eight school example in Section 4.3.
Table 1: $95\%$ posterior credible intervals for $\theta_{i}-\theta_{j}$ for 10 randomly selected pairs of schools.
Method $\theta_{2}-\theta_{5}$ $\theta_{6}-\theta_{7}$ $\theta_{2}-\theta_{4}$ $\theta_{4}-\theta_{8}$ $\theta_{1}-\theta_{2}$
MFVI [-12.08, 18.48] [-20.12, 11.37] [-13.99, 15.79] [-15.37, 14.97] [-13.47, 17.10]
True [-8.50, 14.90] [-17.74, 7.30] [-11.28, 12.40] [-13.02, 12.52] [-9.21, 16.55]
$\lambda=0$ [-7.81, 13.77] [-16.17, 6.49] [-10.42, 12.06] [-11.19, 12.70] [-8.12, 15.81]
$\lambda=1$ [-9.02, 13.54] [-16.71, 7.44] [-11.57, 11.87] [-11.73, 13.33] [-9.09, 15.97]
$\lambda=10$ [-10.43, 13.01] [-15.77, 8.75] [-12.57, 12.82] [-12.21, 14.50] [-10.54, 15.41]
$\lambda=1000$ [-10.51, 12.93] [-15.86, 9.53] [-12.55, 13.19] [-12.71, 14.47] [-10.79, 15.64]
Method $\theta_{2}-\theta_{8}$ $\theta_{3}-\theta_{8}$ $\theta_{5}-\theta_{6}$ $\theta_{2}-\theta_{7}$ $\theta_{3}-\theta_{4}$
MFVI [-14.83, 17.07] [-17.71, 13.88] [-16.77, 14.19] [-17.30, 13.41] [-17.62, 13.49]
True [-12.09, 12.73] [-16.08, 11.05] [-12.62, 10.31] [-14.97, 9.20] [-14.79, 10.45]
$\lambda=0$ [-10.15, 13.41] [-14.54, 11.56] [-11.12, 9.70] [-13.66, 8.31] [-14.33, 10.41]
$\lambda=1$ [-11.26, 13.57] [-14.31, 12.75] [-11.49, 10.27] [-15.28, 8.95] [-14.95, 11.25]
$\lambda=10$ [-11.58, 14.47] [-14.69, 13.74] [-11.15, 10.96] [-15.09, 10.55] [-15.18, 12.35]
$\lambda=1000$ [-12.14, 14.34] [-14.48, 14.32] [-12.00, 11.09] [-14.75, 10.63] [-15.21, 13.12]
Figure 9: Comparison of MCMC and MFVI posterior correlation matrices
Figure 10: Comparison of the approximation error of the standard deviation of the difference in treatment effect under MFVI and $\Xi$-VI with $\lambda=0,1,10,1000$, normalizing flow, full-rank ADVI, and Stein variational gradient descent.
Figure 11: $\Xi$-VI approximation accuracy for the Eight School model across varying $\lambda$ compared with MFVI, normalizing flow, full-rank ADVI, and Stein variational gradient descent, measured in KL divergence and Wasserstein-2 distance. A vertical line at $\lambda=D$ marks the infection point. $\Xi$-VI outperforms all other models in approximation accuracy under both criterions.